### v1

# AllSpark: a multimodal spatiotemporal general model

Abstract
RGB, infrared, multispectral, and other modal data fundamentally represent different observational approaches to the same geographic object. Therefore, leveraging multimodal data is an inherent requirement for comprehending geographic objects. However, for a long time, due to the high heterogeneity in structure and semantics among various spatiotemporal modal data, the joint interpretation of multimodal spatiotemporal data has been an extremely challenging problem. The primary challenge resides in striking a trade-off between the cohesion and autonomy of diverse modalities, and this trade-off exhibits a progressively nonlinear nature as the number of modalities expands. Inspired by the human cognitive system and linguistic philosophy, where perceptual signals from the five senses converge into language, we introduce the Language as Reference Framework (LaRF), a fundamental principle for constructing a multimodal unified model, aiming to strike a trade-off between the cohesion and autonomy among different modalities. Building upon this, we propose a multimodal spatiotemporal general artificial intelligence model, called AllSpark. Our model integrates thirteen different modalities into a unified framework, including one-dimensional (text, code), two-dimensional (RGB, infrared, SAR, multispectral, hyperspectral, tables, graphs, trajectory, oblique photography), and three-dimensional (point clouds, videos) modalities. To achieve modal cohesion, AllSpark uniformly maps diverse modal features to the language modality. In addition, we design modality-specific prompts to guide multi-modal large language models in accurately perceiving multimodal data. To maintain modality autonomy, AllSpark introduces modality-specific encoders to extract the tokens of various spatiotemporal modalities. And modal bridge is employed to achieve dimensional projection from each modality to the language modality. Finally, observing a gap between the model’s interpretation and downstream tasks, we designed task heads to enhance the model’s generalization capability on specific downstream tasks. Experiments indicate that AllSpark, without expert knowledge of the most spatiotemporal modalities and utilizing a unified structure, achieves competitive accuracy in modalities such as RGB and trajectory compared to state-of-the-art models. Moreover, AllSpark showcases excellent adaptability in modalities like MSI, HSI, PointCloud, Table, Code, and Graph. AllSpark demonstrates the potential and possibility of constructing general artificial intelligence with a large language model. This approach contributes to the shift in research paradigm in spatiotemporal intelligence, transitioning from a modality-specific and task-specific paradigm to a general paradigm.
keywords:
Multimodal, Language, Spatiotemporal, General Model
1 Introduction
Benefiting from the increasingly diverse observational methods for spatiotemporal scenes, geographic objects can be described using various spatiotemporal modalities such as RGB, infrared, SAR, multispectral, graphs, and spatiotemporal trajectories. Each modality provides unique information about different aspects of the geographic object. Analogous to the human process of perceiving and understanding the world through multiple modalities such as vision, hearing, and touch, joint interpretation of multimodal data is an inherent requirement for intelligent models to achieve cognition of geographic objects.
However, due to the inherent differences in the mechanisms of each modality, various modalities are often highly heterogeneous in both structure and semantics. For example, in terms of structure, a table is composed of rows and columns. Point clouds are represented by three-dimensional coordinates along with various feature values. Text is constituted by sequences of words. In terms of semantics, RGB imagery reflects the electromagnetic characteristics of visible light bands emitted and reflected by geographic objects, whereas SAR imagery reflects the electromagnetic characteristics of microwaves emitted actively by radar for scattering from geographic objects.
For a long time, constrained by the highly heterogeneity across various modalities as mentioned above, researchers have often developed specific methods based on prior assumptions related to a particular modality or designed multimodal approaches based on a few low-heterogeneity modalities. For instance, in single-modal research, C. R. Qi et al. proposed PointNetqi2017pointnet for the point cloud modality, emphasizing the invariance of point cloud data ordering and the significance of global and local features. For the language modality, Vaswani et al. introduced Transformervaswani2017attention , focusing on the long-range dependencies within word sequences. For the graph modality, Kipf and Welling proposed GCNkipf2016semi , based on the adjacency relationships between nodes in a graph. In multimodal research, fusion of optical and SAR imagery has been widely explored in both traditional and deep learning remote sensingkulkarni2020pixel . Text-visual models have also experienced rapid development in recent yearsradford2021learning ; lu2019vilbert ; kim2021vilt ; alayrac2022flamingo ; li2022blip . The diverse prior assumption associated with each modality have resulted in significant gaps between methods designed for different modalities, making it challenging to perceive and understand different modalities using a unified model.
We believe that the key challenge in addressing this issue lies in striking a trade-off between the cohesion and autonomy of diverse modalities. Cohesion refers to the presence of mutually correlated shared information among modalities, while autonomy signifies the existence of unique information specific to each modality relative to others. If we simply project data from multiple modalities into a common representation space for fusion, emphasizing the cohesion between modalities, it may lead to the loss of modality-specific information, thereby diluting the essence of multimodal collaboration. On the contrary, if we excessively stress the autonomy between modalities, it may hinder the establishment of connections among them, limiting the model’s ability to simultaneously perceive multiple modalities. Moreover, as the number of modalities increases, balancing the cohesion and autonomy becomes progressively more challenging in a nonlinear manner.
We observe that in the process of comprehending the world, humans integrate information from multiple modalities such as hearing, touch, smell, and vision. The concepts formed through the parsing of these modalities are ultimately converged in language. It is through language that humans engage in associating, reasoning, and expressive behaviors. In other words, language precisely encodes human perception and understanding of the world, providing clear definitions and meanings to abstract concepts from each modality. Inspired by this, we propose Language as Reference Framework (LaRF) as a fundamental principle for constructing multimodal models. As illustrated in Figure 1, the abstract concepts derived from each modality should align with language, allowing for joint interpretation in the unified representation space of language.
Figure 1: Language as Reference Framework (LaRF)
Building upon this, we propose a multimodal spatiotemporal general artificial intelligence model, AllSpark, integrating thirteen different modalities into a unified framework, including one-dimensional (text, code), two-dimensional (RGB, infrared, SAR, multispectral, hyperspectral, tables, graphs, trajectory, oblique photography), and three-dimensional (point clouds, videos) modalities.
To achieve modal cohesion, AllSpark uniformly maps diverse modal features to the language modality. In addition, we design modality-specific prompts to guide multi-modal large language models in accurately perceiving multimodal data.
To maintain the autonomy between modalities, AllSpark introduce specific modal encoders for each modality to extract independent tokens. Given the high heterogeneity among modality data and modality encoders, there exists significant dimensional gap between the tokens of each modality and the language modality. To address this issue, we introduce modality bridge, a mechanism from Lynxzeng2023matters to accomplish the dimensional mapping from each modality’s tokens to the language tokens.
Finally, considering the existing gap between the interpretation by the multimodal large language model and the specific downstream tasks, we have designed task heads for each downstream task to enhance the model’s generalization capability. It is noteworthy that, given the powerful interpretation capabilities of the multimodal large language model, we adhere to a lightweight design principle in both the modal encoders and task heads
Experiments demonstrate that AllSpark, without the expert knowledge in most modalities and utilizing a unified structure, achieves competitive accuracy in modalities such as RGB and spatiotemporal trajectories compared to state-of-the-art models. Specifically, in the RGB modality, AllSpark’s accuracy is only 0.84 lower than the optimal model, and in the trajectory modality, the ADE metric differs by only 0.07 compared to the optimal model. Additionally, AllSpark exhibits excellent adaptability in various other modalities, including point clouds, multispectral, hyperspectral, tables, graphs, and code. Theoretically, our proposed model has the potential for seamless extension to arbitrary number of modalities.
In a word, our contributions can be summarized as follows:
  • 1.
    To our best knowledge, we first propose a unified multimodal spatiotemporal general model, AllSpark, successfully integrating 13 spatiotemporal modalities into a single model.
  • 2.
    Inspired by the human cognitive system and linguistic philosophy, we propose the Language as Reference Framework (LaRF), offering a novel solution to balance cohesion and autonomy among multiple modalities. This principle provides a fresh approach to the multimodal unification.
  • 3.
    We provide new possibilities for the paradigm of designing multimodal models in the spatiotemporal domain. MSTAGI demonstrates the potential and capabilities of building general model with large language models, thereby contributing to the shift in research paradigm in spatiotemporal intelligence, transitioning from a modality-specific and task-specific paradigm to a general paradigm.
  • 4.
    Experiments indicate that AllSpark, without expert knowledge of the majority of spatiotemporal modalities and utilizing a unified structure, achieves competitive accuracy in modalities such as RGB and trajectory compared to state-of-the-art models. Moreover, AllSpark showcases excellent adaptability in modalities like MSI, HSI, PointCloud, Table, Code, and Graph. Theoretically, our proposed model has the potential for seamless extension to arbitrary number of modalities.
2 Related Work
Leveraging multimodal data is an inherent requirement for achieving cognitive recognition of geospatial objects. An ideal multimodal model should possess the capability to integrate all modalities for joint interpretation. Hence, a crucial trend in the research of intelligent methods in the spatiotemporal domain is the continual increase in the number of modalities for joint interpretation. Initially, early researchers often constructed single-modal expert models based on prior assumptions about a specific modality, achieving remarkable success within each respective modality. In recent years, with a deeper understanding of single-modal interpretation methods, numerous researchers have attempted to integrate several low-heterogeneity modalities to build multimodal interpretation approaches. However, as the number of modalities increases, the challenge of balancing cohesion and autonomy among the modalities becomes increasingly difficult. In the following, we recall the development of intelligent methods in the spatiotemporal domain from the perspective of the continually increasing number of modalities, and finally, we present the principles and approach of our proposed model, AllSpark.
2.1 Single-Modal Model
For one-dimensional modalities, we focus on Code and Text. Code and Text are fundamentally language modalities, but due to the excellent characteristics of Code, such as strict syntax, unambiguous nature, and its ability to interact with machines, we treat it as a separate modality. Z. Feng et al. have pre-trained a model, CodeBERTfeng2020codebert , which facilitates the mutual transfer of information between Code and natural language modalities. For the Text modality, landmark contributions include Transformervaswani2017attention ,BERTkenton2019bert ,GPT series radford2018improving ; radford2019language ; brown2020language ; openai2023gpt4 , which have inspired a subsequent series of works.
We categorize RGB, multispectral, hyperspectral, infrared, SAR, Table, Graph, Trajectory, and Oblique Photography as two-dimensional modalities.
Among the two-dimensional modalities, the standard three-channel RGB image is one of the most common. For the RGB modality, the ResNethe2016deep proposed by K. He et al., based on the importance of visual global and local information, and the ViTdosovitskiy2020image introduced by A. Dosovitskiy et al., leveraging a global attention mechanism, represent two landmark contributions.
An increase in the number of channels in images leads to the multispectral and hyperspectral modalities. B. Huang et al. proposed the STDCNNhuang2018urban , leveraging the characteristic of a higher number of bands in multispectral images to simultaneously model the global spatial and spectral properties of multispectral images. In comparison to multispectral images, hyperspectral images have even more bands, often reaching hundreds. Based on this, X. Yang et al. introduced the R-3D-CNNyang2018hyperspectral to further enhance the extraction of spectral features.
Beyond an increase in the number of bands, attention has been directed towards images captured in specific spectral bands. For instance, infrared images, widely applied in fields such as agriculture and oceanography, have gained significant attention. H. Chen et al. proposed a shallow convolutional neural network, integrating a decision tree, to meet the demand for real-time intelligent interpretation of a large volume of infrared images in agricultural water resource monitoringchen2020deep . In the case of SAR images, formed by active microwave radar, S. Chen et al. introduced AConvNetchen2016target , a widely used fully convolutional neural network for SAR image intelligent interpretation.
For oblique photography, 3D reconstruction aims to utilize single or multiple two-dimensional images from different angles to reconstruct a three-dimensional model. Y. Yao et al. addressed this task by proposing MVSNetyao2018mvsnet . Utilizing techniques such as three-dimensional convolution, the model establishes correlations between images from different perspectives, thereby reconstructing the three-dimensional model of the target from the features extracted from two-dimensional images.
Beyond the visual domain, Table is one of the commonly used modalities for recording and expressing information. TabNetarik2021tabnet , proposed by S. Ö. Arik et al., employs a sequence attention mechanism to achieve feature selection in the Table modality, thereby enabling interpretable and more efficient learning.
The trajectory modality reflects the temporal changes in the spatial positions of objects. A. Gupta et al. proposed Social GANgupta2018social based on the characteristic of trajectory multiplicity. This model combines historical trajectory information with social context to predict multiple plausible future outcomes.
For the Graph modality, Kipf and Welling introduced the classic Graph Convolutional Network (GCN)kipf2016semi , which is based on the adjacency relationships between nodes in the graph. P. Velicˇkovic et al. proposed GATvelivckovic2017graph , which incorporates attention mechanisms into the Graph modality.
Finally, we turn our attention to the three-dimensional modalities: Video and Point Cloud.
The Video modality is formed by stacking images over time. A model proposed by A. Karpathy et al. extends the connectivity of CNN models over time. Additionally, it utilizes multi-resolution techniques to accelerate the training process, effectively reducing the impact of spatiotemporal redundancy on model efficiency in the Video modalitykarpathy2014large .
The Point Cloud modality captures information about the position, shape, color, texture, and other aspects of three-dimensional objects. C. R. Qi et al. introduced the classic PointNetqi2017pointnet for the Point Cloud modality, emphasizing the importance of invariance through point data permutation and the significance of global and local features. W. Wu et al. extended convolution operations to three-dimensional point clouds with the introduction of PointConvwu2019pointconv .
2.2 Multi-modal Model
While the single-modal methods in Section 2.1 have demonstrated excellent performance within their respective modalities, they often face challenges in generalizing across multiple modalities due to their construction based on specific prior assumptions. Recognizing the intrinsic requirement for intelligent models to utilize multimodal data for geographic object cognition, numerous researchers have endeavored to balance the cohesion and autonomy among modalities to build multimodal models.
For joint interpretation of two modalities, H. Li and X.-J. Wu recognized the enhancement possible through the fusion of RGB and Infrared images, which convey information about color, texture, and temperature. Given the widespread demand and application prospects in military, medical, and other fields, they proposed a dense fusion method for RGB and Infrared images: DenseFuseli2018densefuse . This method, employing a dense module that connects the outputs of all convolutional layers, successfully achieves pixel-level fusion of RGB and Infrared modalities. A. Sadeghian et al. extended the use of historical trajectories for predicting multiple plausible trajectories in Social GANgupta2018social . They introduced RGB images to enhance scene data in Sophiesadeghian2019sophie , achieving better results in trajectory prediction tasks. Recognizing the high complementarity between RGB and SAR images, L. H. Hughes et al. proposed a three-step deep neural network framework that utilizes a universal prediction of matching regions, generates heatmaps, and eliminates outliers to match RGB and SAR imageshughes2020deep . X. Li et al. introduced DTCDNli2021deep , a model that employs a GAN network to migrate RGB and SAR images to the same feature space, facilitating target detection. J. Yang et al. proposed a dual-stream convolutional network that uses high-resolution multispectral images to enhance the spatial resolution of hyperspectral imagesyang2018hyperspectralandmultispectral .
Notably, CLIPradford2021learning proposed by A. Radford et al. associates the RGB modality with the text modality using contrastive learning. With pretraining guided by weak supervisory signals from text, CLIP demonstrates outstanding capabilities in both visual single-modal tasks and visual-text multimodal tasks, inspiring a series of subsequent workslu2019vilbert ; kim2021vilt ; alayrac2022flamingo ; li2022blip . Y. Zhang et al. introduced Meta-transformerzhang2023meta , leveraging the contrastive learning paradigm from CLIP to pretrain a universal backbone network under the visual-language modality. It exhibits multimodal generalization abilities across various modalities such as point clouds, infrared, and hyperspectral. J. Han et al. directly employed a multimodal large language model as a universal backbone network, proposing the One-LLM, successfully unifying eight modalities including Image, Audio, Video, and Pointhan2023onellm . The success of these approaches imply the unique role of language modalities in multimodal model.
Therefore, building upon the aforementioned efforts, we systematically propose the fundamental principle, Language as Reference Framework (LaRF). Guided by this principle, we balance the cohesion and autonomy among diverse modalities and introduce a general intelligent model named AllSpark, which unifies 13 spatiotemporal modalities and possesses the potential to extend to an arbitrary number of modalities.
3 Method
3.1 Theory
We observe that in the process of comprehending the world, humans integrate information from multiple modalities such as hearing, touch, smell, and vision. The concepts formed through the parsing of these modalities are ultimately converged in language. It is through language that humans engage in associating, reasoning, and expressive behaviors. In other words, language precisely encodes human perception and understanding of the world, providing clear definitions and meanings to abstract concepts from each modality.
Inspired by our observation, to balance the cohesion and autonomy among multiple modalities, we introduce the fundamental principle, Language as Reference Framework (LaRF). The significance is:
  1. 1.
    Alignment Capability: Language can accurately encode both cohesion and autonomy information across multiple modalities. Aligning each modality with the language modality enables the attainment of a unified representation in the same feature space, addressing the challenge of high heterogeneity among modalities.
  2. 2.
    Reasoning Capability: Language, being a tool for human thought and expression, inherently possesses the ability to perform complex reasoning. Each modality, when represented in a unified space with LaRF, inherits the reasoning capability of language, unlocking the potential for multimodal joint reasoning.
  3. 3.
    Interpretability: Deep learning have long been characterized as opaque ”black boxes.” However, a multimodal intelligent system constructed based on the LaRF can directly leverage language as a tool. This facilitates the output of interpretable reasoning chains that humans can understand, thereby achieving true explainable artificial intelligence.
  4. 4.
    Interactivity: Language not only aids humans in understanding intelligent models but also facilitates intelligent models in understanding humans. In an intelligent system guided by the LaRF, humans can directly express their needs using natural language. This iterative correction of the model’s output based on human interaction will become a new paradigm for the training and inference of intelligent models.
  5. 5.
    Scalability: The multimodal system guided by the LaRF is agnostic to specific modalities. New modalities only need to establish a mapping to the language model to participate in joint reasoning with other modalities. Therefore, theoretically, a multimodal model based on LaRF can be extended to an arbitrary number of modalities.
In summary, we contend that intelligent models, when engaged in multimodal perception processes akin to those of humans, should adopt language as the reference framework for abstract concepts across modalities. The analysis and reasoning should occur in the feature space of language. As illustrated in the conceptual model in Figure 2, when constructing specific multimodal models, data from each modality should be encoded into modality-independent tokens with contextual information. These modality-specific tokens should align with the tokens of language. A universal intelligent model, primarily based on the Transformer architecture, can then be employed for parsing and reasoning, ultimately generating a comprehensive multimodal representation.
Figure 2: The conceptual model with LaRF
3.2 Overview
Table 1: List of symbols
Module Variants
Symbol
Meaning
Symbol
Meaning
$M$
AllSpark
$m_{i}$
modal i
$f_{i}$ modal encoder for $q$ query vectors of the bridge
$\Phi$ modal bridge $s_{i}$ tokens of $m_{i}$
T text tokenizer $p_{i}$ text prompt for $m_{i}$
$F$
Multi-modal LLM
$\theta$
model parameters
$H_{task}$
Task Head
$y$
label
$L$ loss function $t_{i}$ tokens of $m_{i}$
${Enc}_{n}$
n-layer transformer encoders
$W$
weights of the linear layer
$Emb$
embedding layer
$w_{i}$
word of the code or text modal
${Enc}_{resnet}$
resnet
$Sparse$
sparse module of the TubeViT
$Grouper$
point grouper of the PointBERT
$Conv1d$
1D convolution layer
$FFN$
feed forward network
Guided by the principles mentioned above, the AllSpark we proposed consists of five modules: Modal Encoder, Modal Bridge, Text Tokenizer, Multimodal Large Language Model, and Task Head. The overall architecture is depicted in Figure 3.
Figure 3: AllSpark Architecture
To maintain the autonomy among modalities, we designed modality-specific encoders to encode highly heterogeneous data into modality-independent tokens (Section 3.3 for details). However, the dimens ion of tokens outputted by modal encoders are still inconsistent. To achieve the goal of parsing in the unified representational space of language, we introduced the modal bridge from the Lynxzeng2023matters . The modal bridge aims to project tokens from each modality into the dimension of the Multimodal Large Language Model (Section 3.4 for details). The formalization of this process is as follows:
$s_{i}=\Phi\left(f_{i}\left(m_{i}\right),q\right)$ (1)
Here, $\{m_{RGB},m_{MSI},m_{HSI},\ldots,m_{i},\ldots,m_{Video}\}$ represents inputs from various modalities, where $m_{RGB}$ represents the RGB modality, $m_{MSI}$ represents the MSI modality, and so on. $f_{i}$ is defined as the modal encoder for the $m_{i}$ modality, $\Phi$ represents the modal bridge, $q\in R^{N*D}$represents $N$ learnable vectors of dimension $D$ in the modality bridge, where $D$ is set to 4096, representing the dimensionality of the Multimodal Large Language Model. Each modality’s input data $m_{i}$ is mapped to a token sequence $s_{i}\in R^{N*D}$ of the same dimension as the language model. Table 1 summarizes the main mathematical symbols and their meanings in the model.
To achieve cohesion among modalities, we employ a unified Multimodal Large Language Model to parse data from various modalities. The text tokenizer and Multimodal Large Language Model in AllSpark are based on the visual-language model Lynxzeng2023matters . To extend Lynx to 13 spatiotemporal modalities, we design specific text prompts for each modality to guide the model in correctly parsing information from each modality. Additionally, Lynx incorporates several lightweight multimodal adapter layers internally to accommodate multimodal inputs. We continue this design and do not freeze the parameters of the adapter layers during training to enhance the model’s adaptability to other spatiotemporal modalities. Finally, we acknowledge that there exists a gap between the parsing results of the model and the downstream task. Therefore, we design specific task heads for each task to enhance the model’s generalization capability.
The entire model M can be formalized as:
$M\left(m_{i},p_{i}\right)=H_{task}\left(F\left(s_{i}\oplus T\left(p_{i}\right)\right)\right)$ (2)
Where $p_{i}$ represents the text prompt of $m_{i}$, $T$ denotes the text tokenizer, $\oplus$ indicates the concatenation operation of text tokens and modality tokens in the sequence, $F$ represents the Multimodal Large Language Model, and $H_{task}$ signifies the task head.
All tasks in our experiments are supervised tasks, with $y$ denoting the labels, $L$ representing the loss function, and $\theta$ representing the learnable parameters in our model. The optimization objective of the model can be formalized as:
$\theta_{i}=\mathop{\arg\min}_{\theta}{L\left(y,M\left(m_{i},p_{i}\right),\theta_{i}\right)}$ (3)
3.3 Independent Encoder for Each Modality
The modal encoder aims to encode the raw data of each modality into a token sequence, formalized as: $t_{i}=f_{i}\left(m_{i}\right)$, where $t_{i}\in R^{n*d}$. We have designed different modal encoders for each modality to maintain autonomy among modalities. The following provides individual introductions for each modality:
3.3.1 1D modal
Code & Text: Code is essentially a specialized form of language, but due to its distinct properties—such as having a strict syntax, being unambiguous, and being capable of interacting with machines—we separate it as a distinct modality. Exploring intelligent methods for the Code modality is crucial for reliable AI reasoning and achieving interaction between intelligent models and the real world. Since the Lynx model is a language model, we do not design an additional modal encoder for the Code and Text modalities. Instead, we directly utilize Lynx’s text tokenizer, i.e., $f\left(m_{Text/Code}\right)=T\left(m_{Text/Code}\right)$, where $m_{Text/Code}=\{w_{1},w_{2},w_{3},\ldots\}$ represents the sequence of words in text or code.
3.3.2 2D modal
RGB: RGB imagery represents the visible light spectrum, reflecting the electromagnetic characteristics of objects that emit or reflect visible light waves. It is the most common modality in visual research. RGB imagery is a standard three-band image, i.e., $m_{RGB}\in R^{H*W*3}$. For the modal encoder of this modality, we adopted the visual encoder from the Lynx model: EVA. EVA is a large visual backbone model composed of 40 stacked Transformer Blocks with a width of 1408. During the experiments, AllSpark loaded the official weights of the EVA model and froze them for training.
MSI: Multispectral imagery is a modality extensively studied in the remote sensing field. Typically, it includes more than the three visible light bands present in the RGB modality. MSI incorporates multiple non-visible light bands such as near-infrared, short-wave infrared, coastal atmospheric aerosol, and cirrus bands. Therefore, the number of channels in the imagery is usually greater than the three bands in RGB imagery, i.e., $m_{MSI}\in R^{H*W*C}$, where $C>3$. We extended the PatchEmbed of the standard Vision Transformerdosovitskiy2020image , modifying its channel count to match the number of bands in the input multispectral imagery. This adaptation allows it to serve as the feature encoder for the multispectral modality.
HSI: Hyperspectral imagery increases the number of bands compared to multispectral imagery, often reaching hundreds of bands, with each band containing rich information about the terrain. Unlike RGB and multispectral imagery, where an image serves as a single sample, in hyperspectral imagery, all bands for each pixel are treated as a single sample, i.e., $m_{HSI}\in R^{1*1*C}$. In the modal encoder for hyperspectral imagery, we first use a linear projection layer to expand the feature dimensions for each pixel. This process is formalized as $W*m_{HSI}^{T}$, where $W\in R^{1*d}$ is the weight matrix of the linear projection layer. Subsequently, we use a 12-layer standard Transformer encoder to extract its features. The entire process can be represented as $f\left(m_{HSI}\right)={Enc}_{12}\left(W*m_{HSI}^{T}\right)$.
Table: The Table modality consists of several rows and columns, where each row represents a record, and each column represents a different attribute, i.e., $m_{Table}\in R^{{row}*{col}}$. The modal encoder for the Table modality inherits the design from TabFormerpadhi2021tabular : first, based on the different degrees of discreteness for each column attribute, we use independent Embedding Layers to encode discrete and continuous values separately. Subsequently, we employ a single-layer Transformer Encoder to further extract features. The entire process can be formalized as $f\left(m_{Table}\right)={Enc}_{1}\left(Emb(m_{Table})\right)$.
Trajectory: The trajectory modality reflects the changing information of an object over time and space, composed of a series of two-dimensional coordinate points, i.e., $m_{Trajectory}\in R^{l*2}$, where $l$ represents the sequence length of trajectory points. The encoder for the trajectory modality inherits the design from TUTRshi2023trajectory : first, a linear layer is used to expand the dimensions of the two-dimensional trajectory features. This step can be formalized as ${W*m}_{Trajectory}$, where $W\in R^{d*\left(l*2\right)}$ is the weight matrix of the linear projection layer. Then, we use a 2-layer Transformer Encoder to extract its features. The entire process can be formalized as $f\left(m_{Trajectory}\right)={Enc}_{2}\left({W*m}_{Trajectory}\right)$.
SAR: The Synthetic Aperture Radar (SAR) modality is a type of active remote sensing that reflects the electromagnetic characteristics of objects with respect to microwave backscatter. Due to different polarization modes, the final product of SAR imagery is typically a two-band or single-band image, i.e., $m_{SAR}\in R^{H*W*2}$. Therefore, we designed a simple three-layer convolutional network as its modal encoder.
Infrared: The dataset we selected for infrared is SYSU-MM01, which is, in fact, a multi-modal re-identification dataset comprising both visible and near-infrared modalities. The model needs to simultaneously process infrared and visible light images, i.e., $m_{Infrared}=\{{I_{r}\in R}^{H*W*3},{I_{i}\in R}^{H*W*3}\}$. We inherited and slightly modified the design of DDAGye2020dynamic , discarding the shared ResNet-50 layers’ parameters in DDAG and employing two independently parameterized ResNet-50 networks. These networks extract features from infrared and visible light images, and this process can be formalized as $f\left(m_{Infrared}\right)={Enc}_{resnet}\left(I_{r}\right)\oplus\ {Enc}_{resnet}\left(I_{i}\right)$.
Graph: A graph is constructed from a series of nodes and edges, where the attributes of the nodes and the adjacency attributes of the nodes reflect the majority of the graph’s features, i.e., $m_{Graph}\in R^{K*d}$, where K is the number of nodes, and d is the node feature dimension. In AllSpark, the graph’s modality encoder is based on STAEformerliu2023spatio , with the main design idea being to first use a linear layer to extend its feature dimension and then use several Embedding Layers to separately encode features such as the node’s own characteristics, spatial characteristics, and temporal characteristics in the graph. The entire process can be formalized as $f\left(m_{Graph}\right)={Emb}_{node}\left({W*m}_{Graph}^{T}\right)\oplus{Emb}_{spatial}\left({W*m}_{Graph}^{T}\right)\oplus{Emb}_{time}\left({W*m}_{Graph}^{T}\right)$, where $W\in R^{{hidden}*d}$ is the weight of the linear layer.
Oblique Photography: The oblique photography modality consists of images from multiple views, i.e., $m_{video}=\{V_{1}\in R^{C*H*W},\ V_{3}\in R^{C*H*W},V_{3}\in R^{C*H*W},\ldots\ ,V_{n}\in R^{C*H*W}\}$, where $V_{i}$ represents the image of each view, and n is the number of views. In the encoder for oblique photography modality, we use the standard ViT model to extract features from multiple view images and concatenate them. The process can be formalized as $f\left(m_{video}\right)={Enc}_{12}\left(V_{1}\right)\oplus{Enc}_{12}\left(V_{2}\right)\oplus{Enc}_{12}\left(V_{3}\right)\oplus\ldots\oplus{Enc}_{12}\left(V_{n}\right)$.
3.3.3 3D modal
Video: The video modality can be considered as the stacking of images in the temporal dimension, i.e., $m_{video}\in R^{T*C*H*W}$. The significant characteristic of the video modality is its high redundancy in both temporal and spatial dimensions. To address this issue, the encoder for the video modality in AllSpark is based on TubeViTpiergiovanni2023rethinking , which performs sparse sampling on the highly redundant data in both time and space to reduce redundancy. The features obtained from sparse sampling are then fed into a 6-layer Transformer Encoder for further feature extraction. The process can be formalized as $f\left(m_{video}\right)={Enc}_{6}\left(Sparse(m_{video})\right)$.
PointCloud: PointCloud are typically composed of three-dimensional coordinates and their feature values, $m_{pointCloud}\in R^{K*\left(d+3\right)}$, where K represents the number of three-dimensional points, and d represents the dimensionality of the point cloud features, reflecting information such as the spatial position, shape, color, and texture of objects. The encoder for the point cloud modality inherits the design from PointBERTyu2022point : firstly, point cloud data is grouped and encoded to unify the number of points simultaneously inputted. This step can be represented as $PointGroup=Grouper\left(m_{pointCloud}\right)\in R^{G*N*3}$, where $G$ represents the number of groups, and $N$ represents the number of points in each group. Next, the grouped results are inputted into a one-dimensional convolutional layer to extract feature vectors for each group:$f_{Group}=Conv1d\left(PointGroup\right)\in R^{G*d}$. Finally, the feature vectors for each group are inputted into a standard 12-layer Transformer Encoder to extract their global features. The entire process can be formalized as $f\left(m_{pointCloud}\right)=Enc_{12}\left(Conv1d(Grouper(m_{pointCloud}))\right)$.
3.4 Modal2Language Bridge
The modal bridge, based on the Lynx model, aims to perform dimensional projection from tokens of various modalities to tokens of the language modality. In its implementation, the modal bridge consists of stacked layers, including cross-attention layers and feedforward neural network layers.
In the cross-attention layer, we predefine a learnable query vector $Q\in R^{N*D}$, where $D$ is the internal dimensionality of the language model, and $N$ serves as a hyperparameter that can be flexibly adjusted to accommodate inputs from different modalities. The keys and values in the cross-attention layer are the features outputted by the modal encoders.
The feedforward neural network inherits the classic design from the original Transformer, consisting of two linear layers with an inserted activation layer. The entire process can be formalized as:
$\Phi\left(Q,s_{i}\right)={FFN}\left({softmax}\left(\frac{W_{q}Q^{T}\left(w_{k}s_{i}^{T}\right)^{T}}{\sqrt{D}}\right)*(W_{v}s_{i}^{T})\right)$ (4)
Here, $W_{q}\in R^{D*{hidden}}$, $W_{k}\in R^{d*{hidden}}$, and $W_{v}\in R^{d*{hidden}}$are the linear projection layer weights defined inside the cross-attention layer for $Q$, $K$ and $V$ respectively.
3.5 Task-Guided Text Prompts and Task-Heads
To extend the visual-language multimodal model to 13 spatiotemporal modalities without intervention from modality expert knowledge, we designed specific text prompts and task heads for each modality and task. Text prompts are used to guide the multimodal language model in correctly interpreting each modality’s data, while task heads are employed to match the model’s parsing results with specific downstream tasks.
We manually designed one to four specific text prompts for each modality. During the training process, to enhance model performance, we employed a strategy of diversifying prompts, randomly selecting one prompt for each forward pass. However, during testing, for the sake of result stability and reproducibility, the prompt was fixed to be the first prompt among all prompts. Table 2 provides a list of all the text prompts.
Table 2: List of prompts
Dimension
Modal
Prompt
1D
Text
1.Please determine if this movie review is positive or negative?
Code
2D
RGB
1.This remote sensing image belongs to which of the following categories: [Category list of the NWPU dataset] 2.Describe this remote sensing image briefly 3.Find a word that is most relevant to this remote sensing image 4.Describe the key elements in this remote sensing image
MSI
1.Based on the multi-spectral imagery feature description, please classify this object 2.Given the following multi-spectral imagery characteristics, please output the most fitting scene label
HSI
1.Given the spectral information, can you help determine which class this pixel belongs to? 2.Here is the spectral data for a pixel. Considering the typical characteristics of land cover classes, could you provide a detailed analysis and suggest the most likely class for this pixel? 3.The spectral information for a pixel is given, but the data is noisy. Given the potential variability, which land cover classes should be considered as possible candidates for this pixel?
Table
1.Please utilize the provided air quality indicators to accurately predict the concentrations of PM2.5 in the atmosphere 2.Given the air quality indicators, please provide a prediction for the PM2.5 levels at this particular moment
Oblique Photography
1.Please reconstruct the 3D model from the given images of five different views.
Trajectory
1.Based on their past positions and movements in a crowded environment, predict the future trajectory of a selected pedestrian. 2.Using the pedestrian trajectory data, along with additional information about the surrounding environment, predict the future path of the pedestrian. 3.Given the current and past positions of a pedestrian and their neighboring pedestrians, predict the main pedestrian’s trajectory
SAR
1.Based on the SAR imagery feature description, please classify this object 2.Given the following SAR imagery characteristics, please output the most fitting scene label
Infrared
1.Given an infrared image of a person, find and highlight the same person in a set of visible light images, paying close attention to features like clothing, posture, and gait 2.Analyze the characteristics of a person in a visible light image, such as clothing texture, color, and body shape. Then, locate the person with matching features in a series of infrared images
Graph
1.Given the current traffic data including vehicle flow rate, average speed, and time of day from the METRLA dataset, predict the traffic flow 2.Analyze the historical data on vehicle speeds and flow rates from the METRLA dataset for the past week. Identify any patterns or trends and predict the traffic conditions
3D
Video
1.Please analyze this video and describe the characteristics of the action in detail 2.Watch this short clip and provide a step-by-step description of the action sequence
Point Cloud
1.Classify the provided point cloud sample into the correct category 2.Look at the point cloud data characteristics and classify the object 3.Please analyze the given point cloud dataset and determine which category it belongs to. Focus on the shape and structure evident in the point cloud
Additionally, we noticed that there is still a gap between the multimodal language model’s parsing results and the specific downstream tasks. To enhance the model’s generalization capability on downstream tasks, we designed specific task heads for each task. Moreover, to ensure the transferability of modalities across different tasks, the design principle for task heads is to be as simple and lightweight as possible. The detailed summary of each modality’s task heads can be found in Table 4.
For classification tasks or downstream tasks that can be formalized as classification tasks, we uniformly use a simple single-layer linear layer as the task head. For instance, we implemented standard classification tasks on the RGB, MSI, SAR, Video, and PointCloud modalities. Although the task on the HSI modality is a segmentation task, it can be formalized as a per-pixel classification task. Similarly, for the re-identification task on the Infrared modality, it can also be formalized as a classification task. Therefore, a single-layer linear layer is used as task head for the mentioned modalities.
For regression tasks on the Table, Trajectory, and Graph modalities, we also use a linear layer to perform regression predictions. The only difference from classification tasks is the addition of an unscaled operation without learnable parameters. Since the Lynx model itself is a language model, the Code and Text modalities directly use its native text decoder. The task head for the Oblique Photography modality is based on the 3D reconstruction architecture from the Ada-MVSliu2023deep model.
4 Experiment
4.1 Setup
Our experiments aimed to evaluate the adaptability of AllSpark to various spatiotemporal modalities. To provide a realistic evaluation of its multimodal adaptability and facilitate comparisons with existing work in the field, we followed the principles of simplicity, reproducibility, and choosing datasets and tasks that have been widely studied in the domain. The specific details on dataset selection refer to Section 4.2.
Additionally, for result reproducibility, we employed similar experimental settings across all modalities. We used the AdamW optimizer with a learning rate schedule based on cosine annealing. The hyperparameters were adjusted slightly in terms of training epochs and learning rates for different experiments. Table 3 summarizes the hyperparameter settings for experiments on each modality.
Table 3: Summary of hyperparameters
Dimension
Modal
Max Lr
Max Epochs
Warm-up Epochs
1D
Text
9.0*10-6
5 1
Code
1.0*10-5
4 1
2D
RGB
5.0*10-5
50 5
MSI
2.0*10-5
50 5
HSI
1.0*10-4
30 3
Infrared
5.0*10-5
50 5
SAR
9.0*10-6
30 3
ObliquePhotography 5.0*10-5 30 3
Table
2.0*10-5
30 3
Trajectory
1.0*10-5
30 5
Graph
8.0*10-5
10 2
3D
PointCloud
3.0*10-5
100 10
Video
1.0*10-5
3 1
4.2 Dataset
Table 4 summarizes the data and tasks used in experiments for each modality. Below, we will provide detailed explanations in order:
Text: The IMDB dataset is a binary sentiment analysis dataset, consisting of 50,000 reviews from the Internet Movie Database (IMDb) labeled as positive or negative. Additionally, the dataset includes some unlabeled data. In our experiments, only the labeled data from the IMDB dataset was utilized for supervised sentiment classification tasks.
Code: CodeSearchNet is a large-scale dataset of function code and its documentation from GitHub, covering six programming languages: Go, Java, JavaScript, PHP, Python, and Ruby. The task performed on the code modality is Code Document Generation, and we tested it on the Ruby and JavaScript languages.
RGB: NWPU-RESISC45 is a large-scale open dataset for visible light remote sensing image scene classification. It includes 45 land-use categories such as airplane, baseball diamond, beach, commercial area, etc. Each category has 700 remote sensing images, making a total of 31,500 images in the dataset. The image size is 256*256 pixels, and we selected a version of the dataset split by the official release, using 20% of the data for training.
MSI: The EuroSAT dataset is a multispectral modality dataset for land use and land cover (LULC) classification. The samples are sourced from the Sentinel-2 optical satellite and include all 13 bands. The data is categorized into 10 classes, with a total of 27,000 images. We adopted a random 9:1 split for training and testing.
HSI: The Pavia University dataset is a high-spatial-resolution hyperspectral dataset acquired by the ROSIS sensor. It comprises 103 bands with a size of 610*340 pixels. The dataset includes nine land cover categories such as Asphalt, Meadows, Gravel, etc. We used a 4:6 split for training and testing.
Table: The PRSA dataset is a collection of air quality data from multiple stations in Beijing, containing hourly measurements of air pollutants. The data spans from March 1, 2013, to February 28, 2017, across 12 monitoring stations. In our experiments, we used various features, including time, station information, 4 air pollutant variables (SO2, NO2, CO, O3), and 6 meteorological variables (temperature, pressure, dew point temperature, amount of precipitation, wind speed, wind direction). The task was to predict the concentration of PM2.5. We split the data into training (40%) and testing sets.
Trajectory: The ETH-UCY dataset is one of the widely used benchmarks for pedestrian trajectory prediction, divided into five subsets: ETH, HOTEL, UNIV, ZARA1, and ZARA2. In our experiments, we utilized the ETH subset.
SAR: The MSTAR dataset is a synthetic aperture radar dataset designed for military stationary target recognition, comprising ten categories of military targets. We employed the standard operating conditions (SOC) dataset preprocessing method proposed by S. Chen et al. chen2016target , ensuring that the serial numbers and target configurations are consistent between the test and training sets, while the aspects and depression angles differ.
Infrared: SYSU-MM01 is a cross-modal human re-identification dataset with optical-infrared modalities, including images from 4 RGB cameras and 2 infrared cameras of 491 different individuals. We adopted the official dataset split, with the training set containing 20,284 RGB images and 9,929 infrared images from 296 individuals, while the test set consists of 3,803 infrared images from 96 individuals used as queries and 301 randomly selected RGB images for re-identification targets.
Graph: METR-LA is a traffic data collection from loop detectors on the highways of Los Angeles, spanning from March 1, 2012, to June 30, 2012. The task is traffic flow prediction.
Oblique Photography: WHU-OMVS is an oblique photography dataset designed for 3D reconstruction tasks. The dataset provides imagery from five different viewpoints, along with camera parameters and other relevant information. It consists of six areas, and in our experiments, we use area 1 as the training set and area 2 as the test set.
Video: UCF101 dataset is a human action recognition dataset comprising 101 action classes with a total of 13,320 video clips. The videos have a combined duration of 27 hours and a resolution of 320*240 pixels, sourced from YouTube.
PointCloud: ModelNet40 is a synthetic point cloud dataset consisting of 40 object categories and a total of 12,311 point cloud objects. We follow the official dataset split, with 9,843 objects used for training and 2,468 for testing.
4.3 Results
4.3.1 RGB
We evaluated the performance of AllSpark on the RGB image scene classification task using the NWPU-RESISC45 dataset, with the Top-1 Accuracy as the evaluation metric. AllSpark leverages expert knowledge from Lynx by loading its pretrained weights. Therefore in Table 4, we compare AllSpark with the state-of-the-art models known on the NWPU-RESISC45 dataset. The results indicate that AllSpark outperforms most baseline models, with a margin of only 0.84 compared to the SOTA (95.69). This highlights AllSpark’s exceptional perception and interpretation capabilities in the RGB modality.
Table 4: RGB scene classification with AllSpark(80% as trainset)
Method
Publication
Acc (%)
CNN-CapsNetzhang2019remote
RS2019
89.03
DFAGCNxu2021deep
TNNLS2021
89.29
D-CNN with GoogleNetcheng2018deep
TGRS2018
90.49
D-CNN with VGGNetcheng2018deep
TGRS2018
91.89
SCCovhe2019skip
TNNLS2019
92.1
SeCo-ResNet-50manas2021seasonal
ICCV2021
92.91
MG-CAPwang2020multi
TIP2020
92.95
LSENetbi2021local
TIP2021
93.34
MSANetzhang2021multiscale
JSTARS2021
93.52
IDCCPwang2020invariant
TGRS2021
93.76
MBLANetchen2021remote
TIP2021
94.66
GRMANet-ResNet-50li2021gated
TGRS2021
94.72
EMSNetzhao2023emscnet
TGRS2023
95.37
ViTAE-B + RVSAwang2022advancing
TGRS2022
95.69
AllSpark
ours
94.85
4.3.2 MSI
We evaluated the performance of AllSpark on the MSI scene classification task using the EuroSAT dataset. In the experiment, all 13 spectral bands of the images were simultaneously input into the model. The model’s objective was to correctly classify the images into one of the 10 specified categories, and the evaluation metric chosen was Top-1 Accuracy. AllSpark does not possess expert knowledge in the multispectral modality, so we categorize the baseline models into two groups in Table 5: those with expert knowledge intervention and those without. Expert knowledge intervention refers to baseline models pretraining on large datasets such as BigEarthNet and then fine-tuning on the EuroSAT dataset, while no expert knowledge indicates baseline models trained directly from scratch on the EuroSAT dataset. The results show that our model outperforms most models in the no expert knowledge group, with a margin of only 2.60 compared to the state-of-the-art model (ResNet-152). Furthermore, AllSpark lags behind the best result in the expert knowledge group by only 4.75, demonstrating its excellent adaptability to the multispectral modality.
Table 5: MSI Land Cover Classification with AllSpark
Expert Knowledge
Method
Publication
Accuracy (%)
No
ResNet-18manas2021seasonal
ICCV2021
63.21
ResNet-50nampally2023performance
IGARSS2023
91.13
InceptionNetzhang2020semi
ICIP2020
93.07
EfficientNetgomez2021msmatch
IJSTARS2021
93.94
AllSpark ours 94.03
ResNet-152nampally2023performance
IGARSS2023
96.63
Yes
MoCoV2manas2021seasonal
ICCV2021
89.51
SeComanas2021seasonal
ICCV2021
93.14
SEERgoyal2022vision
Arxiv2022
97.6
DINO-MCwanyan2023dino
Arxiv2023
98.78
4.3.3 HSI
We conducted pixel classification task on the Pavia University dataset for the hyperspectral modality. The model treats all spectral bands of a single pixel as one sample and predicts the land cover category of that pixel. The reported metrics include Overall Accuracy (OA), Average Accuracy (AA), and Kappa. Since AllSpark does not possess expert knowledge in the hyperspectral modality, we compared it with the semi-supervised baselines summarized by D. Uchaev and D. Uchaev uchaev2023small . The results in Table 6 demonstrate that AllSpark outperforms many hyperspectral image classification methods such as IFRF and S-DMM, with a difference of 6.42 compared to the best result in terms of OA, highlighting AllSpark’s superior adaptability to the hyperspectral modality.
Table 6: HSI pixel Classification with AllSpark
Method
OA
AA
Kappa
3D-CNN
75.24 80.26 68.34
CA-GAN
76.81 76.94 71.02
3D VS-CNN
81.63 83.86 76.46
RPNet
84.92 83.26 80.52
S-DMM
88.3 93.76 84.9
IFRF
88.38 85.99 84.97
AllSpark 89.18 86.65 85.32
DCFSL
90.71 90.2 87.73
TC-GAN
93.2 91.6 91
PRNet-RF
95.6 94.96 94.27
4.3.4 Table
Table 7: PM2.5 prediction with AllSpark
Expert Architecture Method PM2.5
RMSE MAE R2
No
GWOxie2020evolving
62.2 40.8 -
TabBERTpadhi2021tabular
32.8 - -
AllSpark 29.03 18.04 0.87
Yes
Stacked ResNet-LSTMcheng2022stacked
40.68 23.75 0.8
CBAM-CNN-Bi LSTMli2022prediction
18.9 11.2 0.94
For the Table modality, we evaluated AllSpark on the regression prediction task using the PRSAzhang2017cautionary dataset. The task involves predicting the concentration of PM2.5 in the air using features such as time, site, 4 air pollutants, and 6 meteorological variables (as detailed in Section 4.2). The performance metrics include Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and R-squared (R2).
Table 7 presents the comparative results between AllSpark and baselines. It’s worth noting that some works specifically focus on the prediction task on the PRSA dataset and design expert models with specific architectures based on dataset characteristics. For example, the CBAM-CNN-Bi LSTM proposed by D. Li et al.li2022prediction uses CNN to extract spatial dependencies between air monitoring stations and Bi-LSTM to capture the temporal dependencies of PM2.5 data. Similarly, the Stacked ResNet-LSTM model proposed by X. Cheng et al.cheng2022stacked employs a stacking LSTM strategy to enhance the extraction of temporal features in PM2.5 data. Therefore, we categorize baseline methods into two types: with expert structures and without expert structures. In models without expert structures, our approach achieves the best performance among the baselines and is slightly inferior to the state-of-the-art method in models with expert structures (CBAM-CNN-Bi LSTM). This reflects the excellent adaptability of AllSpark to the Table modality.
4.3.5 Code
Table 8: Code Document Generation with AllSpark
Expert Knowledge Method MRR
Ruby Javascript
No
BIRNN[68]
0.084 0.153
1D-CNN[68]
0.245 0.352
selfAtt[68]
0.365 0.451
NBoW[68]
0.429 0.461
RoBERTa[67]
0.625 0.606
AllSpark 0.627 0.635
Yes
RoBERTa(Code)[67]
0.661 0.64
CodeBERT[67]
0.693 0.706
GraphCodeBERT[67]
0.732 0.711
For code modality, we evaluated the performance of AllSpark on the Code Document Generation task using the CodeSearchNet dataset. This task involves generating corresponding documents based on the provided function code. We conducted tests for both Ruby and JavaScript languages, using Mean Reciprocal Rank (MRR)guo2020graphcodebert as the evaluation metric.
Similar to the previous modalities, since MSI-AGI does not possess expert knowledge in the Code modality, we categorized the baselines into two groups: with and without expert knowledge. As shown in Table 8, AllSpark achieved the SOTA results in the group without expert knowledge and remained comparable to models with expert knowledge. This demonstrates the strong adaptability of AllSpark to the Code modality.
4.3.6 PointCloud
For pointcloud modality, we evaluated the performance of AllSpark on the ModelNet40 dataset for the classification task, with Top-1 Accuracy as the metric. In the context of single-modal studies focused on point clouds, we observed that due to the unique three-dimensional structure of point cloud data, most works concentrate on designing specific structures to maintain properties like permutation invariance and symmetry in 3D point clouds. However, these structures designed for the unique priors of the modality are challenging to transfer across modalities. Additionally, some methods tend to pretrain on large point cloud datasets to acquire general modal expert knowledge before generalizing to specific downstream tasks to improve performance.
Table 9: PointCloud classification with AllSpark
Method
Expert Architecture Expert Knowledge Acc(%)
PointNetqi2017pointnet
Yes No 89.2
Kd-netklokov2017escape
Yes No 90.6
SPH3D-GCNlei2020spherical
Yes No 91.4
PointNet++qi2017pointnet++
Yes No 91.9
SO-Netli2018so
Yes No 92.5
PointVGGli2021pointvgg
Yes No 93.6
PointBERTyu2022point
No Yes 93.8
PointGPTchen2023pointgpt
Yes Yes 94.9
AllSpark
No No 91.2
The design of AllSpark is based on a general sequence-to-sequence architecture. As shown in Table 9, in the absence of both modality expert architectural designs and modal expert knowledge, AllSpark still outperforms classical networks with point cloud-specific structures (PointNet, Kd-net). It maintains comparability with the state-of-the-art model, PointGPT, which has both a modal expert structure and modal expert knowledge. This suggests that AllSpark holds significant potential for applications in the point cloud modality.
4.3.7 Trajectory
Table 10: Trajectory predicition with AllSpark
Method
ADE/FDE
Social GAN
0.87/1.62
SoPhie
0.70/1.43
STAR
0.36/0.64
SGCN
0.63/1.03
CAGN
0.41/0.65
SIT
0.39/0.62
SocialVAE
0.47/0.76
PCENet
0.54/0.87
AgentFormer
0.45/0.75
MemoNet
0.40/0.61
SocialVAE+FPC
0.41/0.58
TUTR
0.40/0.61
AllSpark
0.43/0.69
For the trajectory modality, we evaluated the performance of AllSpark on the ETH dataset in the trajectory prediction task. This task involves predicting possible two-dimensional trajectories based on a set of two-dimensional coordinate points within a certain time period. We report accuracy using Average Displacement Error (ADE) and Final Displacement Error (FDE)shi2023trajectory . Given the future trajectory$\{x_{t},y_{t}\}_{t=T_{bos}+1}^{T}$ (Ground Truth) and the predicted trajectory $\{\widehat{x_{t}},\widehat{y_{t}}\}_{t=T_{bos}+1}^{T}$, ADE and FDE are used to measure their $L2$ distances, calculated as follows:
${ADE}=\frac{1}{T_{pred}}\sum_{t=T_{bos}}^{T}\sqrt{\left(x_{t}-\widehat{x_{t}}\right)^{2}+\left(y_{t}-\widehat{y_{t}}\right)^{2}}$ (5)
${FDE}=\sqrt{\left(x_{T}-\widehat{x_{T}}\right)^{2}+\left(y_{T}-\widehat{y_{T}}\right)^{2}}$ (6)
In Table 10, AllSpark is compared with the state-of-the-art trajectory prediction models. The results indicate that AllSpark, using a unified structure without trajectory modality expert knowledge, outperforms most expert models. It achieves a prediction accuracy close to the SOTA model (STAR) with a difference of only 0.07 in the ADE metric and 0.11 in the FDE metric. This suggests that AllSpark demonstrates excellent adaptability to the trajectory modality.
4.3.8 SAR
Table 11: SAR classification with AllSpark
Method
Accuracy (%)
EMACHsrinivas2014sar
88
SVMsrinivas2014sar
90
AdaBoostsrinivas2014sar
92
MSRCdong2014sparse
93.6
IGTsrinivas2014sar
95
MSSdong2015classification
96.6
Cond Gausso2001sar
97
M-PMCpark2014modified
98.8
AConvNetschen2016target
99.13
AllSpark
88
For SAR modality, AllSpark’s adaptability was tested on the MSTAR dataset, where the model is required to identify SAR images of ten military targets, and the metric used is Top1 Accuracy. In the experiment, the preprocessing of the MSTAR dataset followed the SOC (standard operating conditions) settings from AConvNetschen2016target . Table 11 presents the comparison results between AllSpark and the state-of-the-art model under these settings. Since AllSpark lacks SAR image modality expert knowledge and the spatial invariance prior to changes in perspective, its performance on the MSTAR dataset is suboptimal, with an accuracy of 88%. This is comparable to the EMACH method. Enhancing the model’s adaptability to SAR images and improving its ability to extract spatial information are directions for our future research.
4.3.9 Infrared
AllSpark’s adaptability to the infrared modality was tested on the SYSU-MM01 dataset in a cross-modal human re-identification task involving RGB and infrared modalities. In this experiment, the model receives both an infrared image and an RGB image and is tasked with identifying and matching the same person across different modality images. Table 12 presents the comparison results between AllSpark and the state-of-the-art re-identification methods, with the evaluation metric being Rank at 20 accuracy. The results indicate that AllSpark (76.31%) achieved similar accuracy compared to HSME (77.95%). While our method did not reach the highest accuracy, it has already surpassed some expert models in RGB-Infrared re-identification (e.g., TONE, HCML) when using a unified architecture without modal expert knowledge, showcasing the potential of AllSpark in handling the infrared modality.
Table 12: Visible-Infrared Re-identification with AllSpark
Method
Rank-20(All-search)
Two-streamwu2017rgb
65.5
One-streamwu2017rgb
66.74
TONEye2018hierarchical
68.6
HCMLye2018hierarchical
69.17
Zero-Padwu2017rgb
71.33
AllSpark 76.31
HSMEhao2019hsme
77.95
BDTRye2019bi
81.07
DDAGye2020dynamic
95.81
4.3.10 Graph
For the Graph modality, AllSpark’s performance was evaluated on the traffic flow prediction task using the METR-LA dataset. The evaluation metrics include RMSE, MAE, and R2. Table 13 presents the comparison between AllSpark and the state-of-the-art results on the METR-LA dataset, with the baseline derived fromliu2023spatio . The results show that AllSpark, without the intervention of modal expert knowledge, is only 0.47 away from the best result in terms of RMSE, demonstrating AllSpark’s excellent adaptability to the Graph modality.
Table 13: Traffic prediction with AllSpark
Method
RMSE MAE R2
HI
6.8 14.2 10.15
GWNet
3.51 7.28 9.96
DCRNN
3.54 7.47 10.32
AGCRN
3.59 7.45 10.47
STGCN
3.6 7.43 10.35
GTS
3.59 7.44 10.25
MTGNN
3.47 7.21 9.7
STNorm
3.57 7.51 10.24
GMAN
3.44 7.35 10.07
PDFormer
3.62 7.47 10.91
STID
3.55 7.55 10.95
STAEformer
3.34 7.02 9.7
AllSpark
3.81 7.52 11.24
4.3.11 Language
Table 14: Text understanding with AllSpark
Method
Accuracy (%)
RoBERTaliu2019roberta
95.3
ULMFiThoward2018universal
95.4
BERTxie2020unsupervised
95.49
Mixed VATsachan2019revisiting ; miyato2016adversarial
95.68
LongFormerbeltagy2020longformer
95.7
XLNetyang2019xlnet
96.8
ERNIE-Doc-Largeding2020ernie
97.1
AllSpark
96.78
We tested AllSpark’s natural language processing capabilities on the IMDB datasetmaas2011learning with the task of binary sentiment classification (positive or negative). AllSpark loaded weights from the Lynx, and therefore, it can be considered to possess expert knowledge in the natural language modality. When compared with the state-of-the-art models on the IMDB dataset, as shown in Table 14, AllSpark outperformed most language models, with only a 0.32 difference from the SOTA result. This highlights AllSpark’s powerful understanding and analysis capabilities in natural language.
4.3.12 Video
For the Video modality, we evaluated AllSpark’s performance on action recognition tasks using the UCF101 dataset. The model is tasked with understanding videos and accurately classifying them into one of the 101 classes, with the evaluation metric being Top1 accuracy. In Table 15, we compared AllSpark with the current state-of-the-art models. Currently, AllSpark’s adaptation to the Video modality is not optimal, showing a significant gap compared to baseline model results. We attribute this to two main reasons: 1. The high redundancy in video information increases the training cost for AllSpark, and we trained it for only 3 epochs on the UCF101 dataset. 2. AllSpark’s model architecture lacks flexibility for three-dimensional data, making it less effective in capturing temporal information.
Table 15: Video classification with AllSpark
Method
Accuracy (%)
OPN
59.6
VCOP
72.4
SpeedNet
81.1
VTHCL
82.1
CVRL
94.4
VideoMAE v1
96.1
VideoMAE v2
99.6
AllSpark
27.5
4.3.13 Oblique Photography
For the Oblique Photography modality, we tested AllSpark’s performance on the 3D reconstruction task using the WHU-OMVS dataset. The model takes five-view images as input, and the goal is to output depth maps for reconstructing 3D models. The evaluation metric used is the Percentage of Accurate Grids in Total (PAG)liu2023deep , calculated by the following formula:
$PAG_{a}=\left(\frac{m_{a}}{m}\right)*100$ (7)
The suffixes in PAG represent different accuracy standards, where ${PAG}_{6}$ signifies an error within 0.6 meters, and ${PAG}_{10}$ indicates an error within 1 meter. Table 16 presents the comparison of AllSpark with popular multi-view 3D reconstruction models on the WHU-OMVS dataset. AllSpark falls short in terms of accuracy compared to modality-specific expert models. We speculate two possible reasons: 1. Dense spatial information gradually diminishes in the deep structure of large models, a point verified in our exploratory experiments involving segmentation, detection, etc.; 2. The model architecture lacks flexibility and cannot connect gradual features like expert 3D reconstruction models such as Ada-MVS. Additionally, it lacks the ability to design specific model structures for processing, severely restricting its performance. Adapting multi-modal large models to dense prediction tasks and optimizing the architecture is one of our future research directions.
Table 16: 3D reconstruction with AllSpark
Method
${PAG}_{6}$
${PAG}_{10}$
MVSNet
81.15 91.44
CasMVSNet
95.45 98.02
Ada-MVS
96.14 98.1
UCSNet
96.25 98.45
AllSpark
6.4 10.4
5 Discussion
5.1 Limitation
Certainly, our work has some limitations, and these will guide our future research directions:
  1. 1.
    Lack of interaction between different modalities. While AllSpark incorporates interaction between infrared and RGB modalities, for other modalities, it essentially uses a unified model for multiple single-modal tasks. This is primarily due to constraints such as the lack of multi-modal paired data, the inflexibility of the model architecture, and the limited richness of multi-modal interaction tasks. In the future, we plan to explore and work more extensively in these directions.
  2. 2.
    Lack of in-depth research on multi-modal alignment methods. While we introduced the basic principles of the Language as Reference Framework (LaRF) and aligned all modalities to the language modality, our approach included only introducing the modality bridge from Lynx as the alignment method. We haven’t conducted in-depth research into the various methods of multi-modal alignment. Moreover, our testing focused on the model’s adaptability to different modalities, and we haven’t explored potential capabilities such as interactive reasoning, zero-shot learning, and multi-modal emergence that the model may possess after alignment with language. In the future, we plan to delve deeper into researching multi-modal alignment methods, and we will design additional tasks and experiments to explore and reveal the upper limits and boundaries of large model reasoning capabilities.
  3. 3.
    Initial work. Our current work is still in its initial exploratory phase, and we haven’t carefully refined AllSpark’s adaptability and performance on each modality. As a result, the model exhibits suboptimal performance on certain modalities, such as oblique photography, SAR, and video. In the future, we plan to conduct more refined and targeted adjustments for each modality to enhance overall performance.
  4. 4.
    Expensive Cost: Multimodal large language models, due to prolonged pretraining, typically possess universal reasoning capabilities in certain modalities. We generalize their applicability to other modalities by utilizing modality bridges to project other modalities onto the language modality. Although we freeze most parameters, fine-tuning even once on 2 A6000 GPUs often requires more than a day. Given the increasing training costs for large models, exploring methods to generalize their universal reasoning abilities remains one of our future research directions.
  5. 5.
    Interesting phenomenon. During our exploratory experiments, we discovered some interesting properties of large models. For instance, spatial information tends to degrade in large models, leading to collapse when performing dense prediction tasks such as segmentation and detection. Additionally, these models struggle to optimize when transferred to a small amount of downstream data. Large models often require smaller hyperparameters and are sensitive to them. These observations might partially reveal the working mechanisms of large models, and we plan to conduct more in-depth investigations into these phenomena in the future.
5.2 Potential of the LaRF
Inspired by the human cognitive system and linguistic philosophy, we propose ”Language as Reference Framework (LaRF)” as the first principle for constructing our unified multimodal model. Its foreseeable potential includes at least the following three points:
  1. 1.
    Efficient Generalization of Large Models. Currently, the computational power and data scale required for training large models are rapidly expanding, and even the cost of fine-tuning large models is becoming prohibitive. Therefore, in the future, training large models for every domain is almost an impossible task. Language, however, holds the potential for achieving efficient generalization of large models, expanding them from their native domains to additional domains at minimal cost. In our proposed AllSpark, we have designed simple text prompts and task heads for each modality, demonstrating significant potential for multimodal expansion. In theory, AllSpark can extend to any number of modalities. In the future, we will conduct more in-depth research on the impact of text prompts and lightweight parameter modules on the generalization of large models.
  2. 2.
    Interpretable Reasoning. Deep learning models have often been referred to as ”black boxes,” indicating that the reasoning process of these models is invisible and challenging to interpret. Some research on the interpretability of deep learning models is often based on complex mathematical models and numerous assumptions, greatly limiting the practical applications of deep learning methods in fields such as clinical medicine, military, and national resources where low fault tolerance or high confidentiality is crucial. However, language, as a tool for human thought and communication, provides models based on LaRF with the potential to directly use natural language for outputting reasoning chains and justifications.
  3. 3.
    Transition from end-to-end to interactive paradigm. The end-to-end paradigm refers to the learning approach where the model takes input and directly outputs results. In recent years, the end-to-end paradigm has become increasingly popular due to its simple and clear architecture and excellent performance. However, it also has clear disadvantages, such as uncontrollable internal operations, the need to optimize the whole for certain problems, and difficulty in pinpointing the cause of issues. A LaRF-based architecture has the potential to achieve an interactive paradigm, where users input raw data and corresponding text prompts, and the model automatically performs relevant operations based on the prompts. Users can even iteratively adjust the text prompts based on the results. Therefore, in terms of both performance and controllability, the interactive paradigm has advantages that are incomparable to the end-to-end paradigm.
6 Conclusion
Leveraging multimodal data is an inherent requirement for intelligent models to achieve geographic object cognition. Inspired by human cognitive systems and linguistic philosophy, we propose that the construction of multimodal models should follow the fundamental principle of Language as Reference Framework (LaRF). Guided by this principle, we use language to balance the cohesion and autonomy of modalities, presenting a unified intelligent model, AllSpark, encompassing 13 spatiotemporal modalities. Experimental results demonstrate that AllSpark exhibits excellent adaptability and application potential across various spatiotemporal modalities, highlighting the feasibility and potential of constructing multimodal models with LaRF. AllSpark remains an initial exploratory work, and in the future, we aim to delve deeper into mechanisms guided by natural language, efficient generalization of large models, and the transition to an interactive paradigm.

### v2

@@
-AllSpark: a multimodal spatiotemporal general model
+AllSpark[Uncaptioned image]: A Multimodal Spatio-Temporal General Intelligence Model with Thirteen Modalities
 Abstract
-RGB, infrared, multispectral, and other modal data fundamentally represent different observational approaches to the same geographic object.
+RGB, infrared, multispectral, and other spatio-temporal modal data fundamentally represent different observational approaches for the same geographic object.
 Therefore, leveraging multimodal data is an inherent requirement for comprehending geographic objects.
-However, for a long time, due to the high heterogeneity in structure and semantics among various spatiotemporal modal data, the joint interpretation of multimodal spatiotemporal data has been an extremely challenging problem.
-The primary challenge resides in striking a trade-off between the cohesion and autonomy of diverse modalities, and this trade-off exhibits a progressively nonlinear nature as the number of modalities expands.
-Inspired by the human cognitive system and linguistic philosophy, where perceptual signals from the five senses converge into language, we introduce the Language as Reference Framework (LaRF), a fundamental principle for constructing a multimodal unified model, aiming to strike a trade-off between the cohesion and autonomy among different modalities.
-Building upon this, we propose a multimodal spatiotemporal general artificial intelligence model, called AllSpark.
+However, due to the high heterogeneity in structure and semantics among various spatio-temporal modal data, the joint interpretation of multimodal spatio-temporal data has long been an extremely challenging problem.
+The primary challenge resides in striking a trade-off between the cohesion and autonomy of diverse modalities.
+This trade-off becomes progressively nonlinear as the number of modalities expands.
+Inspired by the human cognitive system and linguistic philosophy, where perceptual signals from the five senses converge into language, we introduce the Language as Reference Framework (LaRF), a fundamental principle for constructing a multimodal unified model.
+Building upon this, we propose a multimodal spatio-temporal general artificial intelligence model called AllSpark.
 Our model integrates thirteen different modalities into a unified framework, including one-dimensional (text, code), two-dimensional (RGB, infrared, SAR, multispectral, hyperspectral, tables, graphs, trajectory, oblique photography), and three-dimensional (point clouds, videos) modalities.
-To achieve modal cohesion, AllSpark uniformly maps diverse modal features to the language modality.
-In addition, we design modality-specific prompts to guide multi-modal large language models in accurately perceiving multimodal data.
-To maintain modality autonomy, AllSpark introduces modality-specific encoders to extract the tokens of various spatiotemporal modalities.
-And modal bridge is employed to achieve dimensional projection from each modality to the language modality.
-Finally, observing a gap between the model’s interpretation and downstream tasks, we designed task heads to enhance the model’s generalization capability on specific downstream tasks.
-Experiments indicate that AllSpark, without expert knowledge of the most spatiotemporal modalities and utilizing a unified structure, achieves competitive accuracy in modalities such as RGB and trajectory compared to state-of-the-art models.
-Moreover, AllSpark showcases excellent adaptability in modalities like MSI, HSI, PointCloud, Table, Code, and Graph.
+To achieve modal cohesion, AllSpark uniformly maps diverse modal features to the language modality by a modal bridge and a multimodal large language models.
+To maintain modality autonomy, AllSpark introduces modality-specific encoders to extract the tokens of various spatio-temporal modalities.
+Finally, observing a gap between the model‘s interpretability and downstream tasks, we designed modality-specific prompts and task heads to enhance the model’s generalization capability on specific downstream tasks.
+Experiments indicate that AllSpark, without expert knowledge of most spatio-temporal modalities and utilizing a unified structure, achieves competitive accuracy in modalities such as RGB and trajectory compared to state-of-the-art models.
+Moreover, AllSpark shows excellent adaptability in various modalities, such as MSI, HSI, PointCloud, table, code, and graph.
 AllSpark demonstrates the potential and possibility of constructing general artificial intelligence with a large language model.
-This approach contributes to the shift in research paradigm in spatiotemporal intelligence, transitioning from a modality-specific and task-specific paradigm to a general paradigm.
-keywords:
-Multimodal, Language, Spatiotemporal, General Model
-1 Introduction
-Benefiting from the increasingly diverse observational methods for spatiotemporal scenes, geographic objects can be described using various spatiotemporal modalities such as RGB, infrared, SAR, multispectral, graphs, and spatiotemporal trajectories.
+LaRF and AllSpark contribute to the shift in the research paradigm in spatio-temporal intelligence, transitioning from a modality-specific and task-specific paradigm to a general paradigm.
+Index Terms:
+Spatio-temporal Data, Multimodal Machine Learning, Large Language Model, General Intelligence Model.
+I Introduction
+Figure 1: We propose ”Language as Reference Framework (LaRF)” as the fundamental principle for constructing multimodal models.
+Under the guidance of this principle, we introduced a multimodal general intelligence model, AllSpark, which integrates thirteen spatio-temporal modalities.
+Benefiting from the increasingly diverse observational methods available for spatio-temporal scenes, geographic objects can be described using various spatio-temporal modalities, such as RGB, infrared, SAR, multispectral, graph, and spatio-temporal trajectory data.
 Each modality provides unique information about different aspects of the geographic object.
-Analogous to the human process of perceiving and understanding the world through multiple modalities such as vision, hearing, and touch, joint interpretation of multimodal data is an inherent requirement for intelligent models to achieve cognition of geographic objects.
+Analogous to the human process of perceiving and understanding the world through multiple modalities, such as vision, hearing, and touch, joint interpretation of multimodal data is an inherent requirement for intelligent models to achieve cognition of geographic objects.
 However, due to the inherent differences in the mechanisms of each modality, various modalities are often highly heterogeneous in both structure and semantics.
@@
 Point clouds are represented by three-dimensional coordinates along with various feature values.
-Text is constituted by sequences of words.
+A text is composed of sequences of words.
 In terms of semantics, RGB imagery reflects the electromagnetic characteristics of visible light bands emitted and reflected by geographic objects, whereas SAR imagery reflects the electromagnetic characteristics of microwaves emitted actively by radar for scattering from geographic objects.
-For a long time, constrained by the highly heterogeneity across various modalities as mentioned above, researchers have often developed specific methods based on prior assumptions related to a particular modality or designed multimodal approaches based on a few low-heterogeneity modalities.
+For a long time, constrained by the high heterogeneity across various modalities mentioned above, researchers have often developed specific methods based on prior assumptions related to a particular modality or designed multimodal approaches based on a few low-heterogeneity modalities.
 For instance, in single-modal research, C.
@@
 Qi et al.
-proposed PointNetqi2017pointnet for the point cloud modality, emphasizing the invariance of point cloud data ordering and the significance of global and local features.
-For the language modality, Vaswani et al.
-introduced Transformervaswani2017attention , focusing on the long-range dependencies within word sequences.
-For the graph modality, Kipf and Welling proposed GCNkipf2016semi , based on the adjacency relationships between nodes in a graph.
-In multimodal research, fusion of optical and SAR imagery has been widely explored in both traditional and deep learning remote sensingkulkarni2020pixel .
-Text-visual models have also experienced rapid development in recent yearsradford2021learning ; lu2019vilbert ; kim2021vilt ; alayrac2022flamingo ; li2022blip .
-The diverse prior assumption associated with each modality have resulted in significant gaps between methods designed for different modalities, making it challenging to perceive and understand different modalities using a unified model.
+proposed PointNet[1] for the point cloud modality, emphasizing the invariance of point cloud data ordering and the significance of global and local features.
+For language modality, Vaswani et al.
+introduced the transformer[2], which focuses on the long-range dependencies within word sequences.
+For the graph modality, Kipf and Welling proposed the GCN[3] based on the adjacency relationships between nodes in a graph.
+In multimodal research, the fusion of optical and SAR imagery has been widely explored in both traditional and deep learning remote sensing[4].
+Moreover, visual-language models have undergone rapid development in recent years[5, 6, 7, 8, 9].
+The diverse prior assumptions associated with each modality have resulted in significant gaps between methods designed for different modalities, making it challenging to perceive and understand different modalities using a unified model.
 We believe that the key challenge in addressing this issue lies in striking a trade-off between the cohesion and autonomy of diverse modalities.
 Cohesion refers to the presence of mutually correlated shared information among modalities, while autonomy signifies the existence of unique information specific to each modality relative to others.
-If we simply project data from multiple modalities into a common representation space for fusion, emphasizing the cohesion between modalities, it may lead to the loss of modality-specific information, thereby diluting the essence of multimodal collaboration.
-On the contrary, if we excessively stress the autonomy between modalities, it may hinder the establishment of connections among them, limiting the model’s ability to simultaneously perceive multiple modalities.
-Moreover, as the number of modalities increases, balancing the cohesion and autonomy becomes progressively more challenging in a nonlinear manner.
-We observe that in the process of comprehending the world, humans integrate information from multiple modalities such as hearing, touch, smell, and vision.
-The concepts formed through the parsing of these modalities are ultimately converged in language.
-It is through language that humans engage in associating, reasoning, and expressive behaviors.
+If we simply project data from multiple modalities into a common representation space for fusion, emphasizing the cohesion between modalities, this approach may lead to the loss of modality-specific information, thereby diluting the essence of multimodal collaboration.
+In contrast, if we excessively stress the autonomy between modalities, it may hinder the establishment of connections among them, limiting the model’s ability to simultaneously perceive multiple modalities.
+Moreover, as the number of modalities increases, balancing cohesion and autonomy becomes progressively more challenging nonlinearly.
+We observe that in the process of comprehending the world, humans integrate information from multiple modalities, such as hearing, touch, smell, and vision.
+The concepts formed through the parsing of these modalities ultimately converge in language.
+Humans engage in associating, reasoning, and expressive behaviours through language.
 In other words, language precisely encodes human perception and understanding of the world, providing clear definitions and meanings to abstract concepts from each modality.
-Inspired by this, we propose Language as Reference Framework (LaRF) as a fundamental principle for constructing multimodal models.
+Inspired by this, we propose the Language as Reference Framework (LaRF) as a fundamental principle for constructing multimodal models.
 As illustrated in Figure 1, the abstract concepts derived from each modality should align with language, allowing for joint interpretation in the unified representation space of language.
-Figure 1: Language as Reference Framework (LaRF)
-Building upon this, we propose a multimodal spatiotemporal general artificial intelligence model, AllSpark, integrating thirteen different modalities into a unified framework, including one-dimensional (text, code), two-dimensional (RGB, infrared, SAR, multispectral, hyperspectral, tables, graphs, trajectory, oblique photography), and three-dimensional (point clouds, videos) modalities.
+Building upon this, we propose a multimodal spatio-temporal general artificial intelligence model, AllSpark, that integrates thirteen different modalities into a unified framework, including one-dimensional (text, code), two-dimensional (RGB, infrared, SAR, multispectral, hyperspectral, tables, graphs, trajectory, oblique photography), and three-dimensional (point clouds, videos) modalities.
 To achieve modal cohesion, AllSpark uniformly maps diverse modal features to the language modality.
-In addition, we design modality-specific prompts to guide multi-modal large language models in accurately perceiving multimodal data.
-To maintain the autonomy between modalities, AllSpark introduce specific modal encoders for each modality to extract independent tokens.
-Given the high heterogeneity among modality data and modality encoders, there exists significant dimensional gap between the tokens of each modality and the language modality.
-To address this issue, we introduce modality bridge, a mechanism from Lynxzeng2023matters to accomplish the dimensional mapping from each modality’s tokens to the language tokens.
-Finally, considering the existing gap between the interpretation by the multimodal large language model and the specific downstream tasks, we have designed task heads for each downstream task to enhance the model’s generalization capability.
-It is noteworthy that, given the powerful interpretation capabilities of the multimodal large language model, we adhere to a lightweight design principle in both the modal encoders and task heads
-Experiments demonstrate that AllSpark, without the expert knowledge in most modalities and utilizing a unified structure, achieves competitive accuracy in modalities such as RGB and spatiotemporal trajectories compared to state-of-the-art models.
-Specifically, in the RGB modality, AllSpark’s accuracy is only 0.84 lower than the optimal model, and in the trajectory modality, the ADE metric differs by only 0.07 compared to the optimal model.
+In addition, we designed modality-specific prompts to guide multimodal large language models in accurately perceiving multimodal data.
+To maintain the autonomy between modalities, AllSpark introduces specific modal encoders for each modality to extract independent tokens.
+Given the high heterogeneity among modality data and modality encoders, a significant dimensional gap exists between the tokens of each modality and the language modality.
+To address this issue, we introduce a modality bridge, a mechanism from Lynx[10], to accomplish dimensional mapping from each modality’s tokens to the language tokens.
+Finally, considering the existing gap between the interpretability of the multimodal large language model and the specific downstream tasks, we design task heads for each downstream task to enhance the model’s generalization capability.
+Given the powerful interpretability capabilities of the multimodal large language model, we adhere to a lightweight design principle in both the modal encoders and task heads.
+Experiments demonstrate that AllSpark achieves competitive accuracy without expert knowledge in most modalities and utilizing a unified structure in modalities such as RGB and spatio-temporal trajectories compared to state-of-the-art models.
+Specifically, in the RGB modality, the accuracy of AllSpark is only 0.84 lower than that of the SOTA model, and in the trajectory modality, the average displacement error (ADE) metric differs by only 0.07 compared to that of the SOTA model.
 Additionally, AllSpark exhibits excellent adaptability in various other modalities, including point clouds, multispectral, hyperspectral, tables, graphs, and code.
-Theoretically, our proposed model has the potential for seamless extension to arbitrary number of modalities.
-In a word, our contributions can be summarized as follows:
-  • 1.
-    To our best knowledge, we first propose a unified multimodal spatiotemporal general model, AllSpark, successfully integrating 13 spatiotemporal modalities into a single model.
-  • 2.
-    Inspired by the human cognitive system and linguistic philosophy, we propose the Language as Reference Framework (LaRF), offering a novel solution to balance cohesion and autonomy among multiple modalities.
-This principle provides a fresh approach to the multimodal unification.
-  • 3.
-    We provide new possibilities for the paradigm of designing multimodal models in the spatiotemporal domain.
-MSTAGI demonstrates the potential and capabilities of building general model with large language models, thereby contributing to the shift in research paradigm in spatiotemporal intelligence, transitioning from a modality-specific and task-specific paradigm to a general paradigm.
-  • 4.
-    Experiments indicate that AllSpark, without expert knowledge of the majority of spatiotemporal modalities and utilizing a unified structure, achieves competitive accuracy in modalities such as RGB and trajectory compared to state-of-the-art models.
-Moreover, AllSpark showcases excellent adaptability in modalities like MSI, HSI, PointCloud, Table, Code, and Graph.
-Theoretically, our proposed model has the potential for seamless extension to arbitrary number of modalities.
-2 Related Work
+Theoretically, our proposed model has the potential for seamless extension to an arbitrary number of modalities.
+In other words, our contributions can be summarized as follows:
+  • We first propose a unified multimodal spatio-temporal general model, AllSpark, that successfully integrates thirteen spatio-temporal modalities into a single model.
+  • Inspired by the human cognitive system and linguistic philosophy, we propose the language as reference framework (LaRF), which offers a novel solution to balance cohesion and autonomy among multiple modalities.
+This principle provides a fresh approach to multimodal unification.
+  • We propose new approaches for designing multimodal models in the spatio-temporal domain.
+AllSpark demonstrates the potential and ability to construct a general model with large language models, thereby contributing to the shift in the research paradigm in spatio-temporal intelligence, transitioning from a modality-specific and task-specific paradigm to a general paradigm.
+  • Experiments indicate that AllSpark, without expert knowledge of the majority of spatio-temporal modalities and utilizing a unified structure, achieves competitive accuracy in modalities such as RGB and trajectory information compared to state-of-the-art models.
+Moreover, AllSpark shows excellent adaptability in various modalities, such as MSI, HSI, PointCloud, table, code, and graph.
+Theoretically, our proposed model has the potential for seamless extension to an arbitrary number of modalities.
+II Related Work
 Leveraging multimodal data is an inherent requirement for achieving cognitive recognition of geospatial objects.
-An ideal multimodal model should possess the capability to integrate all modalities for joint interpretation.
-Hence, a crucial trend in the research of intelligent methods in the spatiotemporal domain is the continual increase in the number of modalities for joint interpretation.
+An ideal multimodal model should possess the capability to integrate all the modalities for joint interpretation.
+Hence, a crucial trend in the research of intelligent methods in the spatio-temporal domain is the continual increase in the number of modalities available for joint interpretation.
 Initially, early researchers often constructed single-modal expert models based on prior assumptions about a specific modality, achieving remarkable success within each respective modality.
-In recent years, with a deeper understanding of single-modal interpretation methods, numerous researchers have attempted to integrate several low-heterogeneity modalities to build multimodal interpretation approaches.
+In recent years, with a deeper understanding of single-modal interpretation methods, numerous researchers have attempted to integrate several low-heterogeneity modalities to construct multimodal interpretation approaches.
 However, as the number of modalities increases, the challenge of balancing cohesion and autonomy among the modalities becomes increasingly difficult.
-In the following, we recall the development of intelligent methods in the spatiotemporal domain from the perspective of the continually increasing number of modalities, and finally, we present the principles and approach of our proposed model, AllSpark.
-2.1 Single-Modal Model
+In the following, we recall the development of intelligent methods in the spatio-temporal domain from the perspective of the continually increasing number of modalities, and finally, we present the principles and approach of our proposed model, AllSpark.
+II-A Single-Modal Model
 For one-dimensional modalities, we focus on code and text.
-Code and Text are fundamentally language modalities, but due to the excellent characteristics of Code, such as strict syntax, unambiguous nature, and its ability to interact with machines, we treat it as a separate modality.
+Code and text are both fundamental language modalities, but due to the excellent characteristics of code, such as strict syntax, unambiguous nature, and ability to interact with machines, code is treated as separate modality.
 Z.
 Feng et al.
-have pre-trained a model, CodeBERTfeng2020codebert , which facilitates the mutual transfer of information between Code and natural language modalities.
-For the Text modality, landmark contributions include Transformervaswani2017attention ,BERTkenton2019bert ,GPT series radford2018improving ; radford2019language ; brown2020language ; openai2023gpt4 , which have inspired a subsequent series of works.
-We categorize RGB, multispectral, hyperspectral, infrared, SAR, Table, Graph, Trajectory, and Oblique Photography as two-dimensional modalities.
-Among the two-dimensional modalities, the standard three-channel RGB image is one of the most common.
-For the RGB modality, the ResNethe2016deep proposed by K.
-He et al., based on the importance of visual global and local information, and the ViTdosovitskiy2020image introduced by A.
-Dosovitskiy et al., leveraging a global attention mechanism, represent two landmark contributions.
-An increase in the number of channels in images leads to the multispectral and hyperspectral modalities.
+pretrained a model, CodeBERT[11], which facilitates the mutual transfer of information between code and natural language modalities.
+For the text modality, landmark contributions include a transformer[2], BERT[12], and the GPT series [13, 14, 15, 16], which have inspired subsequent series of works.
+We categorize RGB, multispectral, hyperspectral, infrared, SAR, table, graph, trajectory, and oblique photography as two-dimensional (2D) modalities.
+Among the two-dimensional modalities, standard three-channel RGB images are among the most common.
+For the RGB modality, the ResNet[17] proposed by K.
+He et al., which is based on the importance of visual global and local information, and the visual transformer (ViT)[18] introduced by A.
+Dosovitskiy et al., which leverages a global attention mechanism, represent two landmark contributions.
+An increase in the number of channels in images leads to multispectral and hyperspectral modalities.
 B.
 Huang et al.
-proposed the STDCNNhuang2018urban , leveraging the characteristic of a higher number of bands in multispectral images to simultaneously model the global spatial and spectral properties of multispectral images.
+proposed the STDCNN[19], leveraging the characteristic of a greater number of bands in multispectral images to simultaneously model the global spatial and spectral properties of multispectral images.
 In comparison to multispectral images, hyperspectral images have even more bands, often reaching hundreds.
@@
 Yang et al.
-introduced the R-3D-CNNyang2018hyperspectral to further enhance the extraction of spectral features.
-Beyond an increase in the number of bands, attention has been directed towards images captured in specific spectral bands.
+introduced the R-3D-CNN[20] to further enhance the extraction of spectral features.
+Beyond an increase in the number of bands, attention has been given to images captured in specific spectral bands.
 For instance, infrared images, widely applied in fields such as agriculture and oceanography, have gained significant attention.
@@
 Chen et al.
-proposed a shallow convolutional neural network, integrating a decision tree, to meet the demand for real-time intelligent interpretation of a large volume of infrared images in agricultural water resource monitoringchen2020deep .
-In the case of SAR images, formed by active microwave radar, S.
+proposed a shallow convolutional neural network that integrates a decision tree to meet the demand for real-time intelligent interpretation of a large volume of infrared images in agricultural water resource monitoring[21].
+In the case of SAR images formed by active microwave radar, S.
 Chen et al.
-introduced AConvNetchen2016target , a widely used fully convolutional neural network for SAR image intelligent interpretation.
-For oblique photography, 3D reconstruction aims to utilize single or multiple two-dimensional images from different angles to reconstruct a three-dimensional model.
+introduced AConvNet[22], a widely used fully convolutional neural network for intelligent SAR image interpretation.
+For oblique photography, 3D reconstruction aims to utilize single or multiple two-dimensional images from various viewpoints to reconstruct a three-dimensional model.
 Y.
 Yao et al.
-addressed this task by proposing MVSNetyao2018mvsnet .
-Utilizing techniques such as three-dimensional convolution, the model establishes correlations between images from different perspectives, thereby reconstructing the three-dimensional model of the target from the features extracted from two-dimensional images.
-Beyond the visual domain, Table is one of the commonly used modalities for recording and expressing information.
-TabNetarik2021tabnet , proposed by S.
+addressed this task by proposing MVSNet[23].
+By utilizing techniques such as three-dimensional convolution, the model establishes correlations between images from different perspectives, thereby reconstructing a three-dimensional model of the target from the features extracted from two-dimensional images.
+Beyond the visual domain, the table is one of the commonly used modalities for recording and expressing information.
+TabNet[24], proposed by S.
 Ö.
@@
 Gupta et al.
-proposed Social GANgupta2018social based on the characteristic of trajectory multiplicity.
-This model combines historical trajectory information with social context to predict multiple plausible future outcomes.
-For the Graph modality, Kipf and Welling introduced the classic Graph Convolutional Network (GCN)kipf2016semi , which is based on the adjacency relationships between nodes in the graph.
+proposed the social generative adversarial networks (GAN)[25] based on the characteristic of trajectory multiplicity.
+This model combines historical trajectory information with social context information to predict multiple plausible future outcomes.
+For the graph modality, Kipf and Welling introduced the classic graph convolutional network (GCN)[3], which is based on the adjacency relationships between nodes in the graph.
 P.
-Velicˇkovic et al.
-proposed GATvelivckovic2017graph , which incorporates attention mechanisms into the Graph modality.
-Finally, we turn our attention to the three-dimensional modalities: Video and Point Cloud.
+Veličković et al.
+proposed the GAT[26], which incorporates attention mechanisms into the graph modality.
+Finally, we turn our attention to three-dimensional modalities: video and point clouds.
 The video modality is formed by stacking images over time.
@@
 extends the connectivity of CNN models over time.
-Additionally, it utilizes multi-resolution techniques to accelerate the training process, effectively reducing the impact of spatiotemporal redundancy on model efficiency in the Video modalitykarpathy2014large .
-The Point Cloud modality captures information about the position, shape, color, texture, and other aspects of three-dimensional objects.
+Additionally, multiresolution techniques are utilized to accelerate the training process, effectively reducing the impact of spatio-temporal redundancy on model efficiency in the video modality[27].
+The point cloud modality captures information about the position, shape, colour, texture, and other aspects of three-dimensional objects.
 C.
@@
 Qi et al.
-introduced the classic PointNetqi2017pointnet for the Point Cloud modality, emphasizing the importance of invariance through point data permutation and the significance of global and local features.
+introduced the classic PointNet[1] for the point cloud modality, emphasizing the importance of invariance through point data permutation and the significance of global and local features.
 W.
 Wu et al.
-extended convolution operations to three-dimensional point clouds with the introduction of PointConvwu2019pointconv .
-2.2 Multi-modal Model
+extended convolution operations to three-dimensional point clouds with the introduction of PointConv[28].
+II-B Multimodal Model
 While the single-modal methods in Section 2.1 have demonstrated excellent performance within their respective modalities, they often face challenges in generalizing across multiple modalities due to their construction based on specific prior assumptions.
-Recognizing the intrinsic requirement for intelligent models to utilize multimodal data for geographic object cognition, numerous researchers have endeavored to balance the cohesion and autonomy among modalities to build multimodal models.
+Recognizing the intrinsic requirement for intelligent models to utilize multimodal data for geographic object cognition, numerous researchers have endeavoured to balance the cohesion and autonomy among modalities to construct multimodal models.
 For joint interpretation of two modalities, H.
 Li and X.-J.
-Wu recognized the enhancement possible through the fusion of RGB and Infrared images, which convey information about color, texture, and temperature.
-Given the widespread demand and application prospects in military, medical, and other fields, they proposed a dense fusion method for RGB and Infrared images: DenseFuseli2018densefuse .
-This method, employing a dense module that connects the outputs of all convolutional layers, successfully achieves pixel-level fusion of RGB and Infrared modalities.
+Wu recognized the enhancement possible through fusing RGB and infrared images, which convey information about colour, texture, and temperature.
+Given the widespread demand and application prospects in military, medical, and other fields, they proposed a dense fusion method for RGB and infrared images: DenseFuse[29].
+This method, employing a dense module that connects the outputs of all convolutional layers, successfully achieves pixel-level fusion of the RGB and infrared modalities.
 A.
 Sadeghian et al.
-extended the use of historical trajectories for predicting multiple plausible trajectories in Social GANgupta2018social .
-They introduced RGB images to enhance scene data in Sophiesadeghian2019sophie , achieving better results in trajectory prediction tasks.
+extended the social GANs[25] and introduced RGB images to enhance scene data in Sophie[30], achieving better results in trajectory prediction tasks.
 Recognizing the high complementarity between RGB and SAR images, L.
@@
 Hughes et al.
-proposed a three-step deep neural network framework that utilizes a universal prediction of matching regions, generates heatmaps, and eliminates outliers to match RGB and SAR imageshughes2020deep .
+proposed a three-step deep neural network framework that utilizes a universal prediction of matching regions, generates heatmaps, and eliminates outliers to match RGB and SAR images[31].
 X.
 Li et al.
-introduced DTCDNli2021deep , a model that employs a GAN network to migrate RGB and SAR images to the same feature space, facilitating target detection.
+introduced the DTCDN[32], a model that employs a GAN network to migrate RGB and SAR images to the same feature space, facilitating target detection.
 J.
 Yang et al.
-proposed a dual-stream convolutional network that uses high-resolution multispectral images to enhance the spatial resolution of hyperspectral imagesyang2018hyperspectralandmultispectral .
-Notably, CLIPradford2021learning proposed by A.
-Radford et al.
-associates the RGB modality with the text modality using contrastive learning.
-With pretraining guided by weak supervisory signals from text, CLIP demonstrates outstanding capabilities in both visual single-modal tasks and visual-text multimodal tasks, inspiring a series of subsequent workslu2019vilbert ; kim2021vilt ; alayrac2022flamingo ; li2022blip .
+proposed a dual-stream convolutional network that uses high-resolution multispectral images to enhance the spatial resolution of hyperspectral images[33].
+Notably, CLIP[5], proposed by A.
+Radford et al., associates the RGB modality with the text modality using contrastive learning.
+With pretraining guided by weak supervisory signals from text, CLIP has demonstrated outstanding capabilities in both visual single-modal tasks and visual-language multimodal tasks, inspiring a series of subsequent works[6, 7, 8, 9].
 Y.
 Zhang et al.
-introduced Meta-transformerzhang2023meta , leveraging the contrastive learning paradigm from CLIP to pretrain a universal backbone network under the visual-language modality.
-It exhibits multimodal generalization abilities across various modalities such as point clouds, infrared, and hyperspectral.
+introduced a meta-transformer[34], leveraging the contrastive learning paradigm from CLIP to pretrain a universal backbone network under the visual-language modality.
+It exhibits multimodal generalization abilities across various modalities, such as point clouds, infrared, and hyperspectral data.
 J.
 Han et al.
-directly employed a multimodal large language model as a universal backbone network, proposing the One-LLM, successfully unifying eight modalities including Image, Audio, Video, and Pointhan2023onellm .
-The success of these approaches imply the unique role of language modalities in multimodal model.
-Therefore, building upon the aforementioned efforts, we systematically propose the fundamental principle, Language as Reference Framework (LaRF).
-Guided by this principle, we balance the cohesion and autonomy among diverse modalities and introduce a general intelligent model named AllSpark, which unifies 13 spatiotemporal modalities and possesses the potential to extend to an arbitrary number of modalities.
-3 Method
-3.1 Theory
-We observe that in the process of comprehending the world, humans integrate information from multiple modalities such as hearing, touch, smell, and vision.
-The concepts formed through the parsing of these modalities are ultimately converged in language.
-It is through language that humans engage in associating, reasoning, and expressive behaviors.
+directly employed a multimodal large language model as a universal backbone network, proposing the One-LLM, which successfully unifies eight modalities, namely, images, audio, videos, and points[35].
+The success of these approaches implies the unique role of language modalities in multimodal models.
+Therefore, building upon the aforementioned efforts, we systematically propose the fundamental principle of the LaRF.
+Guided by this principle, we balance cohesion and autonomy among diverse modalities and introduce a general intelligent model named AllSpark, which unifies thirteen spatio-temporal modalities and possesses the potential to extend to an arbitrary number of modalities.
+III Method
+III-A Theory
+We observe that in the process of comprehending the world, humans integrate information from multiple modalities, such as hearing, touch, smell, and vision.
+The concepts formed through the parsing of these modalities ultimately converge in language.
+Humans engage in associating, reasoning, and expressive behaviours through language.
 In other words, language precisely encodes human perception and understanding of the world, providing clear definitions and meanings to abstract concepts from each modality.
-Inspired by our observation, to balance the cohesion and autonomy among multiple modalities, we introduce the fundamental principle, Language as Reference Framework (LaRF). The significance is:
+Inspired by our observation, to balance cohesion and autonomy among multiple modalities, we introduce the fundamental principle of Language as Reference Framework (LaRF).
+Figure 2: Guided by the LaRF principle, multimodal data are transformed into a token-context structure similar to language under their respective prior assumptions while maintaining the autonomy of the multimodalities and achieving cohesion among modalities, thus enabling interpretation of multimodal data in the unified language representation space.
+In terms of the cohesion of multimodalities, the high heterogeneity between multiple spatio-temporal modalities is a major challenge, while the LaRF principle defines the alignment anchor between multimodalities as language explicitly.
+As shown in Figure 2, we observe that language is encoded by tokens and their contexts, and this structure can be extended to most spatio-temporal modalities.
+Therefore, we can align highly heterogeneous spatio-temporal modalities to language modalities in structure and semantics, enabling multimodal interpretation in the unified representation space of language.
+In addition, the important guiding role of natural language prompts is also one of the reasons why the LaRF principle can achieve cohesion among modalities.
+In contrast, we can independently encode multiple spatio-temporal modalities into token sequences under their respective prior assumptions, so the LaRF principle does not lead to the loss of modal autonomy.
+More importantly, LaRF is not dependent on specific modalities; therefore, theoretically, as long as token representations of modalities can be obtained, the multimodal model guided by LaRF can be extended to arbitrary modalities.
+In summary, the significance of LaRF is as follows:
   1. 1.
     Alignment Capability: Language can accurately encode both cohesion and autonomy information across multiple modalities.
-Aligning each modality with the language modality enables the attainment of a unified representation in the same feature space, addressing the challenge of high heterogeneity among modalities.
+Aligning each modality with the language modality enables a unified representation in the same feature space, addressing the challenge of high heterogeneity among modalities.
   2. 2.
-    Reasoning Capability: Language, being a tool for human thought and expression, inherently possesses the ability to perform complex reasoning.
+    Reasoning Capability: Language, as a tool for human thought and expression, inherently possesses the ability to perform complex reasoning.
 Each modality, when represented in a unified space with LaRF, inherits the reasoning capability of language, unlocking the potential for multimodal joint reasoning.
   3. 3.
-    Interpretability: Deep learning have long been characterized as opaque ”black boxes.” However, a multimodal intelligent system constructed based on the LaRF can directly leverage language as a tool.
+    Interpretability: Deep learning methods have long been characterized as opaque ”black boxes.” However, a multimodal intelligent system constructed based on the LaRF can directly leverage language as a tool.
 This facilitates the output of interpretable reasoning chains that humans can understand, thereby achieving true explainable artificial intelligence.
@@
     Scalability: The multimodal system guided by the LaRF is agnostic to specific modalities.
-New modalities only need to establish a mapping to the language model to participate in joint reasoning with other modalities.
+New modalities need to establish a mapping to the language model only to participate in joint reasoning with other modalities.
 Therefore, theoretically, a multimodal model based on LaRF can be extended to an arbitrary number of modalities.
-In summary, we contend that intelligent models, when engaged in multimodal perception processes akin to those of humans, should adopt language as the reference framework for abstract concepts across modalities.
-The analysis and reasoning should occur in the feature space of language.
-As illustrated in the conceptual model in Figure 2, when constructing specific multimodal models, data from each modality should be encoded into modality-independent tokens with contextual information.
-These modality-specific tokens should align with the tokens of language.
-A universal intelligent model, primarily based on the Transformer architecture, can then be employed for parsing and reasoning, ultimately generating a comprehensive multimodal representation.
-Figure 2: The conceptual model with LaRF
-3.2 Overview
-Table 1: List of symbols
+Figure 3: AllSpark Architecture.
+Multimodal data are extracted by their respective modal encoders into token sequences.
+After undergoing dimension alignment with modal-specific text prompt tokens through a modal bridge, the text prompt tokens and modal tokens enter a large language multimodal model for interpretation.
+Finally, the interpretation result is matched with downstream tasks by task-specific task heads.
+III-B Overview
+TABLE I: List of symbols
 Module Variants
@@
 modal i
-$f_{i}$ modal encoder for $q$ query vectors of the bridge
-$\Phi$ modal bridge $s_{i}$ tokens of $m_{i}$
-T text tokenizer $p_{i}$ text prompt for $m_{i}$
+$f_{i}$
+modal encoder for $m_{i}$
+$q$ query vectors of the bridge
+$\Phi$
+modal bridge
+$s_{i}$ tokens of $m_{i}$
+$T$
+text tokenizer
+$p_{i}$ text prompt for $m_{i}$
 $F$
-Multi-modal LLM
-$\theta$
-model parameters
+Multimodal LLM
+$\theta$ model parameters
 $H_{task}$
 Task Head
-$y$
-label
-$L$ loss function $t_{i}$ tokens of $m_{i}$
+$y$ label
+$L$
+loss function
+$t_{i}$ tokens of $m_{i}$
 ${Enc}_{n}$
 n-layer transformer encoders
-$W$
-weights of the linear layer
+$W$ weights of the linear layer
 $Emb$
 embedding layer
-$w_{i}$
-word of the code or text modal
+$w_{i}$ word of the code or text
 ${Enc}_{ResNet}$
@@
 feedforward network
-Guided by the principles mentioned above, the AllSpark we proposed consists of five modules: Modal Encoder, Modal Bridge, Text Tokenizer, Multimodal Large Language Model, and Task Head.
+$\sigma$
+softmax
+Guided by the principles mentioned above, the AllSpark we proposed consists of five modules: the modal encoder, modal bridge, text tokenizer, multimodal large language model, and task head.
 The overall architecture is depicted in Figure 3.
-Figure 3: AllSpark Architecture
-To maintain the autonomy among modalities, we designed modality-specific encoders to encode highly heterogeneous data into modality-independent tokens (Section 3.3 for details).
-However, the dimens ion of tokens outputted by modal encoders are still inconsistent.
-To achieve the goal of parsing in the unified representational space of language, we introduced the modal bridge from the Lynxzeng2023matters .
-The modal bridge aims to project tokens from each modality into the dimension of the Multimodal Large Language Model (Section 3.4 for details).
+To maintain autonomy among modalities, we designed modality-specific encoders to encode highly heterogeneous data into modality-independent tokens (for details, see Section III-C).
+However, the dimensions of tokens outputted by modal encoders are still inconsistent.
+To parse in the unified representational space of language, we introduced the modal bridge from the Lynx[10].
+The modal bridge aims to project tokens from each modality into the dimension of the multimodal large language model (Section III-D for details).
 The formalization of this process is as follows:
@@
 Here, $\{m_{RGB},m_{MSI},m_{HSI},\ldots,m_{i},\ldots,m_{Video}\}$ represents inputs from various modalities, where $m_{RGB}$ represents the RGB modality, $m_{MSI}$ represents the MSI modality, and so on.
-$f_{i}$ is defined as the modal encoder for the $m_{i}$ modality, $\Phi$ represents the modal bridge, $q\in R^{N*D}$represents $N$ learnable vectors of dimension $D$ in the modality bridge, where $D$ is set to 4096, representing the dimensionality of the Multimodal Large Language Model.
-Each modality’s input data $m_{i}$ is mapped to a token sequence $s_{i}\in R^{N*D}$ of the same dimension as the language model.
-Table 1 summarizes the main mathematical symbols and their meanings in the model.
+$f_{i}$ is defined as the modal encoder for the $m_{i}$ modality, $\Phi$ represents the modal bridge, and $q\in R^{N*D}$represents $N$ learnable vectors of dimension $D$ in the modality bridge, where $D$ is set to 4096, representing the dimensionality of the multimodal large language model.
+The input data $m_{i}$ of each modality are mapped to a token sequence $s_{i}\in R^{N*D}$ of the same dimension as the language model.
+Table I summarizes the main mathematical symbols and their meanings in the model.
 To achieve cohesion among modalities, we employ a unified multimodal large language model to parse data from various modalities.
-The text tokenizer and Multimodal Large Language Model in AllSpark are based on the visual-language model Lynxzeng2023matters .
-To extend Lynx to 13 spatiotemporal modalities, we design specific text prompts for each modality to guide the model in correctly parsing information from each modality.
+The text tokenizer and multimodal large language model in AllSpark are based on the visual-language model Lynx[10].
+To extend Lynx to thirteen spatio-temporal modalities, we designed specific text prompts for each modality to guide the model in correctly parsing information from each modality.
 Additionally, Lynx incorporates several lightweight multimodal adapter layers internally to accommodate multimodal inputs.
-We continue this design and do not freeze the parameters of the adapter layers during training to enhance the model’s adaptability to other spatiotemporal modalities.
-Finally, we acknowledge that there exists a gap between the parsing results of the model and the downstream task.
+We continue this design and do not freeze the parameters of the adapter layers during training to enhance the model’s adaptability to other spatio-temporal modalities.
+Finally, we acknowledge that a gap exists between the parsing results of the model and those of the downstream task.
 Therefore, we design specific task heads for each task to enhance the model’s generalization capability.
@@
 where $p_{i}$ represents the text prompt of $m_{i}$, $T$ denotes the text tokenizer, $\oplus$ indicates the concatenation operation of text tokens and modality tokens in the sequence, $F$ represents the multimodal large language model, and $H_{task}$ signifies the task head.
-All tasks in our experiments are supervised tasks, with $y$ denoting the labels, $L$ representing the loss function, and $\theta$ representing the learnable parameters in our model.
-The optimization objective of the model can be formalized as:
-$\theta_{i}=\mathop{\arg\min}_{\theta}{L\left(y,M\left(m_{i},p_{i}\right),\theta_{i}\right)}$ (3)
-3.3 Independent Encoder for Each Modality
-The modal encoder aims to encode the raw data of each modality into a token sequence, formalized as: $t_{i}=f_{i}\left(m_{i}\right)$, where $t_{i}\in R^{n*d}$.
-We have designed different modal encoders for each modality to maintain autonomy among modalities.
+All the tasks in our experiments are supervised tasks, with $y$ denoting the labels, $L$ representing the loss function, and $\theta$ representing the learnable parameters in our model.
+The optimization objective of the model can be formalized as follows:
+$\theta_{i}=\mathop{\arg\min}_{\theta}{L\left(y,M\left(m_{i},p_{i}\right);\theta_{i}\right)}$ (3)
+III-C Independent Encoder for Each Modality
+The modal encoder aims to encode the raw data of each modality into a token sequence, formalized as $t_{i}=f_{i}\left(m_{i}\right)$, where $t_{i}\in R^{n*d}$.
+We designed different modal encoders for each modality to maintain autonomy among modalities.
 The following provides individual introductions for each modality:
-3.3.1 1D modal
-Code & Text: Code is essentially a specialized form of language, but due to its distinct properties—such as having a strict syntax, being unambiguous, and being capable of interacting with machines—we separate it as a distinct modality.
+III-C1 1D modal
+Code & Text: Code is essentially a specialized form of language, but due to its distinct properties such as having a strict syntax, being unambiguous, and being capable of interacting with machines.
+we separate it as a distinct modality.
 Exploring intelligent methods for the code modality is crucial for reliable AI reasoning and achieving interaction between intelligent models and the real world.
 Since the Lynx model is a language model, we do not design an additional modal encoder for the code and text modalities.
-Instead, we directly utilize Lynx’s text tokenizer, i.e., $f\left(m_{Text/Code}\right)=T\left(m_{Text/Code}\right)$, where $m_{Text/Code}=\{w_{1},w_{2},w_{3},\ldots\}$ represents the sequence of words in text or code.
-3.3.2 2D modal
-RGB: RGB imagery represents the visible light spectrum, reflecting the electromagnetic characteristics of objects that emit or reflect visible light waves.
-It is the most common modality in visual research.
+Instead, we directly utilize Lynx’s text tokenizer, i.e., $f\left(m_{Text/Code}\right)=T\left(m_{Text/Code}\right)$, where $m_{Text/Code}=\{w_{1},w_{2},w_{3},\ldots\}$ represents the sequence of words in the text or code.
+III-C2 2D modal
+RGB: RGB imagery represents the visible light spectrum and reflects the electromagnetic characteristics of objects that emit or reflect visible light waves.
+It is the most common modality used in visual research.
 RGB imagery is a standard three-band image, i.e., $m_{RGB}\in R^{H*W*3}$.
-For the modal encoder of this modality, we adopted the visual encoder from the Lynx model: EVA.
-EVA is a large visual backbone model composed of 40 stacked Transformer Blocks with a width of 1408.
-During the experiments, AllSpark loaded the official weights of the EVA model and froze them for training.
+For the modal encoder of this modality, we adopted the visual encoder from the Lynx model: the EVA.
+The EVA is a large visual backbone model composed of 40 stacked transformer blocks with a width of 1408.
+During the experiments, AllSpark loaded the official weights of the EVA model and froze them during training.
 MSI: Multispectral imagery is a modality extensively studied in the remote sensing field.
 Typically, it includes more than the three visible light bands present in the RGB modality.
-MSI incorporates multiple non-visible light bands such as near-infrared, short-wave infrared, coastal atmospheric aerosol, and cirrus bands.
+MSIs incorporate multiple nonvisible light bands, such as near-infrared, shortwave infrared, coastal atmospheric aerosol, and cirrus bands.
 Therefore, the number of channels in the imagery is usually greater than the three bands in RGB imagery, i.e., $m_{MSI}\in R^{H*W*C}$, where $C>3$.
-We extended the PatchEmbed of the standard Vision Transformerdosovitskiy2020image , modifying its channel count to match the number of bands in the input multispectral imagery.
+We extended the PatchEmbed of the standard ViT[18], modifying its channel count to match the number of bands in the input multispectral imagery.
 This adaptation allows it to serve as the feature encoder for the multispectral modality.
-HSI: Hyperspectral imagery increases the number of bands compared to multispectral imagery, often reaching hundreds of bands, with each band containing rich information about the terrain.
+HSI: Hyperspectral imagery increases the number of bands compared to multispectral imagery, often reaching hundreds of bands, with each band containing rich information.
 Unlike RGB and multispectral imagery, where an image serves as a single sample, in hyperspectral imagery, all bands for each pixel are treated as a single sample, i.e., $m_{HSI}\in R^{1*1*C}$.
@@
 The entire process can be represented as $f\left(m_{HSI}\right)={Enc}_{12}\left(W*m_{HSI}^{T}\right)$.
-Table: The Table modality consists of several rows and columns, where each row represents a record, and each column represents a different attribute, i.e., $m_{Table}\in R^{{row}*{col}}$.
-The modal encoder for the Table modality inherits the design from TabFormerpadhi2021tabular : first, based on the different degrees of discreteness for each column attribute, we use independent Embedding Layers to encode discrete and continuous values separately.
+Table: The table modality consists of several rows and columns, where each row represents a record and each column represents a different attribute, i.e., $m_{Table}\in R^{{row}*{col}}$.
+The modal encoder for the table modality inherits the design from TabFormer[36]: first, based on the different degrees of discreteness for each column attribute, we use independent embedding layers to encode discrete and continuous values separately.
 Subsequently, we employ a single-layer transformer encoder to further extract features.
 The entire process can be formalized as $f\left(m_{Table}\right)={Enc}_{1}\left(Emb(m_{Table})\right)$.
-Trajectory: The trajectory modality reflects the changing information of an object over time and space, composed of a series of two-dimensional coordinate points, i.e., $m_{Trajectory}\in R^{l*2}$, where $l$ represents the sequence length of trajectory points.
-The encoder for the trajectory modality inherits the design from TUTRshi2023trajectory : first, a linear layer is used to expand the dimensions of the two-dimensional trajectory features.
+Trajectory: The trajectory modality reflects the changing information of an object over time and space and is composed of a series of two-dimensional coordinate points, i.e., $m_{Trajectory}\in R^{l*2}$, where $l$ represents the sequence length of trajectory points.
+The encoder for the trajectory modality inherits the design from TUTR[37]: first, a linear layer is used to expand the dimensions of the two-dimensional trajectory features.
 This step can be formalized as ${W*m}_{Trajectory}$, where $W\in R^{d*\left(l*2\right)}$ is the weight matrix of the linear projection layer.
@@
 SAR: The synthetic aperture radar (SAR) modality is a type of active remote sensing that reflects the electromagnetic characteristics of objects with respect to microwave backscatter.
-Due to different polarization modes, the final product of SAR imagery is typically a two-band or single-band image, i.e., $m_{SAR}\in R^{H*W*2}$.
-Therefore, we designed a simple three-layer convolutional network as its modal encoder.
-Infrared: The dataset we selected for infrared is SYSU-MM01, which is, in fact, a multi-modal re-identification dataset comprising both visible and near-infrared modalities.
+Due to differences in polarization modes, the final product of SAR imagery is typically a two-band or single-band image, i.e., $m_{SAR}\in R^{H*W*2}$.
+Therefore, we designed a simple three-layer convolutional network as the modal encoder.
+Infrared: The dataset we selected for infrared detection was SYSU-MM01, which is, in fact, a multimodal reidentification dataset comprising both visible and near-infrared (NIR) modalities.
 The model needs to simultaneously process infrared and visible light images, i.e., $m_{Infrared}=\{{I_{r}\in R}^{H*W*3},{I_{i}\in R}^{H*W*3}\}$.
-We inherited and slightly modified the design of DDAGye2020dynamic , discarding the shared ResNet-50 layers’ parameters in DDAG and employing two independently parameterized ResNet-50 networks.
+We inherited and slightly modified the design of DDAG[38], discarding the parameters of the shared ResNet-50 layers in DDAG and employing two independently parameterized ResNet-50 networks.
 These networks extract features from infrared and visible light images, and this process can be formalized as $f\left(m_{Infrared}\right)={Enc}_{ResNet}\left(I_{r}\right)\oplus\ {Enc}_{ResNet}\left(I_{i}\right)$.
-Graph: A graph is constructed from a series of nodes and edges, where the attributes of the nodes and the adjacency attributes of the nodes reflect the majority of the graph’s features, i.e., $m_{Graph}\in R^{K*d}$, where K is the number of nodes, and d is the node feature dimension.
-In AllSpark, the graph’s modality encoder is based on STAEformerliu2023spatio , with the main design idea being to first use a linear layer to extend its feature dimension and then use several Embedding Layers to separately encode features such as the node’s own characteristics, spatial characteristics, and temporal characteristics in the graph.
+Graph: A graph is constructed from a series of nodes and edges, where the attributes of the nodes and the adjacency attributes of the nodes reflect the majority of the features of the graph, i.e., $m_{Graph}\in R^{K*d}$, where K is the number of nodes and d is the node feature dimension.
+In AllSpark, the modality encoder of a graph is based on the STAEformer[39], whose main design idea is to first use a linear layer to extend its feature dimension and then use several embedding layers to separately encode features such as the node’s characteristics, spatial characteristics, and temporal characteristics.
 The entire process can be formalized as $f\left(m_{Graph}\right)={Emb}_{node}\left({W*m}_{Graph}^{T}\right)\oplus{Emb}_{spatial}\left({W*m}_{Graph}^{T}\right)\oplus{Emb}_{time}\left({W*m}_{Graph}^{T}\right)$, where $W\in R^{{hidden}*d}$ is the weight of the linear layer.
-Oblique Photography: The oblique photography modality consists of images from multiple views, i.e., $m_{video}=\{V_{1}\in R^{C*H*W},\ V_{3}\in R^{C*H*W},V_{3}\in R^{C*H*W},\ldots\ ,V_{n}\in R^{C*H*W}\}$, where $V_{i}$ represents the image of each view, and n is the number of views.
-In the encoder for oblique photography modality, we use the standard ViT model to extract features from multiple view images and concatenate them.
-The process can be formalized as $f\left(m_{video}\right)={Enc}_{12}\left(V_{1}\right)\oplus{Enc}_{12}\left(V_{2}\right)\oplus{Enc}_{12}\left(V_{3}\right)\oplus\ldots\oplus{Enc}_{12}\left(V_{n}\right)$.
-3.3.3 3D modal
+Oblique Photography: The oblique photography modality consists of images from multiple views, i.e., $m_{Oblique}=\{V_{1}\in R^{C*H*W},\ V_{3}\in R^{C*H*W},V_{3}\in R^{C*H*W},\ldots\ ,V_{n}\in R^{C*H*W}\}$, where $V_{i}$ represents the image of each view and n is the number of views.
+In the encoder for the oblique photography modality, we use the standard ViT model to extract features from multiple view images and concatenate them.
+The process can be formalized as $f\left(m_{Oblique}\right)={Enc}_{12}\left(V_{1}\right)\oplus{Enc}_{12}\left(V_{2}\right)\oplus{Enc}_{12}\left(V_{3}\right)\oplus\ldots\oplus{Enc}_{12}\left(V_{n}\right)$.
+III-C3 3D modal
 Video: The video modality can be considered as the stacking of images in the temporal dimension, i.e., $m_{video}\in R^{T*C*H*W}$.
 The significant characteristic of the video modality is its high redundancy in both temporal and spatial dimensions.
-To address this issue, the encoder for the video modality in AllSpark is based on TubeViTpiergiovanni2023rethinking , which performs sparse sampling on the highly redundant data in both time and space to reduce redundancy.
-The features obtained from sparse sampling are then fed into a 6-layer Transformer Encoder for further feature extraction.
+To address this issue, the encoder for the video modality in AllSpark is based on TubeViT[40], which performs sparse sampling on highly redundant data in both time and space to reduce redundancy.
+The features obtained from sparse sampling are subsequently fed into a 6-layer transformer encoder for further feature extraction.
 The process can be formalized as $f\left(m_{video}\right)={Enc}_{6}\left({Sparse}(m_{video})\right)$.
-PointCloud: PointCloud are typically composed of three-dimensional coordinates and their feature values, $m_{pointCloud}\in R^{K*\left(d+3\right)}$, where K represents the number of three-dimensional points, and d represents the dimensionality of the point cloud features, reflecting information such as the spatial position, shape, color, and texture of objects.
-The encoder for the point cloud modality inherits the design from PointBERTyu2022point : firstly, point cloud data is grouped and encoded to unify the number of points simultaneously inputted.
-This step can be represented as $PointGroup=Grouper\left(m_{pointCloud}\right)\in R^{G*N*3}$, where $G$ represents the number of groups, and $N$ represents the number of points in each group.
-Next, the grouped results are inputted into a one-dimensional convolutional layer to extract feature vectors for each group:$f_{Group}=Conv1d\left(PointGroup\right)\in R^{G*d}$.
-Finally, the feature vectors for each group are inputted into a standard 12-layer Transformer Encoder to extract their global features.
+PointClouds: PointClouds are typically composed of three-dimensional coordinates and feature values $m_{PointCloud}\in R^{K*\left(d+3\right)}$, where K represents the number of three-dimensional points and d represents the dimensionality of the point cloud features, reflecting information such as the spatial position, shape, colour, and texture of objects.
+The encoder for the point cloud modality inherits the design from PointBERT[41]: first, point cloud data are grouped and encoded to unify the number of points simultaneously inputted.
+This step can be represented as $PointGroup=Grouper\left(m_{PointCloud}\right)\in R^{G*N*3}$, where $G$ represents the number of groups and $N$ represents the number of points in each group.
+Next, the grouped results are input into a one-dimensional convolutional layer to extract feature vectors for each group:$f_{Group}=Conv1d\left(PointGroup\right)\in R^{G*d}$.
+Finally, the feature vectors for each group are input into a standard 12-layer transformer encoder to extract their global features.
 The entire process can be formalized as $f\left(m_{PointCloud}\right)=Enc_{12}\left(Conv1d(Grouper(m_{PointCloud}))\right)$.
-3.4 Modal2Language Bridge
+III-D Modal2Language Bridge
+Although the modal encoders have transformed data from various heterogeneous modalities into a unified token sequence, there are still differences in dimensions between different modal tokens, making it difficult to perceive by a multimodal large language model.
 The modal bridge, based on the Lynx model, aims to perform dimensional projection from tokens of various modalities to tokens of the language modality.
 In its implementation, the modal bridge consists of stacked layers, including cross-attention layers and feedforward neural network layers.
-In the cross-attention layer, we predefine a learnable query vector $Q\in R^{N*D}$, where $D$ is the internal dimensionality of the language model, and $N$ serves as a hyperparameter that can be flexibly adjusted to accommodate inputs from different modalities.
+In the cross-attention layer, we predefine a learnable query vector $Q\in R^{N*D}$, where $D$ is the internal dimensionality of the language model and $N$ serves as a hyperparameter that can be flexibly adjusted to accommodate inputs from different modalities.
 The keys and values in the cross-attention layer are the features outputted by the modal encoders.
-The feedforward neural network inherits the classic design from the original Transformer, consisting of two linear layers with an inserted activation layer. The entire process can be formalized as:
-$\Phi\left(Q,s_{i}\right)={FFN}\left({softmax}\left(\frac{W_{q}Q^{T}\left(w_{k}s_{i}^{T}\right)^{T}}{\sqrt{D}}\right)*(W_{v}s_{i}^{T})\right)$ (4)
-Here, $W_{q}\in R^{D*{hidden}}$, $W_{k}\in R^{d*{hidden}}$, and $W_{v}\in R^{d*{hidden}}$are the linear projection layer weights defined inside the cross-attention layer for $Q$, $K$ and $V$ respectively.
-3.5 Task-Guided Text Prompts and Task-Heads
-To extend the visual-language multimodal model to 13 spatiotemporal modalities without intervention from modality expert knowledge, we designed specific text prompts and task heads for each modality and task.
+The feedforward neural network inherits the classic design from the original transformer and consists of two linear layers with an inserted activation layer.
+The entire process can be formalized as follows:
+$\Phi\left(Q,s_{i}\right)={FFN}\left(\sigma\left(\frac{QW_{q}^{T}\left(s_{i}W_{k}^{T}\right)^{T}}{\sqrt{D}}\right)s_{i}W_{v}^{T}\right)$ (4)
+Here, $W_{q}\in R^{D*{hidden}}$, $W_{k}\in R^{d*{hidden}}$, and $W_{v}\in R^{d*{hidden}}$ are the linear projection layer weights defined inside the cross-attention layer for $Q$, $K$ and $V$, respectively.
+$\sigma$ denotes the softmax operation.
+III-E Task-Guided Text Prompts and Task Heads
+To extend the visual-language multimodal model to thirteen spatio-temporal modalities without intervention from modality expert knowledge, we designed specific text prompts and task heads for each modality and task.
 Text prompts are used to guide the multimodal language model in correctly interpreting each modality’s data, while task heads are employed to match the model’s parsing results with specific downstream tasks.
@@
 However, during testing, for the sake of result stability and reproducibility, the prompt was fixed to be the first prompt among all prompts.
-Table 2 provides a list of all the text prompts.
-Table 2: List of prompts
-Dimension
-Modal
-Prompt
-1D
-Text
-1.Please determine if this movie review is positive or negative?
-Code
-2D
-RGB
-1.This remote sensing image belongs to which of the following categories: [Category list of the NWPU dataset] 2.Describe this remote sensing image briefly 3.Find a word that is most relevant to this remote sensing image 4.Describe the key elements in this remote sensing image
-MSI
-1.Based on the multi-spectral imagery feature description, please classify this object 2.Given the following multi-spectral imagery characteristics, please output the most fitting scene label
-HSI
-1.Given the spectral information, can you help determine which class this pixel belongs to?
-2.Here is the spectral data for a pixel.
-Considering the typical characteristics of land cover classes, could you provide a detailed analysis and suggest the most likely class for this pixel?
-3.The spectral information for a pixel is given, but the data is noisy.
-Given the potential variability, which land cover classes should be considered as possible candidates for this pixel?
-Table
-1.Please utilize the provided air quality indicators to accurately predict the concentrations of PM2.5 in the atmosphere 2.Given the air quality indicators, please provide a prediction for the PM2.5 levels at this particular moment
-Oblique Photography
-1.Please reconstruct the 3D model from the given images of five different views.
-Trajectory
-1.Based on their past positions and movements in a crowded environment, predict the future trajectory of a selected pedestrian.
-2.Using the pedestrian trajectory data, along with additional information about the surrounding environment, predict the future path of the pedestrian.
-3.Given the current and past positions of a pedestrian and their neighboring pedestrians, predict the main pedestrian’s trajectory
-SAR
-1.Based on the SAR imagery feature description, please classify this object 2.Given the following SAR imagery characteristics, please output the most fitting scene label
-Infrared
-1.Given an infrared image of a person, find and highlight the same person in a set of visible light images, paying close attention to features like clothing, posture, and gait 2.Analyze the characteristics of a person in a visible light image, such as clothing texture, color, and body shape.
-Then, locate the person with matching features in a series of infrared images
-Graph
-1.Given the current traffic data including vehicle flow rate, average speed, and time of day from the METRLA dataset, predict the traffic flow 2.Analyze the historical data on vehicle speeds and flow rates from the METRLA dataset for the past week.
-Identify any patterns or trends and predict the traffic conditions
-3D
-Video
-1.Please analyze this video and describe the characteristics of the action in detail 2.Watch this short clip and provide a step-by-step description of the action sequence
-Point Cloud
-1.Classify the provided point cloud sample into the correct category 2.Look at the point cloud data characteristics and classify the object 3.Please analyze the given point cloud dataset and determine which category it belongs to.
-Focus on the shape and structure evident in the point cloud
+Table XVI provides a list of all the text prompts.
 Additionally, we noticed that there is still a gap between the multimodal language model’s parsing results and the specific downstream tasks.
@@
 Moreover, to ensure the transferability of modalities across different tasks, the design principle for task heads is to be as simple and lightweight as possible.
-The detailed summary of each modality’s task heads can be found in Table 4.
+A detailed summary of each modality’s task heads can be found in Table 4.
 For classification tasks or downstream tasks that can be formalized as classification tasks, we uniformly use a simple single-layer linear layer as the task head.
 For instance, we implemented standard classification tasks on the RGB, MSI, SAR, video, and PointCloud modalities.
-Although the task on the HSI modality is a segmentation task, it can be formalized as a per-pixel classification task.
-Similarly, for the re-identification task on the Infrared modality, it can also be formalized as a classification task.
-Therefore, a single-layer linear layer is used as task head for the mentioned modalities.
+Although the task involving the HSI modality is segmentation, it can be formalized as a per-pixel classification task.
+Similarly, the reidentification task in the infrared modality can also be formalized as a classification task.
+Therefore, a single-layer linear layer is used as the task head for the mentioned modalities.
 For regression tasks on the table, trajectory, and graph modalities, we also use a linear layer to perform regression predictions.
@@
 Since the Lynx model itself is a language model, the code and text modalities directly use its native text decoder.
-The task head for the Oblique Photography modality is based on the 3D reconstruction architecture from the Ada-MVSliu2023deep model.
-4 Experiment
-4.1 Setup
-Our experiments aimed to evaluate the adaptability of AllSpark to various spatiotemporal modalities.
-To provide a realistic evaluation of its multimodal adaptability and facilitate comparisons with existing work in the field, we followed the principles of simplicity, reproducibility, and choosing datasets and tasks that have been widely studied in the domain.
-The specific details on dataset selection refer to Section 4.2.
-Additionally, for result reproducibility, we employed similar experimental settings across all modalities.
-We used the AdamW optimizer with a learning rate schedule based on cosine annealing.
-The hyperparameters were adjusted slightly in terms of training epochs and learning rates for different experiments.
-Table 3 summarizes the hyperparameter settings for experiments on each modality.
-Table 3: Summary of hyperparameters
+The task head for the oblique photography modality is based on the 3D reconstruction architecture from the Ada-MVS[42] model.
+IV Experiment
+IV-A Setup
+TABLE II: Summary of hyperparameters
 Dimension
@@
 Text
-9.0*10-6
-5 1
+$9.0\times 10^{6}$
+5
+1
 Code
-1.0*10-5
-4 1
+$1.0\times 10^{5}$
+4
+1
 2D
 RGB
-5.0*10-5
-50 5
+$5.0\times 10^{5}$
+50
+5
 MSI
-2.0*10-5
-50 5
+$2.0\times 10^{5}$
+50
+5
 HSI
-1.0*10-4
-30 3
+$1.0\times 10^{4}$
+30
+3
 Infrared
-5.0*10-5
-50 5
+$5.0\times 10^{5}$
+50
+5
 SAR
-9.0*10-6
-30 3
-ObliquePhotography 5.0*10-5 30 3
+$9.0\times 10^{6}$
+30
+3
+Oblique Photography
+$5.0\times 10^{5}$
+30
+3
 Table
-2.0*10-5
-30 3
+$2.0\times 10^{5}$
+30
+3
 Trajectory
-1.0*10-5
-30 5
+$1.0\times 10^{5}$
+30
+5
 Graph
-8.0*10-5
-10 2
+$8.0\times 10^{5}$
+10
+2
 3D
 PointCloud
-3.0*10-5
-100 10
+$3.0\times 10^{5}$
+100
+10
 Video
-1.0*10-5
-3 1
-4.2 Dataset
-Table 4 summarizes the data and tasks used in experiments for each modality. Below, we will provide detailed explanations in order:
-Text: The IMDB dataset is a binary sentiment analysis dataset, consisting of 50,000 reviews from the Internet Movie Database (IMDb) labeled as positive or negative.
-Additionally, the dataset includes some unlabeled data.
-In our experiments, only the labeled data from the IMDB dataset was utilized for supervised sentiment classification tasks.
-Code: CodeSearchNet is a large-scale dataset of function code and its documentation from GitHub, covering six programming languages: Go, Java, JavaScript, PHP, Python, and Ruby.
-The task performed on the code modality is Code Document Generation, and we tested it on the Ruby and JavaScript languages.
-RGB: NWPU-RESISC45 is a large-scale open dataset for visible light remote sensing image scene classification.
-It includes 45 land-use categories such as airplane, baseball diamond, beach, commercial area, etc.
-Each category has 700 remote sensing images, making a total of 31,500 images in the dataset.
-The image size is 256*256 pixels, and we selected a version of the dataset split by the official release, using 20% of the data for training.
-MSI: The EuroSAT dataset is a multispectral modality dataset for land use and land cover (LULC) classification.
+$1.0\times 10^{5}$
+3
+1
+Our experiments evaluated the adaptability of AllSpark to various spatio-temporal modalities.
+To provide a realistic evaluation of its multimodal adaptability and facilitate comparisons with existing work in the field, we followed the principles of simplicity, reproducibility, and choosing datasets and tasks that have been widely studied in the domain.
+The specific details on dataset selection can be found in Section IV-B.
+Additionally, for reproducibility, we employed similar experimental settings across all modalities.
+We used the AdamW optimizer with a learning rate schedule based on cosine annealing.
+The hyperparameters were adjusted slightly in terms of training epochs and learning rates for different experiments.
+Table II summarizes the hyperparameter settings for the experiments on each modality.
+IV-B Dataset
+Table 4 summarizes the data and tasks used in the experiments for each modality. Below, we will provide detailed explanations in order:
+Text: The IMDB[43] dataset is a binary sentiment analysis dataset consisting of 50,000 reviews from the Internet Movie Database (IMDb) labelled as positive or negative.
+Additionally, the dataset includes some unlabelled data.
+In our experiments, only the labelled data from the IMDB dataset were utilized for supervised sentiment classification tasks.
+Code: CodeSearchNet[44] is a large-scale dataset of function code and its documentation from GitHub that covers six programming languages: Go, Java, JavaScript, PHP, Python, and Ruby.
+The task performed on the code modality is code document generation, and we tested it on the Ruby and JavaScript.
+RGB: NWPU-RESISC45[45] is a large-scale open dataset for visible light remote sensing image scene classification.
+The dataset included 45 land use categories, such as airplanes, baseball diamonds, beaches, and commercial areas.
+Each category included 700 remote sensing images, for a total of 31,500 images.
+The image size is 256*256 pixels, and we selected a version of the dataset split by the official release using 20% of the data for training.
+MSI: The EuroSAT[46] dataset is a multispectral dataset for land use and land cover (LULC) classification.
 The samples are sourced from the Sentinel-2 optical satellite and include all 13 bands.
-The data is categorized into 10 classes, with a total of 27,000 images.
+The data are categorized into 10 classes for a total of 27,000 images.
 We adopted a random 9:1 split for training and testing.
@@
 It comprises 103 bands with a size of 610*340 pixels.
-The dataset includes nine land cover categories such as Asphalt, Meadows, Gravel, etc.
+The dataset includes nine land cover categories, such as asphalt, meadows, and gravel.
 We used a 4:6 split for training and testing.
-Table: The PRSA dataset is a collection of air quality data from multiple stations in Beijing, containing hourly measurements of air pollutants.
+Table: The PRSA[47] dataset is a collection of air quality data from multiple stations in Beijing that contains hourly measurements of air pollutants.
 The data spans from March 1, 2013, to February 28, 2017, across 12 monitoring stations.
-In our experiments, we used various features, including time, station information, 4 air pollutant variables (SO2, NO2, CO, O3), and 6 meteorological variables (temperature, pressure, dew point temperature, amount of precipitation, wind speed, wind direction).
+In our experiments, we used various features, including time, station information, 4 air pollutant variables (SO2, NO2, CO, and O3), and 6 meteorological variables (temperature, pressure, dew point temperature, amount of precipitation, wind speed, and wind direction).
 The task was to predict the concentration of PM2.5.
 We split the data into training (40%) and testing sets.
-Trajectory: The ETH-UCY dataset is one of the widely used benchmarks for pedestrian trajectory prediction, divided into five subsets: ETH, HOTEL, UNIV, ZARA1, and ZARA2.
+Trajectory: The ETH-UCY[48, 49] dataset is a widely used benchmark for pedestrian trajectory prediction and is divided into five subsets: ETH, HOTEL, UNIV, ZARA1, and ZARA2.
 In our experiments, we utilized the ETH subset.
-SAR: The MSTAR dataset is a synthetic aperture radar dataset designed for military stationary target recognition, comprising ten categories of military targets.
-We employed the standard operating conditions (SOC) dataset preprocessing method proposed by S.
+SAR: The MSTAR[50] dataset is a synthetic aperture radar dataset designed for military stationary target recognition that comprises ten categories of military targets.
+We employed the standard operating conditions (SOCs) dataset preprocessing method proposed by S.
 Chen et al.
-chen2016target , ensuring that the serial numbers and target configurations are consistent between the test and training sets, while the aspects and depression angles differ.
-Infrared: SYSU-MM01 is a cross-modal human re-identification dataset with optical-infrared modalities, including images from 4 RGB cameras and 2 infrared cameras of 491 different individuals.
-We adopted the official dataset split, with the training set containing 20,284 RGB images and 9,929 infrared images from 296 individuals, while the test set consists of 3,803 infrared images from 96 individuals used as queries and 301 randomly selected RGB images for re-identification targets.
-Graph: METR-LA is a traffic data collection from loop detectors on the highways of Los Angeles, spanning from March 1, 2012, to June 30, 2012. The task is traffic flow prediction.
-Oblique Photography: WHU-OMVS is an oblique photography dataset designed for 3D reconstruction tasks.
+[22], ensuring that the serial numbers and target configurations are consistent between the test and training sets while the aspects and depression angles differ.
+Infrared: SYSU-MM01[51] is a cross-modal human reidentification dataset with optical-infrared modalities that includes images from 4 RGB cameras and 2 infrared cameras of 491 different individuals.
+We adopted the official dataset split, with the training set containing 20,284 RGB images and 9,929 infrared images from 296 individuals and the test set consisting of 3,803 infrared images from 96 individuals used as queries and 301 randomly selected RGB images for reidentification targets.
+Graph: METR-LA is a traffic dataset collected from loop detectors on the Los Angeles highways spanning from March 1, 2012, to June 30, 2012. The task is traffic flow prediction.
+Oblique Photography: WHU-OMVS[42] is an oblique photography dataset designed for 3D reconstruction tasks.
 The dataset provides imagery from five different viewpoints, along with camera parameters and other relevant information.
-It consists of six areas, and in our experiments, we use area 1 as the training set and area 2 as the test set.
-Video: UCF101 dataset is a human action recognition dataset comprising 101 action classes with a total of 13,320 video clips.
-The videos have a combined duration of 27 hours and a resolution of 320*240 pixels, sourced from YouTube.
-PointCloud: ModelNet40 is a synthetic point cloud dataset consisting of 40 object categories and a total of 12,311 point cloud objects.
+It consists of six areas, and in our experiments, area 1 is used as the training set, and area 2 is used as the test set.
+Video: The UCF101[52] dataset is a human action recognition dataset comprising 101 action classes with a total of 13,320 video clips.
+The videos have a combined duration of 27 hours and a resolution of 320*240 pixels and were sourced from YouTube.
+PointCloud: ModelNet40[53] is a synthetic point cloud dataset consisting of 40 object categories and a total of 12,311 point cloud objects.
 We follow the official dataset split, with 9,843 objects used for training and 2,468 for testing.
-4.3 Results
-4.3.1 RGB
-We evaluated the performance of AllSpark on the RGB image scene classification task using the NWPU-RESISC45 dataset, with the Top-1 Accuracy as the evaluation metric.
-AllSpark leverages expert knowledge from Lynx by loading its pretrained weights.
-Therefore in Table 4, we compare AllSpark with the state-of-the-art models known on the NWPU-RESISC45 dataset.
-The results indicate that AllSpark outperforms most baseline models, with a margin of only 0.84 compared to the SOTA (95.69).
-This highlights AllSpark’s exceptional perception and interpretation capabilities in the RGB modality.
-Table 4: RGB scene classification with AllSpark(80% as trainset)
+IV-C Results
+IV-C1 RGB
+TABLE III: rgb image classification with allspark
 Method
@@
 Acc(%)
-CNN-CapsNetzhang2019remote
+CNN-CapsNet[54]
 RS2019
 89.03
-DFAGCNxu2021deep
+DFAGCN[55]
 TNNLS2021
 89.29
-D-CNN with GoogleNetcheng2018deep
+D-CNN with GoogleNet[56]
 TGRS2018
 90.49
-D-CNN with VGGNetcheng2018deep
+D-CNN with VGGNet[56]
 TGRS2018
 91.89
-SCCovhe2019skip
+SCCov[57]
 TNNLS2019
 92.1
-SeCo-ResNet-50manas2021seasonal
+SeCo-ResNet-50[58]
 ICCV2021
 92.91
-MG-CAPwang2020multi
+MG-CAP[59]
 TIP2020
 92.95
-LSENetbi2021local
+LSENet[60]
 TIP2021
 93.34
-MSANetzhang2021multiscale
+MSANet[61]
 JSTARS2021
 93.52
-IDCCPwang2020invariant
+IDCCP[62]
 TGRS2021
 93.76
-MBLANetchen2021remote
+MBLANet[63]
 TIP2021
 94.66
-GRMANet-ResNet-50li2021gated
+GRMANet-ResNet-50[64]
 TGRS2021
 94.72
-EMSNetzhao2023emscnet
+EMSNet[65]
 TGRS2023
 95.37
-ViTAE-B + RVSAwang2022advancing
+ViTAE-B + RVSA[66]
 TGRS2022
@@
 94.85
-4.3.2 MSI
+We evaluated the performance of AllSpark on the RGB image scene classification task using the NWPU-RESISC45 dataset, with the top-1 accuracy as the evaluation metric.
+AllSpark leverages expert knowledge from Lynx by loading its pretrained weights.
+Therefore, in Table III, we compare AllSpark with the state-of-the-art models known on the NWPU-RESISC45 dataset.
+The results indicate that AllSpark outperforms most baseline models, with a margin of only 0.84 compared to that of the SOTA (95.69).
+This highlights AllSpark’s exceptional perception and interpretation capabilities in the RGB modality.
+IV-C2 MSI
 We evaluated the performance of AllSpark on the MSI scene classification task using the EuroSAT dataset.
 In the experiment, all 13 spectral bands of the images were simultaneously input into the model.
-The model’s objective was to correctly classify the images into one of the 10 specified categories, and the evaluation metric chosen was Top-1 Accuracy.
-AllSpark does not possess expert knowledge in the multispectral modality, so we categorize the baseline models into two groups in Table 5: those with expert knowledge intervention and those without.
+The model’s objective was to correctly classify the images into one of the 10 specified categories, and the evaluation metric chosen was the top-1 accuracy.
+AllSpark does not possess expert knowledge in the multispectral modality, so we categorize the baseline models into two groups, as shown in Table IV: those with expert knowledge intervention and those without.
 Expert knowledge intervention refers to baseline models pretraining on large datasets such as BigEarthNet and then fine-tuning on the EuroSAT dataset, while no expert knowledge indicates baseline models trained directly from scratch on the EuroSAT dataset.
@@
 Furthermore, AllSpark lags behind the best result in the expert knowledge group by only 4.75, demonstrating its excellent adaptability to the multispectral modality.
-Table 5: MSI Land Cover Classification with AllSpark
-Expert Knowledge
+TABLE IV: msi land cover classification with allspark
+Expert Know.
 Method
 Publication
-Accuracy (%)
+Acc(%)
 No
-ResNet-18manas2021seasonal
+ResNet-18[58]
 ICCV2021
 63.21
-ResNet-50nampally2023performance
+ResNet-50[67]
 IGARSS2023
 91.13
-InceptionNetzhang2020semi
+InceptionNet[68]
 ICIP2020
 93.07
-EfficientNetgomez2021msmatch
+EfficientNet[69]
 IJSTARS2021
@@
 AllSpark ours 94.03
-ResNet-152nampally2023performance
+ResNet-152[67]
 IGARSS2023
@@
 Yes
-MoCoV2manas2021seasonal
+MoCoV2[58]
 ICCV2021
 89.51
-SeComanas2021seasonal
+SeCo[58]
 ICCV2021
 93.14
-SEERgoyal2022vision
+SEER[70]
 Arxiv2022
 97.6
-DINO-MCwanyan2023dino
+DINO-MC[71]
 Arxiv2023
 98.78
-4.3.3 HSI
-We conducted pixel classification task on the Pavia University dataset for the hyperspectral modality.
+IV-C3 HSI
+We conducted a pixel classification task on the Pavia University dataset for the hyperspectral modality.
 The model treats all spectral bands of a single pixel as one sample and predicts the land cover category of that pixel.
 The reported metrics include overall accuracy (OA), average accuracy (AA), and kappa.
-Since AllSpark does not possess expert knowledge in the hyperspectral modality, we compared it with the semi-supervised baselines summarized by D.
+Since AllSpark does not possess expert knowledge of the hyperspectral modality, we compared it with the semisupervised baselines summarized by D.
 Uchaev and D.
-Uchaev uchaev2023small .
-The results in Table 6 demonstrate that AllSpark outperforms many hyperspectral image classification methods such as IFRF and S-DMM, with a difference of 6.42 compared to the best result in terms of OA, highlighting AllSpark’s superior adaptability to the hyperspectral modality.
-Table 6: HSI pixel Classification with AllSpark
+Uchaev [72].
+The results in Table V demonstrate that AllSpark outperforms many hyperspectral image classification methods, such as IFRF and S-DMM, by a factor of 6.42 compared to the best result in terms of OA, highlighting AllSpark’s superior adaptability to the hyperspectral modality.
+TABLE V: hsi pixel classification with allspark
 Method
@@
 Kappa
-3D-CNN
-75.24 80.26 68.34
-CA-GAN
-76.81 76.94 71.02
-3D VS-CNN
-81.63 83.86 76.46
-RPNet
-84.92 83.26 80.52
-S-DMM
-88.3 93.76 84.9
-IFRF
-88.38 85.99 84.97
+3D-CNN 75.24 80.26 68.34
+CA-GAN 76.81 76.94 71.02
+3D VS-CNN 81.63 83.86 76.46
+RPNet 84.92 83.26 80.52
+S-DMM 88.3 93.76 84.9
+IFRF 88.38 85.99 84.97
 AllSpark 89.18 86.65 85.32
-DCFSL
-90.71 90.2 87.73
-TC-GAN
-93.2 91.6 91
-PRNet-RF
-95.6 94.96 94.27
-4.3.4 Table
-Table 7: PM2.5 prediction with AllSpark
-Expert Architecture Method PM2.5
+DCFSL 90.71 90.2 87.73
+TC-GAN 93.2 91.6 91
+PRNet-RF 95.6 94.96 94.27
+IV-C4 Table
+For the table modality, we evaluated AllSpark on the regression prediction task using the PRSA[47] dataset.
+The task involves predicting the concentration of PM2.5 in the air using features such as time, site, 4 air pollutants, and 6 meteorological variables (as detailed in Section IV-B).
+The performance metrics include the root mean squared error (RMSE), mean absolute error (MAE), and R-squared (R2).
+TABLE VI: PM2.5 prediction with allspark
+Expert Arch. Method PM2.5
 RMSE MAE R2
-No
-GWOxie2020evolving
-62.2 40.8 -
-TabBERTpadhi2021tabular
-32.8 - -
+No GWO[73] 62.2 40.8 -
+TabBERT[36] 32.8 - -
 AllSpark 29.03 18.04 0.87
-Yes
-Stacked ResNet-LSTMcheng2022stacked
-40.68 23.75 0.8
-CBAM-CNN-Bi LSTMli2022prediction
-18.9 11.2 0.94
-For the Table modality, we evaluated AllSpark on the regression prediction task using the PRSAzhang2017cautionary dataset.
-The task involves predicting the concentration of PM2.5 in the air using features such as time, site, 4 air pollutants, and 6 meteorological variables (as detailed in Section 4.2).
-The performance metrics include Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and R-squared (R2).
-Table 7 presents the comparative results between AllSpark and baselines.
-It’s worth noting that some works specifically focus on the prediction task on the PRSA dataset and design expert models with specific architectures based on dataset characteristics.
-For example, the CBAM-CNN-Bi LSTM proposed by D.
-Li et al.li2022prediction uses CNN to extract spatial dependencies between air monitoring stations and Bi-LSTM to capture the temporal dependencies of PM2.5 data.
+Yes Stacked ResNet-LSTM[74] 40.68 23.75 0.8
+CBAM-CNN-Bi-LSTM[75] 18.9 11.2 0.94
+Table VI presents the comparative results between AllSpark and the baselines.
+It is worth noting that some works specifically focus on the prediction task on the PRSA dataset and design expert models with specific architectures based on dataset characteristics.
+For example, the CBAM-CNN-Bi-LSTM proposed by D.
+Li et al.[75] used a CNN to extract spatial dependencies between air monitoring stations and Bi-LSTM to capture the temporal dependencies of PM2.5 data.
 Similarly, the stacked ResNet-LSTM model proposed by X.
-Cheng et al.cheng2022stacked employs a stacking LSTM strategy to enhance the extraction of temporal features in PM2.5 data.
-Therefore, we categorize baseline methods into two types: with expert structures and without expert structures.
-In models without expert structures, our approach achieves the best performance among the baselines and is slightly inferior to the state-of-the-art method in models with expert structures (CBAM-CNN-Bi LSTM).
+Cheng et al.[74] employs a stacking LSTM strategy to enhance the extraction of temporal features in PM2.5 data.
+Therefore, we categorize baseline methods into two types: those with an expert architecture and those without an expert architecture.
+Among the models without expert architecture, our approach achieves the best performance among the baselines and is slightly inferior to the state-of-the-art method for models with expert architecture (CBAM-CNN-Bi-LSTM).
 This reflects the excellent adaptability of AllSpark to the table modality.
-4.3.5 Code
-Table 8: Code Document Generation with AllSpark
-Expert Knowledge Method MRR
+IV-C5 Code
+For the code modality, we evaluated the performance of AllSpark on the code document generation task using the CodeSearchNet dataset.
+This task involves generating corresponding documents based on the provided function code.
+We conducted tests for both the Ruby and JavaScript languages using the mean reciprocal rank (MRR)[76] as the evaluation metric.
+TABLE VII: code document generation with allspark
+Expert Know. Method MRR
 Ruby Javascript
-No
-BIRNN[68]
-0.084 0.153
-1D-CNN[68]
-0.245 0.352
-selfAtt[68]
-0.365 0.451
-NBoW[68]
-0.429 0.461
-RoBERTa[67]
-0.625 0.606
+No BIRNN[68] 0.084 0.153
+1D-CNN[68] 0.245 0.352
+selfAtt[68] 0.365 0.451
+NBoW[68] 0.429 0.461
+RoBERTa[67] 0.625 0.606
 AllSpark 0.627 0.635
-Yes
-RoBERTa(Code)[67]
-0.661 0.64
-CodeBERT[67]
-0.693 0.706
-GraphCodeBERT[67]
-0.732 0.711
-For code modality, we evaluated the performance of AllSpark on the Code Document Generation task using the CodeSearchNet dataset.
-This task involves generating corresponding documents based on the provided function code.
-We conducted tests for both Ruby and JavaScript languages, using Mean Reciprocal Rank (MRR)guo2020graphcodebert as the evaluation metric.
-Similar to the previous modalities, since MSI-AGI does not possess expert knowledge in the Code modality, we categorized the baselines into two groups: with and without expert knowledge.
-As shown in Table 8, AllSpark achieved the SOTA results in the group without expert knowledge and remained comparable to models with expert knowledge.
-This demonstrates the strong adaptability of AllSpark to the Code modality.
-4.3.6 PointCloud
-For pointcloud modality, we evaluated the performance of AllSpark on the ModelNet40 dataset for the classification task, with Top-1 Accuracy as the metric.
-In the context of single-modal studies focused on point clouds, we observed that due to the unique three-dimensional structure of point cloud data, most works concentrate on designing specific structures to maintain properties like permutation invariance and symmetry in 3D point clouds.
+Yes RoBERTa(Code)[67] 0.661 0.64
+CodeBERT[67] 0.693 0.706
+GraphCodeBERT[67] 0.732 0.711
+Like in the previous modalities, since MSI-AGI does not possess expert knowledge of the code modality, we categorized the baselines into two groups: those with and without expert knowledge.
+As shown in Table VII, AllSpark achieved SOTA results in the group without expert knowledge, and the results were comparable to those of models trained with expert knowledge.
+This finding demonstrates the strong adaptability of AllSpark to the code modality.
+IV-C6 PointCloud
+For the PointCloud modality, we evaluated the performance of AllSpark on the ModelNet40 dataset for the classification task, with the top-1 accuracy as the metric.
+In the context of single-modal studies focused on point clouds, we observed that due to the unique three-dimensional structure of point cloud data, most works concentrate on designing specific structures to maintain properties such as permutation invariance and symmetry in 3D point clouds.
 However, these structures designed for the unique priors of the modality are challenging to transfer across modalities.
 Additionally, some methods tend to pretrain on large point cloud datasets to acquire general modal expert knowledge before generalizing to specific downstream tasks to improve performance.
-Table 9: PointCloud classification with AllSpark
+TABLE VIII: pointcloud classification with allspark
 Method
 Expert Architecture Expert Knowledge Acc(%)
-PointNetqi2017pointnet
+PointNet[1]
 Yes No 89.2
-Kd-netklokov2017escape
+Kd-net[77]
 Yes No 90.6
-SPH3D-GCNlei2020spherical
+SPH3D-GCN[78]
 Yes No 91.4
-PointNet++qi2017pointnet++
+PointNet++[79]
 Yes No 91.9
-SO-Netli2018so
+SO-Net[80]
 Yes No 92.5
-PointVGGli2021pointvgg
+PointVGG[81]
 Yes No 93.6
-PointBERTyu2022point
+PointBERT[41]
 No Yes 93.8
-PointGPTchen2023pointgpt
+PointGPT[82]
 Yes Yes 94.9
@@
 No No 91.2
-The design of AllSpark is based on a general sequence-to-sequence architecture.
-As shown in Table 9, in the absence of both modality expert architectural designs and modal expert knowledge, AllSpark still outperforms classical networks with point cloud-specific structures (PointNet, Kd-net).
-It maintains comparability with the state-of-the-art model, PointGPT, which has both a modal expert structure and modal expert knowledge.
-This suggests that AllSpark holds significant potential for applications in the point cloud modality.
-4.3.7 Trajectory
-Table 10: Trajectory predicition with AllSpark
+The AllSpark module is designed based on a general sequence-to-sequence architecture.
+As shown in Table VIII, in the absence of both modality expert architectural designs and modal expert knowledge, AllSpark still outperforms classical networks with point cloud-specific structures (PointNet, Kd-Net).
+This approach maintains comparability with the state-of-the-art PointGPT model, which has both a modal expert structure and modal expert knowledge.
+This finding suggested that AllSpark has significant potential for applications in the point cloud modality.
+IV-C7 Trajectory
+For the trajectory modality, we evaluated the performance of AllSpark on the ETH dataset in the trajectory prediction task.
+This task involves predicting possible two-dimensional trajectories based on a set of two-dimensional coordinate points within a certain time period.
+We report accuracy using the average displacement error (ADE) and final displacement error (FDE)[37].
+Given the future trajectory$\{x_{t},y_{t}\}_{t=T_{bos}+1}^{T}$ (ground truth) and the predicted trajectory $\{\widehat{x_{t}},\widehat{y_{t}}\}_{t=T_{bos}+1}^{T}$, the ADE and FDE are used to measure their $L2$ distances, calculated as follows:
+${ADE}=\frac{1}{T_{pred}}\sum_{t=T_{bos}}^{T}\sqrt{\left(x_{t}-\widehat{x_{t}}\right)^{2}+\left(y_{t}-\widehat{y_{t}}\right)^{2}}$ (5)
+${FDE}=\sqrt{\left(x_{T}-\widehat{x_{T}}\right)^{2}+\left(y_{T}-\widehat{y_{T}}\right)^{2}}$ (6)
+In Table IX, AllSpark is compared with the state-of-the-art trajectory prediction models.
+The results indicate that AllSpark, which uses a unified structure without trajectory modality expert knowledge, outperforms most expert models.
+It achieves a prediction accuracy close to that of the SOTA model (STAR), with a difference of only 0.07 in the ADE metric and 0.11 in the FDE metric.
+This finding suggested that AllSpark demonstrated excellent adaptability to the trajectory modality.
+TABLE IX: trajectory prediction with allspark
 Method
@@
 0.43/0.69
-For the trajectory modality, we evaluated the performance of AllSpark on the ETH dataset in the trajectory prediction task.
-This task involves predicting possible two-dimensional trajectories based on a set of two-dimensional coordinate points within a certain time period.
-We report accuracy using Average Displacement Error (ADE) and Final Displacement Error (FDE)shi2023trajectory .
-Given the future trajectory$\{x_{t},y_{t}\}_{t=T_{bos}+1}^{T}$ (Ground Truth) and the predicted trajectory $\{\widehat{x_{t}},\widehat{y_{t}}\}_{t=T_{bos}+1}^{T}$, ADE and FDE are used to measure their $L2$ distances, calculated as follows:
-${ADE}=\frac{1}{T_{pred}}\sum_{t=T_{bos}}^{T}\sqrt{\left(x_{t}-\widehat{x_{t}}\right)^{2}+\left(y_{t}-\widehat{y_{t}}\right)^{2}}$ (5)
-${FDE}=\sqrt{\left(x_{T}-\widehat{x_{T}}\right)^{2}+\left(y_{T}-\widehat{y_{T}}\right)^{2}}$ (6)
-In Table 10, AllSpark is compared with the state-of-the-art trajectory prediction models.
-The results indicate that AllSpark, using a unified structure without trajectory modality expert knowledge, outperforms most expert models.
-It achieves a prediction accuracy close to the SOTA model (STAR) with a difference of only 0.07 in the ADE metric and 0.11 in the FDE metric.
-This suggests that AllSpark demonstrates excellent adaptability to the trajectory modality.
-4.3.8 SAR
-Table 11: SAR classification with AllSpark
+IV-C8 SAR
+For the SAR modality, the adaptability of AllSpark was tested on the MSTAR dataset, where the model is required to identify SAR images of ten military targets, and the metric used is the top-1 accuracy.
+In the experiment, the preprocessing of the MSTAR dataset followed the SOC settings from AConvNets[22].
+Table X presents the comparison results between AllSpark and the state-of-the-art model under these settings.
+Since AllSpark lacks expert knowledge of SAR image modalities and spatial invariance prior to changes in perspective; its performance on the MSTAR dataset is suboptimal, with an accuracy of 88%.
+This approach is comparable to the EMACH method.
+Enhancing the model’s adaptability to SAR images and improving its ability to extract spatial information are directions for future research.
+TABLE X: sar classification with allspark
 Method
-Accuracy (%)
-EMACHsrinivas2014sar
+Acc(%)
+EMACH[83]
 88
-SVMsrinivas2014sar
+SVM[83]
 90
-AdaBoostsrinivas2014sar
+AdaBoost[83]
 92
-MSRCdong2014sparse
+MSRC[84]
 93.6
-IGTsrinivas2014sar
+IGT[83]
 95
-MSSdong2015classification
+MSS[85]
 96.6
-Cond Gausso2001sar
+Cond Gauss[86]
 97
-M-PMCpark2014modified
+M-PMC[87]
 98.8
-AConvNetschen2016target
+AConvNets[22]
 99.13
@@
 88
-For SAR modality, AllSpark’s adaptability was tested on the MSTAR dataset, where the model is required to identify SAR images of ten military targets, and the metric used is Top1 Accuracy.
-In the experiment, the preprocessing of the MSTAR dataset followed the SOC (standard operating conditions) settings from AConvNetschen2016target .
-Table 11 presents the comparison results between AllSpark and the state-of-the-art model under these settings.
-Since AllSpark lacks SAR image modality expert knowledge and the spatial invariance prior to changes in perspective, its performance on the MSTAR dataset is suboptimal, with an accuracy of 88%.
-This is comparable to the EMACH method.
-Enhancing the model’s adaptability to SAR images and improving its ability to extract spatial information are directions for our future research.
-4.3.9 Infrared
-AllSpark’s adaptability to the infrared modality was tested on the SYSU-MM01 dataset in a cross-modal human re-identification task involving RGB and infrared modalities.
+IV-C9 Infrared
+AllSpark’s adaptability to the infrared modality was tested on the SYSU-MM01 dataset in a cross-modal human reidentification task involving RGB and infrared modalities.
 In this experiment, the model receives both an infrared image and an RGB image and is tasked with identifying and matching the same person across different modality images.
-Table 12 presents the comparison results between AllSpark and the state-of-the-art re-identification methods, with the evaluation metric being Rank at 20 accuracy.
-The results indicate that AllSpark (76.31%) achieved similar accuracy compared to HSME (77.95%).
-While our method did not reach the highest accuracy, it has already surpassed some expert models in RGB-Infrared re-identification (e.g., TONE, HCML) when using a unified architecture without modal expert knowledge, showcasing the potential of AllSpark in handling the infrared modality.
-Table 12: Visible-Infrared Re-identification with AllSpark
+Table XI compares AllSpark and the other state-of-the-art reidentification methods, with the evaluation metric being rank at an accuracy of 20.
+The results indicate that AllSpark (76.31%) achieved similar accuracy to that of HSME (77.95%).
+While our method did not reach the highest accuracy, it surpassed some expert models in RGB-infrared reidentification (e.g., TONE, HCML) when using a unified architecture without modal expert knowledge, demonstrating the potential of AllSpark in handling the infrared modality.
+TABLE XI: visible-infrared reidentification with allspark
 Method
 Rank-20(All-search)
-Two-streamwu2017rgb
+Two-stream[51]
 65.5
-One-streamwu2017rgb
+One-stream[51]
 66.74
-TONEye2018hierarchical
+TONE[88]
 68.6
-HCMLye2018hierarchical
+HCML[88]
 69.17
-Zero-Padwu2017rgb
+Zero-Pad[51]
 71.33
 AllSpark 76.31
-HSMEhao2019hsme
+HSME[89]
 77.95
-BDTRye2019bi
+BDTR[90]
 81.07
-DDAGye2020dynamic
+DDAG[38]
 95.81
-4.3.10 Graph
-For the Graph modality, AllSpark’s performance was evaluated on the traffic flow prediction task using the METR-LA dataset.
-The evaluation metrics include RMSE, MAE, and R2.
-Table 13 presents the comparison between AllSpark and the state-of-the-art results on the METR-LA dataset, with the baseline derived fromliu2023spatio .
-The results show that AllSpark, without the intervention of modal expert knowledge, is only 0.47 away from the best result in terms of RMSE, demonstrating AllSpark’s excellent adaptability to the Graph modality.
-Table 13: Traffic prediction with AllSpark
+IV-C10 Graph
+TABLE XII: traffic prediction with allspark
 Method
@@
 3.81 7.52 11.24
-4.3.11 Language
-Table 14: Text understanding with AllSpark
+For the graph modality, the performance of AllSpark was evaluated on the traffic flow prediction task using the METR-LA dataset.
+The evaluation metrics include the RMSE, MAE, and R2.
+Table XII compares AllSpark and the state-of-the-art methods on the METR-LA dataset, with the baseline derived from[39].
+The results show that AllSpark, without the intervention of modal expert knowledge, is only 0.47 away from the best result in terms of the RMSE, demonstrating its excellent adaptability to the graph modality.
+IV-C11 Language
+We tested AllSpark’s natural language processing capabilities on the IMDB dataset[43] with the task of binary sentiment classification (positive or negative).
+AllSpark loads weights from the Lynx; therefore, it can be considered to possess expert knowledge in the natural language modality.
+Compared with the SOTA models on the IMDB dataset, as shown in Table XIII, AllSpark outperformed most language models, with only a 0.32 difference from the SOTA result.
+This highlights AllSpark’s powerful understanding and analysis capabilities in natural language.
+TABLE XIII: text understanding with allspark
 Method
-Accuracy (%)
-RoBERTaliu2019roberta
+Acc(%)
+RoBERTa[91]
 95.3
-ULMFiThoward2018universal
+ULMFiT[92]
 95.4
-BERTxie2020unsupervised
+BERT[93]
 95.49
-Mixed VATsachan2019revisiting ; miyato2016adversarial
+Mixed VAT[94, 95]
 95.68
-LongFormerbeltagy2020longformer
+LongFormer[96]
 95.7
-XLNetyang2019xlnet
+XLNet[97]
 96.8
-ERNIE-Doc-Largeding2020ernie
+ERNIE-Doc-Large[98]
 97.1
@@
 96.78
-We tested AllSpark’s natural language processing capabilities on the IMDB datasetmaas2011learning with the task of binary sentiment classification (positive or negative).
-AllSpark loaded weights from the Lynx, and therefore, it can be considered to possess expert knowledge in the natural language modality.
-When compared with the state-of-the-art models on the IMDB dataset, as shown in Table 14, AllSpark outperformed most language models, with only a 0.32 difference from the SOTA result.
-This highlights AllSpark’s powerful understanding and analysis capabilities in natural language.
-4.3.12 Video
+IV-C12 Video
 For the video modality, we evaluated AllSpark’s performance on action recognition tasks using the UCF101 dataset.
-The model is tasked with understanding videos and accurately classifying them into one of the 101 classes, with the evaluation metric being Top1 accuracy.
-In Table 15, we compared AllSpark with the current state-of-the-art models.
-Currently, AllSpark’s adaptation to the Video modality is not optimal, showing a significant gap compared to baseline model results.
+The model is tasked with understanding videos and accurately classifying them into one of the 101 classes, with the evaluation metric being the top-1 accuracy.
+In Table XIV, we compare AllSpark with the current state-of-the-art models.
+Currently, AllSpark’s adaptation to the video modality is not optimal, as it shows a significant difference from the baseline model results.
 We attribute this to two main reasons: 1.
-The high redundancy in video information increases the training cost for AllSpark, and we trained it for only 3 epochs on the UCF101 dataset.
+The high redundancy in video information increases the training cost for AllSpark.
+We trained it for only 3 epochs on the UCF101 dataset.
 2.
-AllSpark’s model architecture lacks flexibility for three-dimensional data, making it less effective in capturing temporal information.
-Table 15: Video classification with AllSpark
+AllSpark’s model architecture lacks flexibility for three-dimensional data, making it less effective at capturing temporal information.
+TABLE XIV: video classification with allspark
 Method
-Accuracy (%)
+Acc(%)
 OPN
@@
 27.5
-4.3.13 Oblique Photography
+IV-C13 Oblique Photography
 For the oblique photography modality, we tested AllSpark’s performance on the 3D reconstruction task using the WHU-OMVS dataset.
 The model takes five-view images as input, and the goal is to output depth maps for reconstructing 3D models.
-The evaluation metric used is the Percentage of Accurate Grids in Total (PAG)liu2023deep , calculated by the following formula:
+The evaluation metric used was the percentage of accurate grids in total (PAG)[42], calculated by the following formula:
 $PAG_{a}=\left(\frac{m_{a}}{m}\right)*100$ (7)
-The suffixes in PAG represent different accuracy standards, where ${PAG}_{6}$ signifies an error within 0.6 meters, and ${PAG}_{10}$ indicates an error within 1 meter.
-Table 16 presents the comparison of AllSpark with popular multi-view 3D reconstruction models on the WHU-OMVS dataset.
+The suffixes in the PAG represent different accuracy standards, where ${PAG}_{6}$ signifies an error within 0.6 metres and ${PAG}_{10}$ indicates an error within 1 metre.
+Table XV compares AllSpark with popular multiview 3D reconstruction models on the WHU-OMVS dataset.
 AllSpark falls short in terms of accuracy compared to modality-specific expert models.
@@
 The model architecture lacks flexibility and cannot connect gradual features like expert 3D reconstruction models such as Ada-MVS.
-Additionally, it lacks the ability to design specific model structures for processing, severely restricting its performance.
-Adapting multi-modal large models to dense prediction tasks and optimizing the architecture is one of our future research directions.
-Table 16: 3D reconstruction with AllSpark
+Additionally, these methods lack the ability to design specific model structures for processing, severely restricting their performance.
+Adapting multimodal large models to dense prediction tasks and optimizing the architecture are future research directions.
+TABLE XV: 3D reconstruction with allspark
 Method
@@
 6.4 10.4
-5 Discussion
-5.1 Limitation
-Certainly, our work has some limitations, and these will guide our future research directions:
+V Discussion
+V-A Limitations
+Certainly, our work has several limitations, which will guide our future research directions:
   1. 1.
     Lack of interaction between different modalities.
-While AllSpark incorporates interaction between infrared and RGB modalities, for other modalities, it essentially uses a unified model for multiple single-modal tasks.
-This is primarily due to constraints such as the lack of multi-modal paired data, the inflexibility of the model architecture, and the limited richness of multi-modal interaction tasks.
+While AllSpark incorporates interactions between infrared and RGB modalities, for other modalities, it essentially uses a unified model for multiple single-modal tasks.
+This is primarily due to constraints such as the lack of multimodal paired data, the inflexibility of the model architecture, and the limited richness of multimodal interaction tasks.
 In the future, we plan to explore and work more extensively in these directions.
   2. 2.
-    Lack of in-depth research on multi-modal alignment methods.
-While we introduced the basic principles of the Language as Reference Framework (LaRF) and aligned all modalities to the language modality, our approach included only introducing the modality bridge from Lynx as the alignment method.
-We haven’t conducted in-depth research into the various methods of multi-modal alignment.
-Moreover, our testing focused on the model’s adaptability to different modalities, and we haven’t explored potential capabilities such as interactive reasoning, zero-shot learning, and multi-modal emergence that the model may possess after alignment with language.
-In the future, we plan to delve deeper into researching multi-modal alignment methods, and we will design additional tasks and experiments to explore and reveal the upper limits and boundaries of large model reasoning capabilities.
+    There is a lack of in-depth research on multimodal alignment methods.
+While we introduced the basic principles of the LaRF and aligned all the modalities to the language modality, our approach included only introducing the modality bridge from Lynx as the alignment method.
+We have not conducted in-depth research into the various methods of multimodal alignment.
+Moreover, our testing focused on the model’s adaptability to different modalities, and we have not explored potential capabilities such as interactive reasoning, zero-shot learning, or multimodal emergence that the model may possess after alignment with language.
+In the future, we plan to delve deeper into researching multimodal alignment methods, and we will design additional tasks and experiments to explore and reveal the upper limits and boundaries of large model reasoning capabilities.
   3. 3.
     Initial work.
-Our current work is still in its initial exploratory phase, and we haven’t carefully refined AllSpark’s adaptability and performance on each modality.
-As a result, the model exhibits suboptimal performance on certain modalities, such as oblique photography, SAR, and video.
+Our current work is still in its initial exploratory phase, and we have not carefully refined AllSpark’s adaptability and performance on each modality.
+As a result, the model exhibits suboptimal performance on certain imaging modalities, such as oblique photography, SAR, and video.
 In the future, we plan to conduct more refined and targeted adjustments for each modality to enhance overall performance.
@@
 We generalize their applicability to other modalities by utilizing modality bridges to project other modalities onto the language modality.
-Although we freeze most parameters, fine-tuning even once on 2 A6000 GPUs often requires more than a day.
-Given the increasing training costs for large models, exploring methods to generalize their universal reasoning abilities remains one of our future research directions.
+Although we freeze most of the parameters, fine-tuning even once on 2 A6000 GPUs often requires more than a day.
+Given the increasing training costs for large models, exploring methods to generalize their universal reasoning abilities is one of our future research directions.
   5. 5.
     Interesting phenomenon.
-During our exploratory experiments, we discovered some interesting properties of large models.
+During our exploratory experiments, we discovered several interesting properties of large models.
 For instance, spatial information tends to degrade in large models, leading to collapse when performing dense prediction tasks such as segmentation and detection.
-Additionally, these models struggle to optimize when transferred to a small amount of downstream data.
+Additionally, these models struggle to optimize when transferred to a small quantity of downstream data.
 Large models often require smaller hyperparameters and are sensitive to them.
-These observations might partially reveal the working mechanisms of large models, and we plan to conduct more in-depth investigations into these phenomena in the future.
-5.2 Potential of the LaRF
-Inspired by the human cognitive system and linguistic philosophy, we propose ”Language as Reference Framework (LaRF)” as the first principle for constructing our unified multimodal model.
+These observations might partially reveal the working mechanisms of large models, and we plan to conduct additional in-depth investigations into these phenomena in the future.
+V-B Potential of the LaRF
+Inspired by the human cognitive system and linguistic philosophy, we propose the ”LaRF” as the first principle for constructing our unified multimodal model.
 Its foreseeable potential includes at least the following three points:
@@
 Currently, the computational power and data scale required for training large models are rapidly expanding, and even the cost of fine-tuning large models is becoming prohibitive.
-Therefore, in the future, training large models for every domain is almost an impossible task.
+Therefore, in the future, training large models for every domain will be almost impossible.
 Language, however, holds the potential for achieving efficient generalization of large models, expanding them from their native domains to additional domains at minimal cost.
-In our proposed AllSpark, we have designed simple text prompts and task heads for each modality, demonstrating significant potential for multimodal expansion.
-In theory, AllSpark can extend to any number of modalities.
+With our proposed AllSpark, we designed simple text prompts and task heads for each modality, demonstrating significant potential for multimodal expansion.
+In theory, AllSpark, built on the LaRF principle, can be extended to arbitrary modalities.
 In the future, we will conduct more in-depth research on the impact of text prompts and lightweight parameter modules on the generalization of large models.
@@
 Deep learning models have often been referred to as ”black boxes,” indicating that the reasoning process of these models is invisible and challenging to interpret.
-Some research on the interpretability of deep learning models is often based on complex mathematical models and numerous assumptions, greatly limiting the practical applications of deep learning methods in fields such as clinical medicine, military, and national resources where low fault tolerance or high confidentiality is crucial.
-However, language, as a tool for human thought and communication, provides models based on LaRF with the potential to directly use natural language for outputting reasoning chains and justifications.
+Research on the interpretability of deep learning models often relies on complex mathematical models and numerous assumptions, greatly limiting the practical application of deep learning methods in fields such as clinical medicine, military, and national resources where low fault tolerance or high confidentiality are crucial.
+However, language, as a tool for human thought and communication, provides models based on LaRF with the potential to use natural language directly for outputting reasoning chains and justifications.
   3. 3.
-    Transition from end-to-end to interactive paradigm.
+    Transition from an end-to-end to an interactive paradigm.
 The end-to-end paradigm refers to the learning approach where the model takes input and directly outputs results.
 In recent years, the end-to-end paradigm has become increasingly popular due to its simple and clear architecture and excellent performance.
-However, it also has clear disadvantages, such as uncontrollable internal operations, the need to optimize the whole for certain problems, and difficulty in pinpointing the cause of issues.
-A LaRF-based architecture has the potential to achieve an interactive paradigm, where users input raw data and corresponding text prompts, and the model automatically performs relevant operations based on the prompts.
+However, this approach also has clear disadvantages, such as uncontrollable internal operations, the need to optimize the whole for certain problems, and difficulty in pinpointing the cause of issues.
+A LaRF-based architecture has the potential to achieve an interactive paradigm in which users input raw data and corresponding text prompts and the model automatically performs relevant operations based on the prompts.
 Users can even iteratively adjust the text prompts based on the results.
-Therefore, in terms of both performance and controllability, the interactive paradigm has advantages that are incomparable to the end-to-end paradigm.
-6 Conclusion
+Therefore, in terms of both performance and controllability, the interactive paradigm has advantages that are incomparable to those of the end-to-end paradigm.
+VI Conclusion
 Leveraging multimodal data is an inherent requirement for intelligent models to achieve geographic object cognition.
-Inspired by human cognitive systems and linguistic philosophy, we propose that the construction of multimodal models should follow the fundamental principle of Language as Reference Framework (LaRF).
-Guided by this principle, we use language to balance the cohesion and autonomy of modalities, presenting a unified intelligent model, AllSpark, encompassing 13 spatiotemporal modalities.
-Experimental results demonstrate that AllSpark exhibits excellent adaptability and application potential across various spatiotemporal modalities, highlighting the feasibility and potential of constructing multimodal models with LaRF.
-AllSpark remains an initial exploratory work, and in the future, we aim to delve deeper into mechanisms guided by natural language, efficient generalization of large models, and the transition to an interactive paradigm.
+Inspired by human cognitive systems and linguistic philosophy, we propose that the construction of multimodal models follow the fundamental principle of Language as Reference Framework (LaRF).
+Guided by this principle, we use language to balance the cohesion and autonomy of modalities, presenting a unified intelligent model, AllSpark, encompassing thirteen spatio-temporal modalities.
+The experimental results demonstrated that AllSpark exhibited excellent adaptability and application potential across various spatio-temporal modalities, highlighting the feasibility and potential of constructing multimodal models with LaRF.
+AllSpark remains an initial exploratory work, and in the future, we aim to delve deeper into the mechanisms guided by natural language, the efficient generalization of large models, and the transition to an interactive paradigm.
+TABLE XVI: List of prompts
+Dimension
+Modal Prompt
+1D Text 1.Please determine if this movie review is positive or negative?
+Code -
+2D RGB 1.This remote sensing image belongs to which of the following categories: [Category list of the NWPU dataset] 2.Describe this remote sensing image briefly 3.Find a word that is most relevant to this remote sensing image 4.Describe the key elements in this remote sensing image
+MSI 1.Based on the multi-spectral imagery feature description, please classify this object 2.Given the following multi-spectral imagery characteristics, please output the most fitting scene label
+HSI 1.Given the spectral information, can you help determine which class this pixel belongs to?
+2.Here is the spectral data for a pixel.
+Considering the typical characteristics of land cover classes, could you provide a detailed analysis and suggest the most likely class for this pixel?
+3.The spectral information for a pixel is given, but the data is noisy.
+Given the potential variability, which land cover classes should be considered as possible candidates for this pixel?
+Table 1.Please utilize the provided air quality indicators to accurately predict the concentrations of PM2.5 in the atmosphere 2.Given the air quality indicators, please provide a prediction for the PM2.5 levels at this particular moment
+Oblique Photography 1.Please reconstruct the 3D model from the given images of five different views.
+Trajectory 1.Based on their past positions and movements in a crowded environment, predict the future trajectory of a selected pedestrian.
+2.Using the pedestrian trajectory data, along with additional information about the surrounding environment, predict the future path of the pedestrian.
+3.Given the current and past positions of a pedestrian and their neighboring pedestrians, predict the main pedestrian’s trajectory
+SAR 1.Based on the SAR imagery feature description, please classify this object 2.Given the following SAR imagery characteristics, please output the most fitting scene label
+Infrared 1.Given an infrared image of a person, find and highlight the same person in a set of visible light images, paying close attention to features like clothing, posture, and gait 2.Analyze the characteristics of a person in a visible light image, such as clothing texture, color, and body shape.
+Then, locate the person with matching features in a series of infrared images
+Graph 1.Given the current traffic data including vehicle flow rate, average speed, and time of day from the METRLA dataset, predict the traffic flow 2.Analyze the historical data on vehicle speeds and flow rates from the METRLA dataset for the past week.
+Identify any patterns or trends and predict the traffic conditions
+3D Video 1.Please analyze this video and describe the characteristics of the action in detail 2.Watch this short clip and provide a step-by-step description of the action sequence
+Point Cloud 1.Classify the provided point cloud sample into the correct category 2.Look at the point cloud data characteristics and classify the object 3.Please analyze the given point cloud dataset and determine which category it belongs to.
+Focus on the shape and structure evident in the point cloud

### v3

@@
-AllSpark[Uncaptioned image]: A Multimodal Spatio-Temporal General Intelligence Model with Thirteen Modalities
+AllSpark[Uncaptioned image]: A Multimodal Spatio-Temporal General Intelligence Model with Ten Modalities via Language as a Reference Framework
 Abstract
-RGB, infrared, multispectral, and other spatio-temporal modal data fundamentally represent different observational approaches for the same geographic object.
+RGB, multispectral, point and other spatio-temporal modal data fundamentally represent different observational approaches for the same geographic object.
 Therefore, leveraging multimodal data is an inherent requirement for comprehending geographic objects.
-However, due to the high heterogeneity in structure and semantics among various spatio-temporal modal data, the joint interpretation of multimodal spatio-temporal data has long been an extremely challenging problem.
+However, due to the high heterogeneity in structure and semantics among various spatio-temporal modalities, the joint interpretation of multimodal spatio-temporal data has long been an extremely challenging problem.
 The primary challenge resides in striking a trade-off between the cohesion and autonomy of diverse modalities.
@@
 Inspired by the human cognitive system and linguistic philosophy, where perceptual signals from the five senses converge into language, we introduce the Language as Reference Framework (LaRF), a fundamental principle for constructing a multimodal unified model.
-Building upon this, we propose a multimodal spatio-temporal general artificial intelligence model called AllSpark.
-Our model integrates thirteen different modalities into a unified framework, including one-dimensional (text, code), two-dimensional (RGB, infrared, SAR, multispectral, hyperspectral, tables, graphs, trajectory, oblique photography), and three-dimensional (point clouds, videos) modalities.
-To achieve modal cohesion, AllSpark uniformly maps diverse modal features to the language modality by a modal bridge and a multimodal large language models.
-To maintain modality autonomy, AllSpark introduces modality-specific encoders to extract the tokens of various spatio-temporal modalities.
-Finally, observing a gap between the model‘s interpretability and downstream tasks, we designed modality-specific prompts and task heads to enhance the model’s generalization capability on specific downstream tasks.
-Experiments indicate that AllSpark, without expert knowledge of most spatio-temporal modalities and utilizing a unified structure, achieves competitive accuracy in modalities such as RGB and trajectory compared to state-of-the-art models.
-Moreover, AllSpark shows excellent adaptability in various modalities, such as MSI, HSI, PointCloud, table, code, and graph.
-AllSpark demonstrates the potential and possibility of constructing general artificial intelligence with a large language model.
+Building upon this, we propose AllSpark, a multimodal spatio-temporal general artificial intelligence model.
+Our model integrates ten different modalities into a unified framework, including one-dimensional (language, code, table), two-dimensional (RGB, SAR, multispectral, hyperspectral, graph, trajectory), and three-dimensional (point cloud) modalities.
+To achieve modal cohesion, AllSpark introduces a modal bridge and multimodal large language model (LLM) to map diverse modal features into the language feature space.
+To maintain modality autonomy, AllSpark uses modality-specific encoders to extract the tokens of various spatio-temporal modalities.
+Finally, observing a gap between the model’s interpretability and downstream tasks, we designed modality-specific prompts and task heads, enhancing the model’s generalization capability across specific tasks.
+Experiments indicate that the incorporation of language enables AllSpark to excel in few-shot classification tasks for RGB and point cloud modalities without additional training, surpassing baseline performance by up to 41.82%.
+Additionally, AllSpark, despite lacking expert knowledge in most spatio-temporal modalities and utilizing a unified structure, demonstrates strong adaptability across ten modalities.
 LaRF and AllSpark contribute to the shift in the research paradigm in spatio-temporal intelligence, transitioning from a modality-specific and task-specific paradigm to a general paradigm.
+The source code is available at https://github.com/GeoX-Lab/AllSpark.
 Index Terms:
@@
 I Introduction
-Figure 1: We propose ”Language as Reference Framework (LaRF)” as the fundamental principle for constructing multimodal models.
-Under the guidance of this principle, we introduced a multimodal general intelligence model, AllSpark, which integrates thirteen spatio-temporal modalities.
-Benefiting from the increasingly diverse observational methods available for spatio-temporal scenes, geographic objects can be described using various spatio-temporal modalities, such as RGB, infrared, SAR, multispectral, graph, and spatio-temporal trajectory data.
+TABLE I: AllSpark integrates ten spatio-temporal modalities
+Model 1D 2D 3D
+Language Code Table RGB MSI HSI SAR Traj Graph Point
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Figure 1: AllSpark demonstrates excellent adaptability across up to 10 heterogeneous modalities and shows outstanding few-shot learning capabilities in RGB and point cloud modalities.
+Benefiting from the increasingly diverse observational methods available for spatio-temporal scenes, geographic objects can be described by various spatio-temporal modalities, such as RGB, SAR, multispectral, graph, point cloud and trajectory data [7, 8, 9].
 Each modality provides unique information about different aspects of the geographic object.
@@
 In terms of semantics, RGB imagery reflects the electromagnetic characteristics of visible light bands emitted and reflected by geographic objects, whereas SAR imagery reflects the electromagnetic characteristics of microwaves emitted actively by radar for scattering from geographic objects.
-For a long time, constrained by the high heterogeneity across various modalities mentioned above, researchers have often developed specific methods based on prior assumptions related to a particular modality or designed multimodal approaches based on a few low-heterogeneity modalities.
+For a long time, constrained by the high heterogeneity across various modalities mentioned above, researchers have often developed specific methods based on prior assumptions related to a particular modality or designed multimodal approaches for a few low-heterogeneity modalities.
 For instance, in single-modal research, C.
@@
 introduced the transformer[11], which focuses on the long-range dependencies within word sequences.
-For the graph modality, Kipf and Welling proposed the GCN[3] based on the adjacency relationships between nodes in a graph.
+For the graph modality, Kipf and Welling proposed the GCN[12, 13] based on the adjacency relationships between nodes in a graph.
 In multimodal research, the fusion of optical and SAR imagery has been widely explored in both traditional and deep learning remote sensing[14].
@@
 We believe that the key challenge in addressing this issue lies in striking a trade-off between the cohesion and autonomy of diverse modalities.
-Cohesion refers to the presence of mutually correlated shared information among modalities, while autonomy signifies the existence of unique information specific to each modality relative to others.
-If we simply project data from multiple modalities into a common representation space for fusion, emphasizing the cohesion between modalities, this approach may lead to the loss of modality-specific information, thereby diluting the essence of multimodal collaboration.
+In our paper, ”cohesion” refers to the presence of mutually correlated shared information among modalities.
+For instance, both RGB and SAR images may describe the contour of the same object.
+”Autonomy,” on the other hand, refers to the existence of unique information specific to each modality relative to others.
+For example, the RGB modality can describe an object’s color and texture, while the SAR modality can capture the object’s scattering properties in relation to radar waves.
+Cohesion forms the foundation for the interrelation between modalities, while autonomy highlights the value of multimodal joint interpretation—gaining a complete understanding of an object by integrating multiple modalities.
+If we merely project data from different modalities into a shared representation space to emphasize inter-modality cohesion, this approach risks losing modality-specific information, ultimately undermining the unique contributions of each modality and weakening the core value of multimodal collaboration.
 In contrast, if we excessively stress the autonomy between modalities, it may hinder the establishment of connections among them, limiting the model’s ability to simultaneously perceive multiple modalities.
@@
 Inspired by this, we propose the Language as Reference Framework (LaRF) as a fundamental principle for constructing multimodal models.
-As illustrated in Figure 1, the abstract concepts derived from each modality should align with language, allowing for joint interpretation in the unified representation space of language.
-Building upon this, we propose a multimodal spatio-temporal general artificial intelligence model, AllSpark, that integrates thirteen different modalities into a unified framework, including one-dimensional (text, code), two-dimensional (RGB, infrared, SAR, multispectral, hyperspectral, tables, graphs, trajectory, oblique photography), and three-dimensional (point clouds, videos) modalities.
-To achieve modal cohesion, AllSpark uniformly maps diverse modal features to the language modality.
-In addition, we designed modality-specific prompts to guide multimodal large language models in accurately perceiving multimodal data.
+It means that the abstract concepts derived from each modality should align with language, enabling joint interpretation in the unified representation space of language.
+Building upon this, we propose a multimodal spatio-temporal general intelligence model [20], AllSpark, that integrates ten different modalities into a unified framework, including one-dimensional (language, code, table), two-dimensional (RGB, SAR, multispectral, hyperspectral, graph, trajectory), and three-dimensional (point cloud) modalities.
+As shown in Table I and Figure 1, previous work either overlooked some important spatio-temporal modalities, such as hyperspectral and trajectory, or focused solely on natural images without considering remote sensing imagery.
+AllSpark covers a broader range of spatio-temporal modalities, such as multispectral, hyperspectral, graph, trajectory, and more, while demonstrating excellent few-shot learning capabilities and modality adaptability.
+To achieve modal cohesion, AllSpark uniformly maps diverse modal features to the language feature space.
 To maintain the autonomy between modalities, AllSpark introduces specific modal encoders for each modality to extract independent tokens.
 Given the high heterogeneity among modality data and modality encoders, a significant dimensional gap exists between the tokens of each modality and the language modality.
-To address this issue, we introduce a modality bridge, a mechanism from Lynx[10], to accomplish dimensional mapping from each modality’s tokens to the language tokens.
-Finally, considering the existing gap between the interpretability of the multimodal large language model and the specific downstream tasks, we design task heads for each downstream task to enhance the model’s generalization capability.
-Given the powerful interpretability capabilities of the multimodal large language model, we adhere to a lightweight design principle in both the modal encoders and task heads.
-Experiments demonstrate that AllSpark achieves competitive accuracy without expert knowledge in most modalities and utilizing a unified structure in modalities such as RGB and spatio-temporal trajectories compared to state-of-the-art models.
+To address this issue, we introduce a modality bridge, a mechanism from perceiver[21], to accomplish dimensional mapping from each modality’s tokens to the language tokens [22].
+Finally, considering the existing gap between the interpretability of the multimodal large language model and the specific downstream tasks, we design task heads and modality-specific text prompts for each downstream task to enhance the model’s generalization capability.
+Given the powerful interpretability capabilities of the multimodal large language model, we adhere to a lightweight design principle in task heads.
+Experiments demonstrate that AllSpark, despite lacking expert knowledge in most spatio-temporal modalities and utilizing a unified structure, achieves competitive accuracy in modalities such as RGB and spatio-temporal trajectories compared to state-of-the-art models.
 Specifically, in the RGB modality, the accuracy of AllSpark is only 0.84 lower than that of the SOTA model, and in the trajectory modality, the average displacement error (ADE) metric differs by only 0.07 compared to that of the SOTA model.
-Additionally, AllSpark exhibits excellent adaptability in various other modalities, including point clouds, multispectral, hyperspectral, tables, graphs, and code.
+Additionally, AllSpark exhibits excellent adaptability in various other modalities, including point cloud, multispectral, hyperspectral, table, graph, and code.
 Theoretically, our proposed model has the potential for seamless extension to an arbitrary number of modalities.
 In other words, our contributions can be summarized as follows:
-  • We first propose a unified multimodal spatio-temporal general model, AllSpark, that successfully integrates thirteen spatio-temporal modalities into a single model.
+  • We first propose a unified multimodal spatio-temporal general model, AllSpark, that successfully integrates ten spatio-temporal modalities into a single model.
   • Inspired by the human cognitive system and linguistic philosophy, we propose the language as reference framework (LaRF), which offers a novel solution to balance cohesion and autonomy among multiple modalities.
-This principle provides a fresh approach to multimodal unification.
-  • We propose new approaches for designing multimodal models in the spatio-temporal domain.
-AllSpark demonstrates the potential and ability to construct a general model with large language models, thereby contributing to the shift in the research paradigm in spatio-temporal intelligence, transitioning from a modality-specific and task-specific paradigm to a general paradigm.
-  • Experiments indicate that AllSpark, without expert knowledge of the majority of spatio-temporal modalities and utilizing a unified structure, achieves competitive accuracy in modalities such as RGB and trajectory information compared to state-of-the-art models.
-Moreover, AllSpark shows excellent adaptability in various modalities, such as MSI, HSI, PointCloud, table, code, and graph.
+  • Experiments indicate that AllSpark demonstrates strong few-shot learning capabilities and supports ten modalities.
 Theoretically, our proposed model has the potential for seamless extension to an arbitrary number of modalities.
@@
 II-A Single-Modal Model
-For one-dimensional modalities, we focus on code and text.
-Code and text are both fundamental language modalities, but due to the excellent characteristics of code, such as strict syntax, unambiguous nature, and ability to interact with machines, code is treated as separate modality.
+For one-dimensional modalities, we focus on code, language and table.
+Given the excellent characteristics of code, such as strict syntax, unambiguous nature, and ability to interact with machines, code is treated as a separate modality.
 Z.
@@
 pretrained a model, CodeBERT[23], which facilitates the mutual transfer of information between code and natural language modalities.
-For the text modality, landmark contributions include a transformer[2], BERT[12], and the GPT series [13, 14, 15, 16], which have inspired subsequent series of works.
-We categorize RGB, multispectral, hyperspectral, infrared, SAR, table, graph, trajectory, and oblique photography as two-dimensional (2D) modalities.
+For the language modality, landmark contributions include a transformer[11], BERT[24], and the GPT series [25, 26, 27, 28], which have inspired subsequent series of works.
+The table is one of the commonly used modalities for recording and expressing information.
+TabNet[29], proposed by S.
+Ö.
+Arik et al., employs a sequence attention mechanism to achieve feature selection in the Table modality, thereby enabling interpretable and more efficient learning.
+We categorize RGB, multispectral, hyperspectral, SAR, graph, trajectory as two-dimensional (2D) modalities.
 Among the two-dimensional modalities, standard three-channel RGB images are among the most common.
 For the RGB modality, the ResNet[30] proposed by K.
-He et al., which is based on the importance of visual global and local information, and the visual transformer (ViT)[18] introduced by A.
+He et al., which is based on the importance of visual global and local information, and the vision transformer (ViT)[31] introduced by A.
 Dosovitskiy et al., which leverages a global attention mechanism, represent two landmark contributions.
@@
 introduced the R-3D-CNN[33] to further enhance the extraction of spectral features.
-Beyond an increase in the number of bands, attention has been given to images captured in specific spectral bands.
-For instance, infrared images, widely applied in fields such as agriculture and oceanography, have gained significant attention.
-H.
-Chen et al.
-proposed a shallow convolutional neural network that integrates a decision tree to meet the demand for real-time intelligent interpretation of a large volume of infrared images in agricultural water resource monitoring[21].
-In the case of SAR images formed by active microwave radar, S.
-Chen et al.
-introduced AConvNet[22], a widely used fully convolutional neural network for intelligent SAR image interpretation.
-For oblique photography, 3D reconstruction aims to utilize single or multiple two-dimensional images from various viewpoints to reconstruct a three-dimensional model.
-Y.
-Yao et al.
-addressed this task by proposing MVSNet[23].
-By utilizing techniques such as three-dimensional convolution, the model establishes correlations between images from different perspectives, thereby reconstructing a three-dimensional model of the target from the features extracted from two-dimensional images.
-Beyond the visual domain, the table is one of the commonly used modalities for recording and expressing information.
-TabNet[24], proposed by S.
-Ö.
-Arik et al., employs a sequence attention mechanism to achieve feature selection in the Table modality, thereby enabling interpretable and more efficient learning.
+In the case of SAR images formed by active microwave radar, S. Chen et al. introduced AConvNet[34], a widely used fully convolutional neural network for intelligent SAR image interpretation.
 The trajectory modality reflects the temporal changes in the spatial positions of objects.
@@
 proposed the GAT[36], which incorporates attention mechanisms into the graph modality.
-Finally, we turn our attention to three-dimensional modalities: video and point clouds.
-The video modality is formed by stacking images over time.
-A model proposed by A.
-Karpathy et al.
-extends the connectivity of CNN models over time.
-Additionally, multiresolution techniques are utilized to accelerate the training process, effectively reducing the impact of spatio-temporal redundancy on model efficiency in the video modality[27].
+Finally, we turn our attention to three-dimensional modalities: point cloud.
 The point cloud modality captures information about the position, shape, colour, texture, and other aspects of three-dimensional objects.
@@
 Recognizing the intrinsic requirement for intelligent models to utilize multimodal data for geographic object cognition, numerous researchers have endeavoured to balance the cohesion and autonomy among modalities to construct multimodal models.
-For joint interpretation of two modalities, H.
-Li and X.-J.
-Wu recognized the enhancement possible through fusing RGB and infrared images, which convey information about colour, texture, and temperature.
-Given the widespread demand and application prospects in military, medical, and other fields, they proposed a dense fusion method for RGB and infrared images: DenseFuse[29].
-This method, employing a dense module that connects the outputs of all convolutional layers, successfully achieves pixel-level fusion of the RGB and infrared modalities.
 A.
@@
 proposed a dual-stream convolutional network that uses high-resolution multispectral images to enhance the spatial resolution of hyperspectral images[41].
+To achieve joint interpretation of multimodal data, traditional multimodal models typically design specific architecture based on the priors of certain modalities.
+For example, Hang et al.
+proposed Coupled CNNs [42], which consist of two CNN networks to extract spectral-spatial features from hyperspectral data and elevation information from LiDAR data.
+They ultimately use both feature-level and decision-level fusion methods to integrate the heterogeneous features of the two modalities.
+Similarly, Zhang et al.
+proposed SLA-Net [43], which designs specific network structures to extract and fuse spatial information and morphological characteristics from hyperspectral imagery.
+These methods are typically designed for specific purposes and modalities, making it difficult to extend them to more modalities.
+Additionally, Gao et al.
+proposed DFINet [44], which extracts self-correlation and cross-correlation between multimodal data to deeply fuse hyperspectral and multispectral modality features.
+Likewise, Hong et al.
+introduced S2FL [45], which extracts Modality-Specific Subspaces for each modality and a Shared Subspace for all modalities, finally obtaining multimodal interpretation results through a unified projection.
+The issue with such methods is the lack of a unified alignment reference, limiting them to collaboration between a few modalities.
+As the number of modalities increases, the complexity of aligning multiple modalities will grow exponentially.
 Notably, CLIP[15], proposed by A.
-Radford et al., associates the RGB modality with the text modality using contrastive learning.
+Radford et al., associates the RGB modality with the text modality using contrastive learning [46, 47].
 With pretraining guided by weak supervisory signals from text, CLIP has demonstrated outstanding capabilities in both visual single-modal tasks and visual-language multimodal tasks, inspiring a series of subsequent works[16, 17, 18, 19].
@@
 introduced a meta-transformer[6], leveraging the contrastive learning paradigm from CLIP to pretrain a universal backbone network under the visual-language modality.
-It exhibits multimodal generalization abilities across various modalities, such as point clouds, infrared, and hyperspectral data.
+It exhibits multimodal generalization abilities across various modalities, such as point cloud, infrared, and hyperspectral data.
 J.
@@
 The success of these approaches implies the unique role of language modalities in multimodal models.
-Therefore, building upon the aforementioned efforts, we systematically propose the fundamental principle of the LaRF.
-Guided by this principle, we balance cohesion and autonomy among diverse modalities and introduce a general intelligent model named AllSpark, which unifies thirteen spatio-temporal modalities and possesses the potential to extend to an arbitrary number of modalities.
+Building upon the aforementioned efforts, we systematically propose the fundamental principle of the LaRF.
+Guided by this principle, we balance cohesion and autonomy among diverse modalities and introduce a general intelligent model named AllSpark, which unifies ten spatio-temporal modalities and possesses the potential to extend to an arbitrary number of modalities.
 III Method
-III-A Theory
+III-A Language as Reference Framework
 We observe that in the process of comprehending the world, humans integrate information from multiple modalities, such as hearing, touch, smell, and vision.
@@
 In other words, language precisely encodes human perception and understanding of the world, providing clear definitions and meanings to abstract concepts from each modality.
-Inspired by our observation, to balance cohesion and autonomy among multiple modalities, we introduce the fundamental principle of Language as Reference Framework (LaRF).
-Figure 2: Guided by the LaRF principle, multimodal data are transformed into a token-context structure similar to language under their respective prior assumptions while maintaining the autonomy of the multimodalities and achieving cohesion among modalities, thus enabling interpretation of multimodal data in the unified language representation space.
+Inspired by our observation, we introduce the fundamental principle of Language as Reference Framework (LaRF) to balance cohesion and autonomy among multiple modalities.
+Figure 2: Guided by the LaRF principle, multimodal data are transformed into a token-context structure akin to language, based on their respective prior assumptions.
+This approach preserves the autonomy of each modality while achieving cohesion between them, enabling the interpretation of multimodal data within a unified language representation space.
 In terms of the cohesion of multimodalities, the high heterogeneity between multiple spatio-temporal modalities is a major challenge, while the LaRF principle defines the alignment anchor between multimodalities as language explicitly.
@@
 Therefore, we can align highly heterogeneous spatio-temporal modalities to language modalities in structure and semantics, enabling multimodal interpretation in the unified representation space of language.
-In addition, the important guiding role of natural language prompts is also one of the reasons why the LaRF principle can achieve cohesion among modalities.
+Additionally, the pivotal role of natural language prompts is a key factor in the LaRF principle’s ability to achieve cohesion across modalities.
 In contrast, we can independently encode multiple spatio-temporal modalities into token sequences under their respective prior assumptions, so the LaRF principle does not lead to the loss of modal autonomy.
@@
   3. 3.
-    Interpretability: Deep learning methods have long been characterized as opaque ”black boxes.” However, a multimodal intelligent system constructed based on the LaRF can directly leverage language as a tool.
+    Interpretability: Deep learning methods have long been characterized as ”black boxes.” However, a multimodal intelligent system constructed based on the LaRF can directly leverage language as a tool.
 This facilitates the output of interpretable reasoning chains that humans can understand, thereby achieving true explainable artificial intelligence.
@@
 Multimodal data are extracted by their respective modal encoders into token sequences.
-After undergoing dimension alignment with modal-specific text prompt tokens through a modal bridge, the text prompt tokens and modal tokens enter a large language multimodal model for interpretation.
-Finally, the interpretation result is matched with downstream tasks by task-specific task heads.
+Following dimension alignment with modality-specific text prompt tokens via a modal bridge, both the text prompt tokens and modality tokens are passed into a large language multimodal model for interpretation.
+The interpretation results are then aligned with downstream tasks through task-specific heads.
 III-B Overview
-TABLE I: List of symbols
+TABLE II: List of symbols
 Module Variants
-Symbol
-Meaning
-Symbol
-Meaning
-$M$
-AllSpark
-$m_{i}$
-modal i
-$f_{i}$
-modal encoder for $m_{i}$
-$q$ query vectors of the bridge
-$\Phi$
-modal bridge
-$s_{i}$ tokens of $m_{i}$
-$T$
-text tokenizer
-$p_{i}$ text prompt for $m_{i}$
-$F$
-Multimodal LLM
-$\theta$ model parameters
-$H_{task}$
-Task Head
-$y$ label
-$L$
-loss function
-$t_{i}$ tokens of $m_{i}$
-${Enc}_{n}$
-n-layer transformer encoders
-$W$ weights of the linear layer
-$Emb$
-embedding layer
-$w_{i}$ word of the code or text
-${Enc}_{ResNet}$
-ResNet
-$Sparse$
-sparse module of the TubeViT
-$Grouper$
-point grouper of the PointBERT
-$Conv1d$
-1D convolution layer
-$FFN$
-feedforward network
-$\sigma$
-softmax
-Guided by the principles mentioned above, the AllSpark we proposed consists of five modules: the modal encoder, modal bridge, text tokenizer, multimodal large language model, and task head.
+Symbol Meaning Symbol Meaning
+$M$ AllSpark $m_{i}$ modal i
+$f_{i}$ modal encoder for $m_{i}$ $q$ query vectors of the bridge
+$\Phi$ modal bridge $s_{i}$ tokens of $m_{i}$
+$T$ text tokenizer $p_{i}$ text prompt for $m_{i}$
+$F$ Multimodal LLM $\theta$ model parameters
+$H_{task}$ Task Head $y$ label
+$L$ loss function $t_{i}$ tokens of $m_{i}$
+${Enc}_{n}$ n-layer transformer encoders $W$ weights of the linear layer
+$Emb$ embedding layer $w_{i}$ word of the code or text
+${Enc}_{ResNet}$ ResNet
+$Grouper$ point grouper of the PointBERT[48]
+$Conv1d$ 1D convolution layer
+$FFN$ feedforward network
+$\sigma$ softmax
+Guided by the principles mentioned above, AllSpark consists of five modules: the modal encoder, modal bridge, text tokenizer, multimodal large language model, and task head.
 The overall architecture is depicted in Figure 3.
@@
 $s_{i}=\Phi\left(f_{i}\left(m_{i}\right),q\right)$ (1)
-Here, $\{m_{RGB},m_{MSI},m_{HSI},\ldots,m_{i},\ldots,m_{Video}\}$ represents inputs from various modalities, where $m_{RGB}$ represents the RGB modality, $m_{MSI}$ represents the MSI modality, and so on.
+Here, $\{m_{RGB},m_{MSI},m_{HSI},\ldots,m_{i},\ldots\}$ represents inputs from various modalities, where $m_{RGB}$ represents the RGB modality, $m_{MSI}$ represents the multispectral modality, and so on.
 $f_{i}$ is defined as the modal encoder for the $m_{i}$ modality, $\Phi$ represents the modal bridge, and $q\in R^{N*D}$represents $N$ learnable vectors of dimension $D$ in the modality bridge, where $D$ is set to 4096, representing the dimensionality of the multimodal large language model.
 The input data $m_{i}$ of each modality are mapped to a token sequence $s_{i}\in R^{N*D}$ of the same dimension as the language model.
-Table I summarizes the main mathematical symbols and their meanings in the model.
+Table II summarizes the main mathematical symbols and their meanings in the model.
 To achieve cohesion among modalities, we employ a unified multimodal large language model to parse data from various modalities.
 The text tokenizer and multimodal large language model in AllSpark are based on the visual-language model Lynx[49].
-To extend Lynx to thirteen spatio-temporal modalities, we designed specific text prompts for each modality to guide the model in correctly parsing information from each modality.
+To extend Lynx to ten spatio-temporal modalities, we designed specific text prompts for each modality to guide the model in correctly parsing information from each modality.
 Additionally, Lynx incorporates several lightweight multimodal adapter layers internally to accommodate multimodal inputs.
@@
 III-C1 1D modal
-Code & Text: Code is essentially a specialized form of language, but due to its distinct properties such as having a strict syntax, being unambiguous, and being capable of interacting with machines.
-we separate it as a distinct modality.
+Code & Language: Code is essentially a specialized form of language.
+But due to its distinct properties such as having a strict syntax, being unambiguous, and being capable of interacting with machines, we separate it as a distinct modality.
+Therefore, to avoid ambiguity, we will use ”text” and ”language” interchangeably to distinguish between the natural language and code modalities.
 Exploring intelligent methods for the code modality is crucial for reliable AI reasoning and achieving interaction between intelligent models and the real world.
@@
 Instead, we directly utilize Lynx’s text tokenizer, i.e., $f\left(m_{Text/Code}\right)=T\left(m_{Text/Code}\right)$, where $m_{Text/Code}=\{w_{1},w_{2},w_{3},\ldots\}$ represents the sequence of words in the text or code.
+Table: A table can be viewed as a sequence of rows, with each row containing several fields, or columns, i.e., $m_{Table}\in R^{{row}*{col}}$.
+The modal encoder for the table modality inherits the design from TabFormer[50]: first, based on the different degrees of discreteness for each column attribute, we use independent embedding layers to encode discrete and continuous values separately.
+Subsequently, we employ a single-layer transformer encoder to further extract features.
+The entire process can be formalized as $f\left(m_{Table}\right)={Enc}_{1}\left(Emb(m_{Table})\right)$.
 III-C2 2D modal
 RGB: RGB imagery represents the visible light spectrum and reflects the electromagnetic characteristics of objects that emit or reflect visible light waves.
-It is the most common modality used in visual research.
+It is the most common modality in the field of computer vision.
 RGB imagery is a standard three-band image, i.e., $m_{RGB}\in R^{H*W*3}$.
-For the modal encoder of this modality, we adopted the visual encoder from the Lynx model: the EVA.
-The EVA is a large visual backbone model composed of 40 stacked transformer blocks with a width of 1408.
+For the modal encoder of this modality, we adopted the visual encoder from the Lynx model: EVA[51].
+The EVA is a large visual model composed of 40 stacked transformer blocks with a width of 1408.
 During the experiments, AllSpark loaded the official weights of the EVA model and froze them during training.
 MSI: Multispectral imagery is a modality extensively studied in the remote sensing field.
-Typically, it includes more than the three visible light bands present in the RGB modality.
-MSIs incorporate multiple nonvisible light bands, such as near-infrared, shortwave infrared, coastal atmospheric aerosol, and cirrus bands.
-Therefore, the number of channels in the imagery is usually greater than the three bands in RGB imagery, i.e., $m_{MSI}\in R^{H*W*C}$, where $C>3$.
+It incorporate multiple nonvisible light bands, such as near-infrared, shortwave infrared, coastal atmospheric aerosol, and cirrus bands.
+Therefore, the number of channels in MSIs is usually greater than the three bands in RGB imagery, i.e., $m_{MSI}\in R^{H*W*C}$, where $C>3$.
 We extended the PatchEmbed of the standard ViT[31], modifying its channel count to match the number of bands in the input multispectral imagery.
@@
 The entire process can be represented as $f\left(m_{HSI}\right)={Enc}_{12}\left(W*m_{HSI}^{T}\right)$.
-Table: The table modality consists of several rows and columns, where each row represents a record and each column represents a different attribute, i.e., $m_{Table}\in R^{{row}*{col}}$.
-The modal encoder for the table modality inherits the design from TabFormer[36]: first, based on the different degrees of discreteness for each column attribute, we use independent embedding layers to encode discrete and continuous values separately.
-Subsequently, we employ a single-layer transformer encoder to further extract features.
-The entire process can be formalized as $f\left(m_{Table}\right)={Enc}_{1}\left(Emb(m_{Table})\right)$.
 Trajectory: The trajectory modality reflects the changing information of an object over time and space and is composed of a series of two-dimensional coordinate points, i.e., $m_{Trajectory}\in R^{l*2}$, where $l$ represents the sequence length of trajectory points.
@@
 Therefore, we designed a simple three-layer convolutional network as the modal encoder.
-Infrared: The dataset we selected for infrared detection was SYSU-MM01, which is, in fact, a multimodal reidentification dataset comprising both visible and near-infrared (NIR) modalities.
-The model needs to simultaneously process infrared and visible light images, i.e., $m_{Infrared}=\{{I_{r}\in R}^{H*W*3},{I_{i}\in R}^{H*W*3}\}$.
-We inherited and slightly modified the design of DDAG[38], discarding the parameters of the shared ResNet-50 layers in DDAG and employing two independently parameterized ResNet-50 networks.
-These networks extract features from infrared and visible light images, and this process can be formalized as $f\left(m_{Infrared}\right)={Enc}_{ResNet}\left(I_{r}\right)\oplus\ {Enc}_{ResNet}\left(I_{i}\right)$.
 Graph: A graph is constructed from a series of nodes and edges, where the attributes of the nodes and the adjacency attributes of the nodes reflect the majority of the features of the graph, i.e., $m_{Graph}\in R^{K*d}$, where K is the number of nodes and d is the node feature dimension.
-In AllSpark, the modality encoder of a graph is based on the STAEformer[39], whose main design idea is to first use a linear layer to extend its feature dimension and then use several embedding layers to separately encode features such as the node’s characteristics, spatial characteristics, and temporal characteristics.
+In AllSpark, the modality encoder of a graph is based on the STAEformer[53, 54], whose main design idea is to first use a linear layer to extend its feature dimension and then use several embedding layers to separately encode features such as the node’s characteristics, spatial characteristics, and temporal characteristics.
 The entire process can be formalized as $f\left(m_{Graph}\right)={Emb}_{node}\left({W*m}_{Graph}^{T}\right)\oplus{Emb}_{spatial}\left({W*m}_{Graph}^{T}\right)\oplus{Emb}_{time}\left({W*m}_{Graph}^{T}\right)$, where $W\in R^{{hidden}*d}$ is the weight of the linear layer.
-Oblique Photography: The oblique photography modality consists of images from multiple views, i.e., $m_{Oblique}=\{V_{1}\in R^{C*H*W},\ V_{3}\in R^{C*H*W},V_{3}\in R^{C*H*W},\ldots\ ,V_{n}\in R^{C*H*W}\}$, where $V_{i}$ represents the image of each view and n is the number of views.
-In the encoder for the oblique photography modality, we use the standard ViT model to extract features from multiple view images and concatenate them.
-The process can be formalized as $f\left(m_{Oblique}\right)={Enc}_{12}\left(V_{1}\right)\oplus{Enc}_{12}\left(V_{2}\right)\oplus{Enc}_{12}\left(V_{3}\right)\oplus\ldots\oplus{Enc}_{12}\left(V_{n}\right)$.
 III-C3 3D modal
-Video: The video modality can be considered as the stacking of images in the temporal dimension, i.e., $m_{video}\in R^{T*C*H*W}$.
-The significant characteristic of the video modality is its high redundancy in both temporal and spatial dimensions.
-To address this issue, the encoder for the video modality in AllSpark is based on TubeViT[40], which performs sparse sampling on highly redundant data in both time and space to reduce redundancy.
-The features obtained from sparse sampling are subsequently fed into a 6-layer transformer encoder for further feature extraction.
-The process can be formalized as $f\left(m_{video}\right)={Enc}_{6}\left({Sparse}(m_{video})\right)$.
-PointClouds: PointClouds are typically composed of three-dimensional coordinates and feature values $m_{PointCloud}\in R^{K*\left(d+3\right)}$, where K represents the number of three-dimensional points and d represents the dimensionality of the point cloud features, reflecting information such as the spatial position, shape, colour, and texture of objects.
+Point Cloud: A point cloud is typically composed of three-dimensional coordinates and feature values $m_{PointCloud}\in R^{K*\left(d+3\right)}$, where K represents the number of three-dimensional points and d represents the dimensionality of the point cloud features, reflecting information such as the spatial position, shape, colour, and texture of objects.
 The encoder for the point cloud modality inherits the design from PointBERT[48]: first, point cloud data are grouped and encoded to unify the number of points simultaneously inputted.
@@
 Although the modal encoders have transformed data from various heterogeneous modalities into a unified token sequence, there are still differences in dimensions between different modal tokens, making it difficult to perceive by a multimodal large language model.
-The modal bridge, based on the Lynx model, aims to perform dimensional projection from tokens of various modalities to tokens of the language modality.
-In its implementation, the modal bridge consists of stacked layers, including cross-attention layers and feedforward neural network layers.
+The modal bridge, based on the Perceiver[21], aims to perform dimensional projection from tokens of various modalities to tokens of the language modality.
+In its implementation, the modal bridge consists of stacked cross-attention layers and feedforward neural network layers.
 In the cross-attention layer, we predefine a learnable query vector $Q\in R^{N*D}$, where $D$ is the internal dimensionality of the language model and $N$ serves as a hyperparameter that can be flexibly adjusted to accommodate inputs from different modalities.
@@
 III-E Task-Guided Text Prompts and Task Heads
-To extend the visual-language multimodal model to thirteen spatio-temporal modalities without intervention from modality expert knowledge, we designed specific text prompts and task heads for each modality and task.
+To extend the visual-language multimodal model to ten spatio-temporal modalities without intervention from modality expert knowledge, we designed specific text prompts and task heads for each modality and task.
 Text prompts are used to guide the multimodal language model in correctly interpreting each modality’s data, while task heads are employed to match the model’s parsing results with specific downstream tasks.
@@
 However, during testing, for the sake of result stability and reproducibility, the prompt was fixed to be the first prompt among all prompts.
-Table XVI provides a list of all the text prompts.
-Additionally, we noticed that there is still a gap between the multimodal language model’s parsing results and the specific downstream tasks.
-To enhance the model’s generalization capability on downstream tasks, we designed specific task heads for each task.
-Moreover, to ensure the transferability of modalities across different tasks, the design principle for task heads is to be as simple and lightweight as possible.
-A detailed summary of each modality’s task heads can be found in Table 4.
+Table XVII provides a list of all the text prompts.
+To ensure the transferability of modalities across different tasks, the design principle for task heads is to be as simple and lightweight as possible.
 For classification tasks or downstream tasks that can be formalized as classification tasks, we uniformly use a simple single-layer linear layer as the task head.
-For instance, we implemented standard classification tasks on the RGB, MSI, SAR, video, and PointCloud modalities.
+For instance, we implemented standard classification tasks on the RGB, MSI, SAR, and point cloud modalities.
 Although the task involving the HSI modality is segmentation, it can be formalized as a per-pixel classification task.
-Similarly, the reidentification task in the infrared modality can also be formalized as a classification task.
 Therefore, a single-layer linear layer is used as the task head for the mentioned modalities.
@@
 The only difference from classification tasks is the addition of an unscaled operation without learnable parameters.
-Since the Lynx model itself is a language model, the code and text modalities directly use its native text decoder.
-The task head for the oblique photography modality is based on the 3D reconstruction architecture from the Ada-MVS[42] model.
+Since the Lynx itself is a language model, the code and text modalities directly use its native text decoder.
+TABLE III: Summary of hyperparameters
+Dimension Modal Max Lr Max Epochs Warm-up Epochs
+1D Language $9.0\times 10^{6}$ 5 1
+Code $1.0\times 10^{5}$ 4 1
+Table $2.0\times 10^{5}$ 30 3
+2D RGB $5.0\times 10^{5}$ 50 5
+MSI $2.0\times 10^{5}$ 50 5
+HSI $1.0\times 10^{4}$ 30 3
+SAR $9.0\times 10^{6}$ 30 3
+Trajectory $1.0\times 10^{5}$ 30 5
+Graph $8.0\times 10^{5}$ 10 2
+3D Point Cloud $3.0\times 10^{5}$ 100 10
 IV Experiment
 IV-A Setup
-TABLE II: Summary of hyperparameters
-Dimension
-Modal
-Max Lr
-Max Epochs
-Warm-up Epochs
-1D
-Text
-$9.0\times 10^{6}$
-5
-1
-Code
-$1.0\times 10^{5}$
-4
-1
-2D
-RGB
-$5.0\times 10^{5}$
-50
-5
-MSI
-$2.0\times 10^{5}$
-50
-5
-HSI
-$1.0\times 10^{4}$
-30
-3
-Infrared
-$5.0\times 10^{5}$
-50
-5
-SAR
-$9.0\times 10^{6}$
-30
-3
-Oblique Photography
-$5.0\times 10^{5}$
-30
-3
-Table
-$2.0\times 10^{5}$
-30
-3
-Trajectory
-$1.0\times 10^{5}$
-30
-5
-Graph
-$8.0\times 10^{5}$
-10
-2
-3D
-PointCloud
-$3.0\times 10^{5}$
-100
-10
-Video
-$1.0\times 10^{5}$
-3
-1
-Our experiments evaluated the adaptability of AllSpark to various spatio-temporal modalities.
-To provide a realistic evaluation of its multimodal adaptability and facilitate comparisons with existing work in the field, we followed the principles of simplicity, reproducibility, and choosing datasets and tasks that have been widely studied in the domain.
+Our experiments aim to demonstrate: (1) the distinct advantages of AllSpark over traditional models, and (2) AllSpark’s ability to understand 10 spatio-temporal modalities simultaneously.
+For the former, we believe that since language as reference framework, AllSpark’s advantage over traditional models lies in the richer semantic meaning of its features.
+In large language models[55, 26, 27, 56], a common way to measure this property is through few-shot learning.
+Therefore, we evaluated AllSpark’s performance on few-shot classification tasks in the RGB and point cloud modalities without any extra training steps.
+The experimental results can be found in Section IV-C.
+For the latter, we select a task for each modality to conduct the evaluation.
+Following the principles of simplicity and reproducibility, we choose the widely studied datasets in each modality’s respective field and employ similar experimental settings across all modalities.
+Specifically, we use the AdamW optimizer with a learning rate schedule based on cosine annealing.
+The hyperparameters are adjusted slightly in terms of training epochs and learning rates for different experiments.
 The specific details on dataset selection can be found in Section IV-B.
-Additionally, for reproducibility, we employed similar experimental settings across all modalities.
-We used the AdamW optimizer with a learning rate schedule based on cosine annealing.
-The hyperparameters were adjusted slightly in terms of training epochs and learning rates for different experiments.
-Table II summarizes the hyperparameter settings for the experiments on each modality.
+Table III summarizes the hyperparameter settings for the experiments on each modality.
+The experimental results can be found in Section IV-D.
 IV-B Dataset
-Table 4 summarizes the data and tasks used in the experiments for each modality. Below, we will provide detailed explanations in order:
-Text: The IMDB[43] dataset is a binary sentiment analysis dataset consisting of 50,000 reviews from the Internet Movie Database (IMDb) labelled as positive or negative.
+Below, we will provide detailed explanations of datasets in order:
+Language: The IMDB[57] dataset is a binary sentiment analysis dataset consisting of 50,000 reviews from the Internet Movie Database (IMDb) labelled as positive or negative.
 Additionally, the dataset includes some unlabelled data.
@@
 The task performed on the code modality is code document generation, and we tested it on the Ruby and JavaScript.
+TABLE IV: The few-shot classification results for the RGB modality
+Model Training-free UC Merced RS19
+5way-1shot 5way-5shot 5way-1shot 5way-5shot
+MatchingNet[59] 48.18 67.39 67.68 85.01
+RelationNet[60] 50.07 65.22 65.01 79.75
+ProtoNet[61] 53.85 71.23 76.36 85.00
+DLA-MatchNet[62] 53.76 63.01 68.27 79.89
+SPNet[63] 57.64 73.52 81.06 88.04
+AllSpark 95.58 97.64 97.16 98.94
+TABLE V: The few-shot classification results for the point modality
+Model Training-free ShapeNet ScanObject
+5way-1shot 5way-5shot 5way-1shot 5way-5shot
+FSLGNN[64] 64.98 76.14 29.91 32.77
+RelationNet[60] 65.88 76.25 45.32 55.43
+ProtoNet[61] 65.96 78.77 44.75 59.81
+AllSpark 67.20 82.12 35.70 53.33
+Table: The PRSA[65] dataset is a collection of air quality data from multiple stations in Beijing that contains hourly measurements of air pollutants.
+The data spans from March 1, 2013, to February 28, 2017, across 12 monitoring stations.
+In our experiments, we used various features, including time, station information, 4 air pollutant variables (SO2, NO2, CO, and O3), and 6 meteorological variables (temperature, pressure, dew point temperature, amount of precipitation, wind speed, and wind direction).
+The task was to predict the concentration of PM2.5.
+We split the data into training (40%) and testing sets.
 RGB: NWPU-RESISC45[66] is a large-scale open dataset for visible light remote sensing image scene classification.
@@
 We used a 4:6 split for training and testing.
-Table: The PRSA[47] dataset is a collection of air quality data from multiple stations in Beijing that contains hourly measurements of air pollutants.
-The data spans from March 1, 2013, to February 28, 2017, across 12 monitoring stations.
-In our experiments, we used various features, including time, station information, 4 air pollutant variables (SO2, NO2, CO, and O3), and 6 meteorological variables (temperature, pressure, dew point temperature, amount of precipitation, wind speed, and wind direction).
-The task was to predict the concentration of PM2.5.
-We split the data into training (40%) and testing sets.
 Trajectory: The ETH-UCY[68, 69] dataset is a widely used benchmark for pedestrian trajectory prediction and is divided into five subsets: ETH, HOTEL, UNIV, ZARA1, and ZARA2.
@@
 [34], ensuring that the serial numbers and target configurations are consistent between the test and training sets while the aspects and depression angles differ.
-Infrared: SYSU-MM01[51] is a cross-modal human reidentification dataset with optical-infrared modalities that includes images from 4 RGB cameras and 2 infrared cameras of 491 different individuals.
-We adopted the official dataset split, with the training set containing 20,284 RGB images and 9,929 infrared images from 296 individuals and the test set consisting of 3,803 infrared images from 96 individuals used as queries and 301 randomly selected RGB images for reidentification targets.
 Graph: METR-LA is a traffic dataset collected from loop detectors on the Los Angeles highways spanning from March 1, 2012, to June 30, 2012. The task is traffic flow prediction.
-Oblique Photography: WHU-OMVS[42] is an oblique photography dataset designed for 3D reconstruction tasks.
-The dataset provides imagery from five different viewpoints, along with camera parameters and other relevant information.
-It consists of six areas, and in our experiments, area 1 is used as the training set, and area 2 is used as the test set.
-Video: The UCF101[52] dataset is a human action recognition dataset comprising 101 action classes with a total of 13,320 video clips.
-The videos have a combined duration of 27 hours and a resolution of 320*240 pixels and were sourced from YouTube.
 Point cloud: ModelNet40[71] is a synthetic point cloud dataset consisting of 40 object categories and a total of 12,311 point cloud objects.
 We follow the official dataset split, with 9,843 objects used for training and 2,468 for testing.
-IV-C Results
+IV-C Few-shot learning
+To demonstrate AllSpark’s unique advantages over traditional models, we tested its few-shot performance on the RGB and point cloud modalities.
+It is worth noting that traditional few-shot learning methods typically require additional training steps.
+For instance, common approaches like ProtoNet[61] and MatchingNet[59] involve randomly splitting support and query sets on the training set for supervised training, a step known as meta-learning.
+Thanks to the integration of natural language, AllSpark requires no extra training step and can directly evaluate few-shot classification accuracy on the test set, significantly outperforming baseline models.
 IV-C1 RGB
-TABLE III: rgb image classification with allspark
-Method
-Publication
-Acc(%)
-CNN-CapsNet[54]
-RS2019
-89.03
-DFAGCN[55]
-TNNLS2021
-89.29
-D-CNN with GoogleNet[56]
-TGRS2018
-90.49
-D-CNN with VGGNet[56]
-TGRS2018
-91.89
-SCCov[57]
-TNNLS2019
-92.1
-SeCo-ResNet-50[58]
-ICCV2021
-92.91
-MG-CAP[59]
-TIP2020
-92.95
-LSENet[60]
-TIP2021
-93.34
-MSANet[61]
-JSTARS2021
-93.52
-IDCCP[62]
-TGRS2021
-93.76
-MBLANet[63]
-TIP2021
-94.66
-GRMANet-ResNet-50[64]
-TGRS2021
-94.72
-EMSNet[65]
-TGRS2023
-95.37
-ViTAE-B + RVSA[66]
-TGRS2022
-95.69
-AllSpark
-ours
-94.85
+We evaluated AllSpark’s RGB image few-shot classification performance on the UC-Merced[72] and WHU-RS19[73, 74] datasets, following the dataset splits from [62].
+Specifically, the UC-Merced dataset uses six classes—Beach, Golf course, Mobile home park, River, Sparse residential, and Tennis court—as the test set, while the WHU-RS19 dataset uses five classes—Commercial, Meadow, Pond, River, and Viaduct—as the test set.
+AllSpark does not require meta-learning on the training set and is evaluated on 600 episodes directly on the test set, with 15 query samples per episode.
+As shown in the Table IV, we report the results for both 5-way 1-shot and 5-way 5-shot settings.
+The experiments demonstrate that AllSpark significantly outperforms baseline models without requiring any training, highlighting the advantage of LaRF.
+IV-C2 Point
+We evaluated AllSpark’s point cloud few-shot classification performance on the ShapeNet[75] and ScanObjectNN[76] datasets, following the dataset settings from [77].
+Similar to the RGB modality, AllSpark does not require meta-learning on the training set and is evaluated on 700 episodes directly on the test set, with 15 query samples per episode.
+As shown in the Table V, we report the results for both 5-way 1-shot and 5-way 5-shot settings.
+The experiments demonstrate that AllSpark surpasses most baseline models even on previously unseen point cloud datasets, showcasing its outstanding semantic richness and generalization capabilities.
+IV-D Ability to understand 10 spatio-temporal modalities
+IV-D1 RGB
 We evaluated the performance of AllSpark on the RGB image scene classification task using the NWPU-RESISC45 dataset, with the top-1 accuracy as the evaluation metric.
 AllSpark leverages expert knowledge from Lynx by loading its pretrained weights.
-Therefore, in Table III, we compare AllSpark with the state-of-the-art models known on the NWPU-RESISC45 dataset.
+Therefore, in Table VI, we compare AllSpark with the state-of-the-art models.
 The results indicate that AllSpark outperforms most baseline models, with a margin of only 0.84 compared to that of the SOTA (95.69).
 This highlights AllSpark’s exceptional perception and interpretation capabilities in the RGB modality.
-IV-C2 MSI
+TABLE VI: RGB image classification with allspark
+Method Publication Acc(%)
+CNN-CapsNet[78] RS2019 89.03
+DFAGCN[79] TNNLS2021 89.29
+D-CNN with GoogleNet[80] TGRS2018 90.49
+D-CNN with VGGNet[80] TGRS2018 91.89
+SCCov[81] TNNLS2019 92.1
+SeCo-ResNet-50[82] ICCV2021 92.91
+MG-CAP[83] TIP2020 92.95
+LSENet[84] TIP2021 93.34
+MSANet[85] JSTARS2021 93.52
+IDCCP[86] TGRS2021 93.76
+MBLANet[87] TIP2021 94.66
+GRMANet-ResNet-50[88] TGRS2021 94.72
+EMSNet[89] TGRS2023 95.37
+ViTAE-B + RVSA[90] TGRS2022 95.69
+AllSpark ours 94.85
+IV-D2 MSI
 We evaluated the performance of AllSpark on the MSI scene classification task using the EuroSAT dataset.
@@
 The model’s objective was to correctly classify the images into one of the 10 specified categories, and the evaluation metric chosen was the top-1 accuracy.
-AllSpark does not possess expert knowledge in the multispectral modality, so we categorize the baseline models into two groups, as shown in Table IV: those with expert knowledge intervention and those without.
+AllSpark does not possess expert knowledge in the multispectral modality, so we categorize the baseline models into two groups, as shown in Table VII: those with expert knowledge intervention and those without.
 Expert knowledge intervention refers to baseline models pretraining on large datasets such as BigEarthNet and then fine-tuning on the EuroSAT dataset, while no expert knowledge indicates baseline models trained directly from scratch on the EuroSAT dataset.
@@
 Furthermore, AllSpark lags behind the best result in the expert knowledge group by only 4.75, demonstrating its excellent adaptability to the multispectral modality.
-TABLE IV: msi land cover classification with allspark
-Expert Know.
-Method
-Publication
-Acc(%)
-No
-ResNet-18[58]
-ICCV2021
-63.21
-ResNet-50[67]
-IGARSS2023
-91.13
-InceptionNet[68]
-ICIP2020
-93.07
-EfficientNet[69]
-IJSTARS2021
-93.94
+TABLE VII: MSI land cover classification with allspark
+Expert Know. Method Publication Acc(%)
+ResNet-18[82] ICCV2021 63.21
+ResNet-50[91] IGARSS2023 91.13
+InceptionNet[92] ICIP2020 93.07
+EfficientNet[93] IJSTARS2021 93.94
 AllSpark ours 94.03
-ResNet-152[67]
-IGARSS2023
-96.63
-Yes
-MoCoV2[58]
-ICCV2021
-89.51
-SeCo[58]
-ICCV2021
-93.14
-SEER[70]
-Arxiv2022
-97.6
-DINO-MC[71]
-Arxiv2023
-98.78
-IV-C3 HSI
+ResNet-152[91] IGARSS2023 96.63
+MoCoV2[82] ICCV2021 89.51
+SeCo[82] ICCV2021 93.14
+SEER[94] Arxiv2022 97.6
+DINO-MC[95] Arxiv2023 98.78
+IV-D3 HSI
 We conducted a pixel classification task on the Pavia University dataset for the hyperspectral modality.
@@
 Uchaev [96].
-The results in Table V demonstrate that AllSpark outperforms many hyperspectral image classification methods, such as IFRF and S-DMM, by a factor of 6.42 compared to the best result in terms of OA, highlighting AllSpark’s superior adaptability to the hyperspectral modality.
-TABLE V: hsi pixel classification with allspark
-Method
-OA
-AA
-Kappa
+The results in Table VIII demonstrate that AllSpark outperforms many hyperspectral image classification methods, such as IFRF and S-DMM, by a factor of 6.42 compared to the best result in terms of OA, highlighting AllSpark’s superior adaptability to the hyperspectral modality.
+TABLE VIII: HSI pixel classification with allspark
+Method OA AA Kappa
 3D-CNN 75.24 80.26 68.34
@@
 PRNet-RF 95.6 94.96 94.27
-IV-C4 Table
+IV-D4 Table
 For the table modality, we evaluated AllSpark on the regression prediction task using the PRSA[65] dataset.
@@
 The performance metrics include the root mean squared error (RMSE), mean absolute error (MAE), and R-squared (R2).
-TABLE VI: PM2.5 prediction with allspark
+TABLE IX: PM2.5 prediction with allspark
 Expert Arch. Method PM2.5
 RMSE MAE R2
-No GWO[73] 62.2 40.8 -
+GWO[97] 62.2 40.8 -
 TabBERT[50] 32.8 - -
 AllSpark 29.03 18.04 0.87
-Yes Stacked ResNet-LSTM[74] 40.68 23.75 0.8
+Stacked ResNet-LSTM[98] 40.68 23.75 0.8
 CBAM-CNN-Bi-LSTM[99] 18.9 11.2 0.94
-Table VI presents the comparative results between AllSpark and the baselines.
+Table IX presents the comparative results between AllSpark and the baselines.
 It is worth noting that some works specifically focus on the prediction task on the PRSA dataset and design expert models with specific architectures based on dataset characteristics.
@@
 This reflects the excellent adaptability of AllSpark to the table modality.
-IV-C5 Code
+IV-D5 Code
 For the code modality, we evaluated the performance of AllSpark on the code document generation task using the CodeSearchNet dataset.
@@
 We conducted tests for both the Ruby and JavaScript languages using the mean reciprocal rank (MRR)[100] as the evaluation metric.
-TABLE VII: code document generation with allspark
+TABLE X: code document generation with allspark
 Expert Know. Method MRR
 Ruby Javascript
-No BIRNN[68] 0.084 0.153
+BIRNN[68] 0.084 0.153
 1D-CNN[68] 0.245 0.352
@@
 AllSpark 0.627 0.635
-Yes RoBERTa(Code)[67] 0.661 0.64
+RoBERTa(Code)[67] 0.661 0.64
 CodeBERT[67] 0.693 0.706
@@
 Like in the previous modalities, since MSI-AGI does not possess expert knowledge of the code modality, we categorized the baselines into two groups: those with and without expert knowledge.
-As shown in Table VII, AllSpark achieved SOTA results in the group without expert knowledge, and the results were comparable to those of models trained with expert knowledge.
+As shown in Table X, AllSpark achieved SOTA results in the group without expert knowledge, and the results were comparable to those of models trained with expert knowledge.
 This finding demonstrates the strong adaptability of AllSpark to the code modality.
-IV-C6 PointCloud
+IV-D6 Point Cloud
 For the point cloud modality, we evaluated the performance of AllSpark on the ModelNet40 dataset for the classification task, with the top-1 accuracy as the metric.
@@
 Additionally, some methods tend to pretrain on large point cloud datasets to acquire general modal expert knowledge before generalizing to specific downstream tasks to improve performance.
-TABLE VIII: pointcloud classification with allspark
-Method
-Expert Architecture Expert Knowledge Acc(%)
-PointNet[1]
-Yes No 89.2
-Kd-net[77]
-Yes No 90.6
-SPH3D-GCN[78]
-Yes No 91.4
-PointNet++[79]
-Yes No 91.9
-SO-Net[80]
-Yes No 92.5
-PointVGG[81]
-Yes No 93.6
-PointBERT[41]
-No Yes 93.8
-PointGPT[82]
-Yes Yes 94.9
-AllSpark
-No No 91.2
+TABLE XI: pointcloud classification with allspark
+Method Expert Architecture Expert Knowledge Acc(%)
+PointNet[10] 89.2
+Kd-net[101] 90.6
+SPH3D-GCN[102] 91.4
+PointNet++[103] 91.9
+SO-Net[104] 92.5
+PointVGG[105] 93.6
+PointBERT[48] 93.8
+PointGPT[106] 94.9
+AllSpark 91.2
 The AllSpark module is designed based on a general sequence-to-sequence architecture.
-As shown in Table VIII, in the absence of both modality expert architectural designs and modal expert knowledge, AllSpark still outperforms classical networks with point cloud-specific structures (PointNet, Kd-Net).
+As shown in Table XI, in the absence of both modality expert architectural designs and modal expert knowledge, AllSpark still outperforms classical networks with point cloud-specific structures (PointNet, Kd-Net).
 This approach maintains comparability with the state-of-the-art PointGPT model, which has both a modal expert structure and modal expert knowledge.
 This finding suggested that AllSpark has significant potential for applications in the point cloud modality.
-IV-C7 Trajectory
+IV-D7 Trajectory
 For the trajectory modality, we evaluated the performance of AllSpark on the ETH dataset in the trajectory prediction task.
@@
 ${FDE}=\sqrt{\left(x_{T}-\widehat{x_{T}}\right)^{2}+\left(y_{T}-\widehat{y_{T}}\right)^{2}}$ (6)
-In Table IX, AllSpark is compared with the state-of-the-art trajectory prediction models.
+In Table XII, AllSpark is compared with the state-of-the-art trajectory prediction models.
 The results indicate that AllSpark, which uses a unified structure without trajectory modality expert knowledge, outperforms most expert models.
@@
 This finding suggested that AllSpark demonstrated excellent adaptability to the trajectory modality.
-TABLE IX: trajectory prediction with allspark
-Method
-ADE/FDE
-Social GAN
-0.87/1.62
-SoPhie
-0.70/1.43
-STAR
-0.36/0.64
-SGCN
-0.63/1.03
-CAGN
-0.41/0.65
-SIT
-0.39/0.62
-SocialVAE
-0.47/0.76
-PCENet
-0.54/0.87
-AgentFormer
-0.45/0.75
-MemoNet
-0.40/0.61
-SocialVAE+FPC
-0.41/0.58
-TUTR
-0.40/0.61
-AllSpark
-0.43/0.69
-IV-C8 SAR
+TABLE XII: trajectory prediction with allspark
+Method ADE/FDE
+Social GAN 0.87/1.62
+SoPhie 0.70/1.43
+STAR 0.36/0.64
+SGCN 0.63/1.03
+CAGN 0.41/0.65
+SIT 0.39/0.62
+SocialVAE 0.47/0.76
+PCENet 0.54/0.87
+AgentFormer 0.45/0.75
+MemoNet 0.40/0.61
+SocialVAE+FPC 0.41/0.58
+TUTR 0.40/0.61
+AllSpark 0.43/0.69
+IV-D8 SAR
 For the SAR modality, the adaptability of AllSpark was tested on the MSTAR dataset, where the model is required to identify SAR images of ten military targets, and the metric used is the top-1 accuracy.
 In the experiment, the preprocessing of the MSTAR dataset followed the SOC settings from AConvNets[34].
-Table X presents the comparison results between AllSpark and the state-of-the-art model under these settings.
-Since AllSpark lacks expert knowledge of SAR image modalities and spatial invariance prior to changes in perspective; its performance on the MSTAR dataset is suboptimal, with an accuracy of 88%.
-This approach is comparable to the EMACH method.
-Enhancing the model’s adaptability to SAR images and improving its ability to extract spatial information are directions for future research.
-TABLE X: sar classification with allspark
-Method
-Acc(%)
-EMACH[83]
-88
-SVM[83]
-90
-AdaBoost[83]
-92
-MSRC[84]
-93.6
-IGT[83]
-95
-MSS[85]
-96.6
-Cond Gauss[86]
-97
-M-PMC[87]
-98.8
-AConvNets[22]
-99.13
-AllSpark
-88
-IV-C9 Infrared
-AllSpark’s adaptability to the infrared modality was tested on the SYSU-MM01 dataset in a cross-modal human reidentification task involving RGB and infrared modalities.
-In this experiment, the model receives both an infrared image and an RGB image and is tasked with identifying and matching the same person across different modality images.
-Table XI compares AllSpark and the other state-of-the-art reidentification methods, with the evaluation metric being rank at an accuracy of 20.
-The results indicate that AllSpark (76.31%) achieved similar accuracy to that of HSME (77.95%).
-While our method did not reach the highest accuracy, it surpassed some expert models in RGB-infrared reidentification (e.g., TONE, HCML) when using a unified architecture without modal expert knowledge, demonstrating the potential of AllSpark in handling the infrared modality.
-TABLE XI: visible-infrared reidentification with allspark
-Method
-Rank-20(All-search)
-Two-stream[51]
-65.5
-One-stream[51]
-66.74
-TONE[88]
-68.6
-HCML[88]
-69.17
-Zero-Pad[51]
-71.33
-AllSpark 76.31
-HSME[89]
-77.95
-BDTR[90]
-81.07
-DDAG[38]
-95.81
-IV-C10 Graph
-TABLE XII: traffic prediction with allspark
-Method
-RMSE MAE R2
-HI
-6.8 14.2 10.15
-GWNet
-3.51 7.28 9.96
-DCRNN
-3.54 7.47 10.32
-AGCRN
-3.59 7.45 10.47
-STGCN
-3.6 7.43 10.35
-GTS
-3.59 7.44 10.25
-MTGNN
-3.47 7.21 9.7
-STNorm
-3.57 7.51 10.24
-GMAN
-3.44 7.35 10.07
-PDFormer
-3.62 7.47 10.91
-STID
-3.55 7.55 10.95
-STAEformer
-3.34 7.02 9.7
-AllSpark
-3.81 7.52 11.24
+Table XIII presents the comparison results between AllSpark and the state-of-the-art model under these settings.
+AllSpark achieves 97.24% top1-accuracy, only 1.89% lower than the SOTA model.
+The experiments demonstrate that AllSpark is capable of effectively understanding SAR imagery.
+TABLE XIII: sar classification with allspark
+Method Acc(%)
+EMACH[107] 88
+SVM[107] 90
+AdaBoost[107] 92
+MSRC[108] 93.6
+IGT[107] 95
+MSS[109] 96.6
+Cond Gauss[110] 97
+M-PMC[111] 98.8
+AConvNets[34] 99.13
+AllSpark 97.24
+IV-D9 Graph
 For the graph modality, the performance of AllSpark was evaluated on the traffic flow prediction task using the METR-LA dataset.
 The evaluation metrics include the RMSE, MAE, and R2.
-Table XII compares AllSpark and the state-of-the-art methods on the METR-LA dataset, with the baseline derived from[39].
+Table XIV compares AllSpark and the state-of-the-art methods on the METR-LA dataset, with the baseline derived from[53].
 The results show that AllSpark, without the intervention of modal expert knowledge, is only 0.47 away from the best result in terms of the RMSE, demonstrating its excellent adaptability to the graph modality.
-IV-C11 Language
+IV-D10 Language
 We tested AllSpark’s natural language processing capabilities on the IMDB dataset[57] with the task of binary sentiment classification (positive or negative).
 AllSpark loads weights from the Lynx; therefore, it can be considered to possess expert knowledge in the natural language modality.
-Compared with the SOTA models on the IMDB dataset, as shown in Table XIII, AllSpark outperformed most language models, with only a 0.32 difference from the SOTA result.
+Compared with the SOTA models on the IMDB dataset, as shown in Table XV, AllSpark outperformed most language models, trailing the SOTA result by only 0.32.
 This highlights AllSpark’s powerful understanding and analysis capabilities in natural language.
-TABLE XIII: text understanding with allspark
-Method
-Acc(%)
-RoBERTa[91]
-95.3
-ULMFiT[92]
-95.4
-BERT[93]
-95.49
-Mixed VAT[94, 95]
-95.68
-LongFormer[96]
-95.7
-XLNet[97]
-96.8
-ERNIE-Doc-Large[98]
-97.1
-AllSpark
-96.78
-IV-C12 Video
-For the video modality, we evaluated AllSpark’s performance on action recognition tasks using the UCF101 dataset.
-The model is tasked with understanding videos and accurately classifying them into one of the 101 classes, with the evaluation metric being the top-1 accuracy.
-In Table XIV, we compare AllSpark with the current state-of-the-art models.
-Currently, AllSpark’s adaptation to the video modality is not optimal, as it shows a significant difference from the baseline model results.
-We attribute this to two main reasons: 1.
-The high redundancy in video information increases the training cost for AllSpark.
-We trained it for only 3 epochs on the UCF101 dataset.
-2.
-AllSpark’s model architecture lacks flexibility for three-dimensional data, making it less effective at capturing temporal information.
-TABLE XIV: video classification with allspark
-Method
-Acc(%)
-OPN
-59.6
-VCOP
-72.4
-SpeedNet
-81.1
-VTHCL
-82.1
-CVRL
-94.4
-VideoMAE v1
-96.1
-VideoMAE v2
-99.6
-AllSpark
-27.5
-IV-C13 Oblique Photography
-For the oblique photography modality, we tested AllSpark’s performance on the 3D reconstruction task using the WHU-OMVS dataset.
-The model takes five-view images as input, and the goal is to output depth maps for reconstructing 3D models.
-The evaluation metric used was the percentage of accurate grids in total (PAG)[42], calculated by the following formula:
-$PAG_{a}=\left(\frac{m_{a}}{m}\right)*100$ (7)
-The suffixes in the PAG represent different accuracy standards, where ${PAG}_{6}$ signifies an error within 0.6 metres and ${PAG}_{10}$ indicates an error within 1 metre.
-Table XV compares AllSpark with popular multiview 3D reconstruction models on the WHU-OMVS dataset.
-AllSpark falls short in terms of accuracy compared to modality-specific expert models.
-We speculate two possible reasons: 1.
-Dense spatial information gradually diminishes in the deep structure of large models, a point verified in our exploratory experiments involving segmentation, detection, etc.; 2.
-The model architecture lacks flexibility and cannot connect gradual features like expert 3D reconstruction models such as Ada-MVS.
-Additionally, these methods lack the ability to design specific model structures for processing, severely restricting their performance.
-Adapting multimodal large models to dense prediction tasks and optimizing the architecture are future research directions.
-TABLE XV: 3D reconstruction with allspark
-Method
-${PAG}_{6}$
-${PAG}_{10}$
-MVSNet
-81.15 91.44
-CasMVSNet
-95.45 98.02
-Ada-MVS
-96.14 98.1
-UCSNet
-96.25 98.45
-AllSpark
-6.4 10.4
+TABLE XIV: traffic prediction with allspark
+Method RMSE MAE R2
+HI 6.8 14.2 10.15
+GWNet 3.51 7.28 9.96
+DCRNN 3.54 7.47 10.32
+AGCRN 3.59 7.45 10.47
+STGCN 3.6 7.43 10.35
+GTS 3.59 7.44 10.25
+MTGNN 3.47 7.21 9.7
+STNorm 3.57 7.51 10.24
+GMAN 3.44 7.35 10.07
+PDFormer 3.62 7.47 10.91
+STID 3.55 7.55 10.95
+STAEformer 3.34 7.02 9.7
+AllSpark 3.81 7.52 11.24
+TABLE XV: text understanding with allspark
+Method Acc(%)
+RoBERTa[112] 95.3
+ULMFiT[113] 95.4
+BERT[114] 95.49
+Mixed VAT[115, 116] 95.68
+LongFormer[117] 95.7
+XLNet[118] 96.8
+ERNIE-Doc-Large[119] 97.1
+AllSpark 96.78
+IV-E Training and Inference Costs
+As the scale of model parameters grows, the training and inference costs of large language models have been rapidly increasing.
+In this section, we provide detailed the training and inference costs of AllSpark for reference.
+All experiments are conducted on two NVIDIA A6000 48G GPUs, using a numerical precision of torch.float32.
+The hyperparameter settings are detailed in Section IV-A.
+Due to the presence of modality-specific encoders, the activated parameters in AllSpark varies when processing different modalities.
+Table XVI summarizes the total parameters, trainable parameters, training time, and inference computational cost (measured in MACs) for each modality.
+TABLE XVI: training and inference costs
+Modality Total params.(M) Trainable params.(M) Training time Inference MACs(G)
+RGB 8176.02 689.80 $\sim$18 hours 1695.98
+MSI 7455.14 978.87 $\sim$30 hours 363.02
+HSI 7450.58 974.31 $\sim$21 hours 374.61
+Table 7486.67 1010.40 $\sim$44 hours 410.00
+Code 6876.02 268.66 $\sim$59 hours 3606.88
+Point 7594.81 1118.54 $\sim$49 hours 557.38
+Trajectory 7381.00 904.73 $\sim$39 hours 538.58
+SAR 7452.58 976.31 $\sim$87 hours 373.63
+Graph 7223.91 747.64 $\sim$53 hours 1846.28
+Language 6876.02 268.66 $\sim$34 hours 3606.88
 V Discussion
@@
     Lack of interaction between different modalities.
-While AllSpark incorporates interactions between infrared and RGB modalities, for other modalities, it essentially uses a unified model for multiple single-modal tasks.
-This is primarily due to constraints such as the lack of multimodal paired data, the inflexibility of the model architecture, and the limited richness of multimodal interaction tasks.
-In the future, we plan to explore and work more extensively in these directions.
+AllSpark only facilitates interactions between the language and other modalities, without involving more interactions, such as RGB and point cloud, or hyperspectral and multispectral.
+This is primarily due to constraints such as the lack of multimodal paired data.
+However, collecting large-scale paired data for ten modalities is nearly impossible, so we attempt to use language as the alignment reference for each modality, achieving indirect alignment between modalities using unpaired data.
+AllSpark represents the initial effort in this approach and has demonstrated strong adaptability across various modalities.
+Also, the adversarial examples also affect the robustness of the proposed model [120].
+In the future, we plan to explore and expand our efforts in these directions in the further.
   2. 2.
-    There is a lack of in-depth research on multimodal alignment methods.
-While we introduced the basic principles of the LaRF and aligned all the modalities to the language modality, our approach included only introducing the modality bridge from Lynx as the alignment method.
-We have not conducted in-depth research into the various methods of multimodal alignment.
-Moreover, our testing focused on the model’s adaptability to different modalities, and we have not explored potential capabilities such as interactive reasoning, zero-shot learning, or multimodal emergence that the model may possess after alignment with language.
-In the future, we plan to delve deeper into researching multimodal alignment methods, and we will design additional tasks and experiments to explore and reveal the upper limits and boundaries of large model reasoning capabilities.
-  3. 3.
     Initial work.
 Our current work is still in its initial exploratory phase, and we have not carefully refined AllSpark’s adaptability and performance on each modality.
-As a result, the model exhibits suboptimal performance on certain imaging modalities, such as oblique photography, SAR, and video.
+As a result, the model exhibits suboptimal performance on certain modalities, such as oblique photography, and video.
+The experimental results can be found in Table XIX and Table XVIII.
 In the future, we plan to conduct more refined and targeted adjustments for each modality to enhance overall performance.
-  4. 4.
+  3. 3.
     Expensive cost: Multimodal large language models, due to prolonged pretraining, typically possess universal reasoning capabilities in certain modalities.
 We generalize their applicability to other modalities by utilizing modality bridges to project other modalities onto the language modality.
-Although we freeze most of the parameters, fine-tuning even once on 2 A6000 GPUs often requires more than a day.
+As shown in Section IV-E, although we freeze most of the parameters, fine-tuning even once on 2 A6000 GPUs often requires more than a day.
 Given the increasing training costs for large models, exploring methods to generalize their universal reasoning abilities is one of our future research directions.
-  5. 5.
+  4. 4.
     Interesting phenomenon.
@@
 Inspired by human cognitive systems and linguistic philosophy, we propose that the construction of multimodal models follow the fundamental principle of Language as Reference Framework (LaRF).
-Guided by this principle, we use language to balance the cohesion and autonomy of modalities, presenting a unified intelligent model, AllSpark, encompassing thirteen spatio-temporal modalities.
+Guided by this principle, we use language to balance the cohesion and autonomy of modalities, presenting a unified intelligent model, AllSpark, encompassing ten spatio-temporal modalities.
 The experimental results demonstrated that AllSpark exhibited excellent adaptability and application potential across various spatio-temporal modalities, highlighting the feasibility and potential of constructing multimodal models with LaRF.
 AllSpark remains an initial exploratory work, and in the future, we aim to delve deeper into the mechanisms guided by natural language, the efficient generalization of large models, and the transition to an interactive paradigm.
-TABLE XVI: List of prompts
-Dimension
-Modal Prompt
-1D Text 1.Please determine if this movie review is positive or negative?
+Appendix A Appendix
+A-A List of prompts
+The text prompts we used are listed in Table XVII.
+TABLE XVII: List of prompts
+Dimension Modal Prompts
+1D Language 1.Please determine if this movie review is positive or negative?
 Code -
-2D RGB 1.This remote sensing image belongs to which of the following categories: [Category list of the NWPU dataset] 2.Describe this remote sensing image briefly 3.Find a word that is most relevant to this remote sensing image 4.Describe the key elements in this remote sensing image
-MSI 1.Based on the multi-spectral imagery feature description, please classify this object 2.Given the following multi-spectral imagery characteristics, please output the most fitting scene label
-HSI 1.Given the spectral information, can you help determine which class this pixel belongs to?
-2.Here is the spectral data for a pixel.
-Considering the typical characteristics of land cover classes, could you provide a detailed analysis and suggest the most likely class for this pixel?
-3.The spectral information for a pixel is given, but the data is noisy.
-Given the potential variability, which land cover classes should be considered as possible candidates for this pixel?
-Table 1.Please utilize the provided air quality indicators to accurately predict the concentrations of PM2.5 in the atmosphere 2.Given the air quality indicators, please provide a prediction for the PM2.5 levels at this particular moment
-Oblique Photography 1.Please reconstruct the 3D model from the given images of five different views.
-Trajectory 1.Based on their past positions and movements in a crowded environment, predict the future trajectory of a selected pedestrian.
-2.Using the pedestrian trajectory data, along with additional information about the surrounding environment, predict the future path of the pedestrian.
-3.Given the current and past positions of a pedestrian and their neighboring pedestrians, predict the main pedestrian’s trajectory
-SAR 1.Based on the SAR imagery feature description, please classify this object 2.Given the following SAR imagery characteristics, please output the most fitting scene label
-Infrared 1.Given an infrared image of a person, find and highlight the same person in a set of visible light images, paying close attention to features like clothing, posture, and gait 2.Analyze the characteristics of a person in a visible light image, such as clothing texture, color, and body shape.
-Then, locate the person with matching features in a series of infrared images
-Graph 1.Given the current traffic data including vehicle flow rate, average speed, and time of day from the METRLA dataset, predict the traffic flow 2.Analyze the historical data on vehicle speeds and flow rates from the METRLA dataset for the past week.
-Identify any patterns or trends and predict the traffic conditions
-3D Video 1.Please analyze this video and describe the characteristics of the action in detail 2.Watch this short clip and provide a step-by-step description of the action sequence
-Point Cloud 1.Classify the provided point cloud sample into the correct category 2.Look at the point cloud data characteristics and classify the object 3.Please analyze the given point cloud dataset and determine which category it belongs to.
-Focus on the shape and structure evident in the point cloud
+Table
+1.Please utilize the provided air quality indicators to accurately
+predict the concentrations of PM2.5 in the atmosphere
+2.Given the air quality indicators, please provide a prediction
+for the PM2.5 levels at this particular moment
+2D RGB
+1.This remote sensing image belongs to which of the following
+categories: [Category list of the NWPU dataset]
+2.Describe this remote sensing image briefly3.Find a word that
+is most relevant to this remote sensing image
+3.Find a word that is most relevant to this remote sensing image
+4.Describe the key elements in this remote sensing image
+MSI
+1.Based on the multi-spectral imagery feature description, please
+classify this object
+2.Given the following multi-spectral imagery characteristics, please
+output the most fitting scene label
+HSI
+1.Given the spectral information, can you help determine which
+class this pixel belongs to?
+2.Here is the spectral data for a pixel. Considering the typical
+characteristics of land cover classes, could you provide a detailed
+analysis and suggest the most likely class for this pixel?
+3.The spectral information for a pixel is given, but the data is
+noisy. Given the potential variability, which land cover classes
+should be considered as possible candidates for this pixel?
+Trajectory
+1.Based on their past positions and movements in a crowded
+environment, predict the future trajectory of a selected pedestrian
+2.Using the pedestrian trajectory data, along with additional
+information about the surrounding environment, predict the
+future path of the pedestrian
+3.Given the current and past positions of a pedestrian and their
+neighboring pedestrians, predict the main pedestrian’s trajectory
+SAR
+1.Based on the SAR imagery feature description, please classify
+this object
+2.Given the following SAR imagery characteristics, please output
+the most fitting scene label
+Graph
+1.Given the current traffic data including vehicle flow rate, average
+speed, and time of day from the METRLA dataset, predict the
+traffic flow
+2.Analyze the historical data on vehicle speeds and flow rates from
+the METRLA dataset for the past week. Identify any patterns or
+trends and predict the traffic conditions
+3D Point Cloud
+1.Classify the provided point cloud sample into the correct category
+2.Look at the point cloud data characteristics and classify the object
+3.Please analyze the given point cloud dataset and determine which
+category it belongs to. Focus on the shape and structure evident in
+the point cloud
+A-B Video
+For the video modality, we evaluated AllSpark’s performance on action recognition tasks using the UCF101[121] dataset.
+The UCF101 dataset is a human action recognition dataset comprising 101 action classes with a total of 13,320 video clips.
+The videos have a combined duration of 27 hours and a resolution of 320*240 pixels and were sourced from YouTube.
+The model is tasked with understanding videos and accurately classifying them into one of the 101 classes, with the evaluation metric being the top-1 accuracy.
+In Table XVIII, we compare AllSpark with the current state-of-the-art models.
+Currently, AllSpark’s adaptation to the video modality is not optimal, as it shows a significant difference from the baseline model results.
+We attribute this to two main reasons: 1.
+The high redundancy in video information increases the training cost for AllSpark.
+We trained it for only 3 epochs on the dataset.
+2.
+AllSpark’s model architecture lacks flexibility for three-dimensional data, making it less effective at capturing temporal information.
+TABLE XVIII: video classification with allspark
+Method Acc(%)
+OPN 59.6
+VCOP 72.4
+SpeedNet 81.1
+VTHCL 82.1
+CVRL 94.4
+VideoMAE v1 96.1
+VideoMAE v2 99.6
+AllSpark 27.5
+A-C Oblique Photography
+For the oblique photography modality, we tested AllSpark’s performance on the 3D reconstruction task using the WHU-OMVS[122] dataset.
+WHU-OMVS is an oblique photography dataset designed for 3D reconstruction tasks.
+The dataset provides imagery from five different viewpoints, along with camera parameters and other relevant information.
+It consists of six areas, and in our experiments, area 1 is used as the training set, and area 2 is used as the test set.
+The model takes five-view images as input, and the goal is to output depth maps for reconstructing 3D models.
+The evaluation metric used was the percentage of accurate grids in total (PAG)[122], calculated by the following formula:
+$PAG_{a}=\left(\frac{m_{a}}{m}\right)*100$ (7)
+The suffixes in the PAG represent different accuracy standards, where ${PAG}_{6}$ signifies an error within 0.6 metres and ${PAG}_{10}$ indicates an error within 1 metre.
+Table XIX compares AllSpark with popular multiview 3D reconstruction models on the WHU-OMVS dataset.
+AllSpark falls short in terms of accuracy compared to modality-specific expert models.
+We speculate two possible reasons: 1.
+Dense spatial information gradually diminishes in the deep structure of large models, a point verified in our exploratory experiments involving segmentation, detection, etc.; 2.
+The model architecture lacks flexibility and cannot connect gradual features like expert 3D reconstruction models such as Ada-MVS.
+Additionally, these methods lack the ability to design specific model structures for processing, severely restricting their performance.
+Adapting multimodal large models to dense prediction tasks and optimizing the architecture are future research directions.
+TABLE XIX: 3D reconstruction with allspark
+Method ${PAG}_{6}$ ${PAG}_{10}$
+MVSNet 81.15 91.44
+CasMVSNet 95.45 98.02
+Ada-MVS 96.14 98.1
+UCSNet 96.25 98.45
+AllSpark 6.4 10.4